### 21 Jun 2023 v2

# Optimal rates of approximation by shallow ReLUk neural networks and applications to nonparametric regression

Abstract
We study the approximation capacity of some variation spaces corresponding to shallow ReLUk neural networks. It is shown that sufficiently smooth functions are contained in these spaces with finite variation norms. For functions with less smoothness, the approximation rates in terms of the variation norm are established. Using these results, we are able to prove the optimal approximation rates in terms of the number of neurons for shallow ReLUk neural networks. It is also shown how these results can be used to derive approximation bounds for deep neural networks and convolutional neural networks (CNNs). As applications, we study convergence rates for nonparametric regression using three ReLU neural network models: shallow neural network, over-parameterized neural network, and CNN. In particular, we show that shallow neural networks can achieve the minimax optimal rates for learning Hölder functions, which complements recent results for deep neural networks. It is also proven that over-parameterized (deep or shallow) neural networks can achieve nearly optimal rates for nonparametric regression.
Keywords: Neural Network, Approximation Rate, Nonparametric Regression, Spherical Harmonic
Mathematics Subject Classification: 41A25, 62G08, 68T07
1 Introduction
Neural networks generate very popular function classes used in machine learning algorithms [Anthony and Bartlett, 2009; LeCun et al., 2015]. The fundamental building blocks of neural networks are ridge functions (also called neurons) of the form $x\in\mathbb{R}^{d}\mapsto\rho((x^{\intercal},1)v)$, where $\rho:\mathbb{R}\to\mathbb{R}$ is a continuous activation function and $v\in\mathbb{R}^{d+1}$ is a trainable parameter. It is well-known that a shallow neural network with non-polynomial activation
$f(x)=\sum_{i=1}^{N}a_{i}\rho((x^{\intercal},1)v_{i}),$ (1.1)
is universal in the sense that it can approximate any continuous function on any compact set with desired accuracy when the number of neurons $N$ is sufficiently large [Cybenko, 1989; Hornik, 1991; Pinkus, 1999]. The approximation and statistical properties of neural networks with different architectures have also been widely studied in the literature [Yarotsky, 2017; Zhou, 2020b; Bartlett et al., 2019; Schmidt-Hieber, 2020], especially when $\rho$ is a sigmoidal activation or $\rho$ is the ReLUk function $\sigma_{k}(t)=\max\{0,t\}^{k}$, the $k$-power of the rectified linear unit (ReLU) with $k\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
The main focus of this paper is rates of approximation by neural networks. For classical smooth function classes, such as Hölder functions, Mhaskar [1996] (see also [Pinkus, 1999, Theorem 6.8]) presented the optimal approximation rates for shallow neural networks, when the activation function $\rho\in C^{\infty}(\Omega)$ is not a polynomial on some open interval $\Omega$ (ReLUk does not satisfy this condition). Similar results have been established for deep ReLU neural networks in recent works [Yarotsky, 2017, 2018; Shen et al., 2020; Lu et al., 2021]. All these approximation rates are obtained by using the idea that one can construct neural networks to approximate polynomials efficiently. There is another line of works [Barron, 1993; Makovoz, 1996; Klusowski and Barron, 2018; Siegel and Xu, 2020, 2022] studying the approximation rates for functions of certain integral forms (such as (1.2)) by using a random sampling argument due to Maurey [Pisier, 1981]. In particular, Barron [1993] derived dimension independent approximation rate for sigmoid type activations and functions $h$, whose Fourier transform $\widehat{h}$ satisfies $\int_{\mathbb{R}^{d}}|\omega||\widehat{h}(\omega)|d\omega<\infty$. This result has been improved and generalized to ReLU in recent articles [Klusowski and Barron, 2018; Siegel and Xu, 2020, 2022].
In this paper, we give a connection of these two lines of approximation results for neural networks (i.e. the constructive approximation of smooth functions and the random approximation of integral representation). Specifically, we consider the following function class defined on the unit ball $\mathbb{B}^{d}$ of $\mathbb{R}^{d}$ induced by vectors on unit sphere $\mathbb{S}^{d}$ of $\mathbb{R}^{d+1}$ as
$\mathcal{F}_{\sigma_{k}}(M):=\left\{f(x)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu(v):\|\mu\|\leq M,x\in\mathbb{B}^{d}\right\},$ (1.2)
which can be regarded as a shallow ReLUk neural network with infinite width [Bach, 2017]. The variation norm restriction $\|\mu\|\leq M$ gives a constraint on the size of the weights in the network. We study how well $\mathcal{F}_{\sigma_{k}}(M)$ approximates the unit ball of Hölder class $\mathcal{H}^{\alpha}$ with smoothness index $\alpha>0$ as $M\to\infty$. Roughly speaking, our main theorem shows that, if $\alpha>(d+2k+1)/2$, then $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma_{k}}(M)$ for some constant $M$ depending on $k,d,\alpha$, and if $\alpha<(d+2k+1)/2$, we obtain the approximation bound
$\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim M^{-\frac{2\alpha}{d+2k+1-2\alpha}},$
where, for two quantities $X$ and $Y$, $X\lesssim Y$ (or $Y\gtrsim X$) denotes the statement that $X\leq CY$ for some constant $C>0$ (we will also denote $X\asymp Y$ when $X\lesssim Y\lesssim X$). In other words, sufficiently smooth functions are always contained in the shallow neural network space $\mathcal{F}_{\sigma_{k}}(M)$. And, for less smooth functions, we can characterize the approximation error by the variation norm. Furthermore, combining our result with the random approximation bounds from Bach [2017]; Siegel and Xu [2022], we are able to prove that shallow ReLUk neural network of the form (1.1) achieves the optimal approximation rate $\mathcal{O}(N^{-\alpha/d})$ for $\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$, which generalizes the result of [Mhaskar, 1996] to ReLUk.
In addition to shallow neural networks, we can also apply our results to derive approximation bounds for multi-layer neural networks and convolutional neural networks (CNNs) when $k=1$ (ReLU activation $\sigma:=\sigma_{1}$). These approximation bounds can then be used to study the performances of machine learning algorithms based on neural networks [Anthony and Bartlett, 2009]. Here, we illustrate the idea by studying the nonparametric regression problem. The goal is to learn a function $h$ from a hypothesis space $\mathcal{H}$ from its noisy samples
$Y_{i}=h(X_{i})+\eta_{i},\quad i=1,\dots,n,$
where $X_{i}$ is sampled from an unknown probability distribution $\mu$ and $\eta_{i}$ is Gaussian noise. One popular algorithm for solving this problem is the empirical least square minimization
$\operatorname*{argmin}_{f\in\mathcal{F}_{n}}\frac{1}{n}\sum_{i=1}^{n}|f(X_{i})-Y_{i}|^{2},$
where $\mathcal{F}_{n}$ is an appropriately chosen function class. For instance, in deep learning, $\mathcal{F}_{n}$ is parameterized by deep neural networks and one solves the minimization by (stochastic) gradient descent methods. Assuming that we can compute a minimizer $f_{n}^{*}\in\mathcal{F}_{n}$, the performance of the algorithm is often measured by the square loss $\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}$. A fundamental question in learning theory is to determine the convergence rate of the error $\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\to 0$ as the sample size $n\to\infty$. The error can be decomposed into two components: approximation error and generalization error (also called estimation error). For neural network models $\mathcal{F}_{n}$, our results provide bounds for the approximation errors, while the generalization errors can be bounded by the complexity of the models [Shalev-Shwartz and Ben-David, 2014; Mohri et al., 2018]. We study the cases $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$ or $\mathcal{H}=\mathcal{F}_{\sigma}(1)$ for three ReLU neural network models: shallow neural network, over-parameterized neural network, and CNN. The models and our contributions are summarized as follows:
  1. (1)
    Shallow ReLU neural network $\mathcal{F}_{\sigma}(N,M)$, where $N$ is the number of neurons and $M$ is a bound for the variation norm that measures the size of the weights. We prove optimal approximation rates (in terms of $N$) for this model. It is also shown that this model can achieve the optimal convergence rates for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$, which complements the recent results for deep neural networks [Schmidt-Hieber, 2020; Kohler and Langer, 2021].
  2. (2)
    Over-parameterized (deep or shallow) ReLU neural network $\mathcal{NN}(W,L,M)$ studied in [Jiao et al., 2023], where $W,L$ are the width and depth respectively, and $M$ is a constraint on the weight matrices. For fixed depth $L$, the generalization error for this model can be controlled by $M$ [Jiao et al., 2023]. When $\mathcal{H}=\mathcal{H}^{\alpha}$, we characterize the approximation error by $M$, and allow the width $W$ to be arbitrary large so that the model can be over-parameterized (the number of parameters is larger than the number of samples). When $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, we can simply increase the width to reduce the approximation error so that the model can also be over-parameterized. Our result shows that this model can achieve nearly optimal convergence rates for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$. Both the approximation and convergence rates improve the results of [Jiao et al., 2023].
  3. (3)
    Sparse convolutional ReLU neural network $\mathcal{CNN}(s,L)$ introduced by Zhou [2020b], where $L$ is the depth and $s\geq 2$ is a fixed integer that controls the filter length. This model is shown to be universal for approximation [Zhou, 2020b] and universal consistent for regression [Lin et al., 2022]. We improve the approximation bound in [Zhou, 2020b] and give new convergence rates of this model for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$.
The approximation rates and convergence rates of nonparametric regression for these models are summarized in Table 1.1, where we use the notation $a\lor b:=\max\{a,b\}$.
Model Approximation Nonparametric regression
$\mathcal{H}^{\alpha},\alpha<\frac{d+3}{2}$ $\mathcal{F}_{\sigma}(1)$ $\mathcal{H}^{\alpha},\alpha<\frac{d+3}{2}$ $\mathcal{F}_{\sigma}(1)$
$\mathcal{F}_{\sigma}(N,M)$ $N^{-\frac{\alpha}{d}}\lor M^{-\frac{2\alpha}{d+3-2\alpha}}$ $N^{-\frac{1}{2}-\frac{3}{2d}}$ $n^{-\frac{2\alpha}{d+2\alpha}}$ $n^{-\frac{d+3}{2d+3}}$
$\mathcal{NN}(W,L,M)$ $W^{-\frac{\alpha}{d}}\lor M^{-\frac{2\alpha}{d+3-2\alpha}}$ $W^{-\frac{1}{2}-\frac{3}{2d}}$ $n^{-\frac{2\alpha}{d+3+2\alpha}}$ $n^{-\frac{1}{2}}$
$\mathcal{CNN}(s,L)$ $L^{-\frac{\alpha}{d}}$ $L^{-\frac{1}{2}-\frac{3}{2d}}$ $n^{-\frac{\alpha}{d+\alpha}}$ $n^{-\frac{d+3}{3d+3}}$
Table 1.1: Approximation rates and convergence rates of nonparametric regression for three neural network models, ignoring logarithmic factors.
The rest of the paper is organized as follows. In Section 2, we present our approximation results for shallow neural networks. Section 3 gives a proof of our main theorem. In Section 4, we apply our approximation results to study there neural network models and derive convergence rates for nonparametric regression using these models. Section 5 concludes this paper with a discussion on possible future directions of research.
2 Approximation rates for shallow neural networks
Let us begin with some notations for function classes. Let $\mathbb{B}^{d}=\{x\in\mathbb{R}^{d}:\|x\|_{2}\leq 1\}$ and $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:\|x\|_{2}=1\}$ be the unit ball and the unit sphere of $\mathbb{R}^{d}$. We are interested in functions of the integral form
$f(x)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu(v),\quad x\in\mathbb{B}^{d},$ (2.1)
where $\mu$ is a signed Radon measure on $\mathbb{S}^{d}$ with finite total variation $\|\mu\|:=|\mu|(\mathbb{S}^{d})<\infty$ and $\sigma_{k}(t):=\max\{t,0\}^{k}$ with $k\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ is the ReLUk function (when $k=0$, $\sigma_{0}(t)$ is the Heaviside function). For simplicity, we will also denote the ReLU function by $\sigma:=\sigma_{1}$. The variation norm $\gamma(f)$ of $f$ is the infimum of $\|\mu\|$ over all decompositions of $f$ as (2.1) [Bach, 2017]. By the compactness of $\mathbb{S}^{d}$, the infimum can be attained by some signed measure $\mu$. We denote $\mathcal{F}_{\sigma_{k}}(M)$ as the function class that contains all functions $f$ in the form (2.1) whose variation norm $\gamma(f)\leq M$, see (1.2). The class $\mathcal{F}_{\sigma_{k}}(M)$ can be thought of as an infinitely wide neural network with a constraint on its weights. The variation spaces corresponding to shallow neural networks have been studied by many researchers. We refer the reader to [Savarese et al., 2019; Ongie et al., 2020; Bartolucci et al., 2023; Parhi and Nowak, 2022, 2023; Siegel and Xu, 2020, 2022, 2023] for several other definitions and characterizations of these spaces.
We will also need the notion of classical smoothness of functions on Euclidean space. Given a smoothness index $\alpha>0$, we write $\alpha=r+\beta$ where $r\in\mathbb{N}_{0}$ and $\beta\in(0,1]$. Let $C^{r,\beta}(\mathbb{R}^{d})$ be the Hölder space with the norm
$\|f\|_{C^{r,\beta}(\mathbb{R}^{d})}:=\max\left\{\|f\|_{C^{r}(\mathbb{R}^{d})},\max_{\|s\|_{1}=r}|\partial^{s}f|_{C^{0,\beta}(\mathbb{R}^{d})}\right\},$
where $s=(s_{1},\dots,s_{d})\in\mathbb{N}_{0}^{d}$ is an multi-index and
$\|f\|_{C^{r}(\mathbb{R}^{d})}:=\max_{\|s\|_{1}\leq r}\|\partial^{s}f\|_{L^{\infty}(\mathbb{R}^{d})},\quad|f|_{C^{0,\beta}(\mathbb{R}^{d})}:=\sup_{x\neq y\in\mathbb{R}^{d}}\frac{|f(x)-f(y)|}{\|x-y\|_{2}^{\beta}},$
where we use $\|\cdot\|_{L^{\infty}}$ to denote the supremum norm, since we mainly consider continuous functions. We write $C^{r,\beta}(\mathbb{B}^{d})$ for the Banach space of all restrictions to $\mathbb{B}^{d}$ of functions in $C^{r,\beta}(\mathbb{R}^{d})$. The norm is given by $\|f\|_{C^{r,\beta}(\mathbb{B}^{d})}=\inf\{\|g\|_{C^{r,\beta}(\mathbb{R}^{d})}:g\in C^{r,\beta}(\mathbb{R}^{d})\mbox{ and }g=f\mbox{ on }\mathbb{B}^{d}\}$. For convenience, we will denote the unit ball of $C^{r,\beta}(\mathbb{B}^{d})$ by $\mathcal{H}^{\alpha}:=\{f\in C^{r,\beta}(\mathbb{B}^{d}):\|f\|_{C^{r,\beta}(\mathbb{B}^{d})}\leq 1\}$. Note that, for $\alpha=1$, $\mathcal{H}^{\alpha}$ is Lipschitz continuous function class.
Due to the universality of shallow neural networks [Pinkus, 1999], $\mathcal{F}_{\sigma_{k}}(M)$ can approximate any continuous functions on $\mathbb{B}^{d}$ if $M$ is sufficiently large. Our main theorem estimates the rate of this approximation for Hölder class.
Theorem 2.1.
Let $k\in\mathbb{N}_{0}$, $d\in\mathbb{N}$ and $\alpha>0$. If $\alpha>(d+2k+1)/2$ or $\alpha=(d+2k+1)/2$ is an even integer, then $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma_{k}}(M)$ for some constant $M$ depending on $k,d,\alpha$. Otherwise,
$\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim\begin{cases}\exp(-\alpha M^{2}),&\mbox{ if }\alpha=(d+2k+1)/2\mbox{ and }\alpha/2\notin\mathbb{N},\\ M^{-\frac{2\alpha}{d+2k+1-2\alpha}},&\mbox{ if }\alpha<(d+2k+1)/2,\end{cases}$
where the implied constants only depend on $k,d,\alpha$.
The proof of Theorem 2.1 is deferred to the next section. Our proof uses similar ideas as Bach [2017, Proposition 3], which obtained the same approximation rate for $\alpha=1$. The conclusion is more complicated for the critical value $\alpha=(d+2k+1)/2$. We think this is due to the proof technique and conjecture that $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma_{k}}(M)$ for all $\alpha\geq(d+2k+1)/2$, see Remark 3.4. Nevertheless, in practical applications of machine learning, the dimension $d$ is large and it is reasonable to expect that $\alpha<(d+2k+1)/2$.
In order to apply Theorem 2.1 to shallow neural networks with finite neurons, we can approximate $\mathcal{F}_{\sigma_{k}}(M)$ by the subclass
$\mathcal{F}_{\sigma_{k}}(N,M):=\left\{f(x)=\sum_{i=1}^{N}a_{i}\sigma_{k}((x^{\intercal},1)v_{i}):v_{i}\in\mathbb{S}^{d},\sum_{i=1}^{N}|a_{i}|\leq M\right\},$
where we restrict the measure $\mu$ to be a discrete one supported on at most $N$ points. The next proposition shows that any function in $\mathcal{F}_{\sigma_{k}}(M)$ is the limit of functions in $\mathcal{F}_{\sigma_{k}}(N,M)$ as $N\to\infty$.
Proposition 2.2.
Let $\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$ denote the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$ in $L^{\infty}(\mathbb{B}^{d})$. If $k\in\mathbb{N}$, then $\mathcal{F}_{\sigma_{k}}(1)=\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$. If $k=0$, then $\mathcal{F}_{\sigma_{0}}(1)\subseteq\widetilde{\mathcal{F}}_{\sigma_{0}}(1)$.
Proof.
We first show that $\mathcal{F}_{\sigma_{k}}(1)\subseteq\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$ for $k\in\mathbb{N}_{0}$. For any $f\in\mathcal{F}_{\sigma_{k}}(1)$ with the integral form $f(x)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu(v)$, we can decompose $f$ as
$f(x)$ $=\|\mu_{+}\|\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)\frac{d\mu_{+}(v)}{\|\mu_{+}\|}-\|\mu_{-}\|\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)\frac{d\mu_{+}(v)}{\|\mu_{-}\|}$
$=:\|\mu_{+}\|f_{+}(x)-\|\mu_{-}\|f_{-}(x),$
where $\mu_{+}$ and $\mu_{-}$ are the positive and negative parts of $\mu$. If $f_{+},f_{-}\in\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$, then $f\in\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$. Hence, without loss of generality, we can assume $\mu$ is a probability measure. We are going to approximate $f$ by uniform laws of large numbers. Let $\{v_{i}\}_{i=1}^{N}$ be $N$ i.i.d. samples from $\mu$. By symmetrization argument (see Wainwright [2019, Theorem 4.10] for example), we can bound the expected approximation error by Rademacher complexity [Bartlett and Mendelson, 2002]:
$\mathbb{E}\left[\sup_{x\in\mathbb{B}^{d}}\left|f(x)-\frac{1}{N}\sum_{i=1}^{N}\sigma_{k}((x^{\intercal},1)v_{i})\right|\right]\leq 2\mathbb{E}\left[\sup_{x\in\mathbb{B}^{d}}\left|\frac{1}{N}\sum_{i=1}^{N}\epsilon_{i}\sigma_{k}((x^{\intercal},1)v_{i})\right|\right]=:\mathcal{E}_{k}(N),$
where $(\epsilon_{1},\dots,\epsilon_{N})$ is an i.i.d. sequence of Rademacher random variables. For $k\in\mathbb{N}$, the Lipschitz constant of $\sigma_{k}$ on $\sqrt{2},\sqrt{2}]$ is $k2^{(k-1)/2}$. By the contraction property of Rademacher complexity [Ledoux and Talagrand, 1991, Corollary 3.17],
$\mathcal{E}_{k}(N)$ $\leq\frac{k2^{(k+1)/2}}{N}\mathbb{E}\left[\sup_{x\in\mathbb{B}^{d}}\left|\sum_{i=1}^{N}\epsilon_{i}(x^{\intercal},1)v_{i}\right|\right]\leq\frac{k2^{k/2+1}}{N}\mathbb{E}\left[\left\|\sum_{i=1}^{N}\epsilon_{i}v_{i}\right\|_{2}\right]$
$\leq\frac{k2^{k/2+1}}{N}\sqrt{\mathbb{E}\left[\left\|\sum_{i=1}^{N}\epsilon_{i}v_{i}\right\|_{2}^{2}\right]}=\frac{k2^{k/2+1}}{N}\sqrt{\mathbb{E}\left[\sum_{i=1}^{N}\left\|v_{i}\right\|_{2}^{2}\right]}=\frac{k2^{k/2+1}}{\sqrt{N}}.$
For $k=0$, the VC dimension of the function class $\{f_{x}(v)=\sigma_{0}((x^{\intercal},1)v):x\in\mathbb{B}^{d}\}$ is at most $d$ [Wainwright, 2019, Proposition 4.20]. Thus, we have the bound $\mathcal{E}_{0}(N)\lesssim\sqrt{d/N}$ by Wainwright [2019, Example 5.24]. Hence, $f$ is in the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$.
Next, we show that $\widetilde{\mathcal{F}}_{\sigma_{k}}(1)\subseteq\mathcal{F}_{\sigma_{k}}(1)$ for $k\in\mathbb{N}$. Since $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)\subseteq\mathcal{F}_{\sigma_{k}}(1)$, we only need to show that $\mathcal{F}_{\sigma_{k}}(1)$ is closed in $L^{\infty}(\mathbb{B}^{d})$. Let $f_{n}(x)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu_{n}(v)$, where $\|\mu_{n}\|\leq 1$, be a convergent sequence with limit $f\in L^{\infty}(\mathbb{B}^{d})$. By the compactness of $\mathbb{S}^{d}$ and Prokhorov’s theorem, there exists a weakly convergent subsequence $\mu_{n_{i}}\to\mu$. In particular, $\|\mu\|\leq 1$ and for any $x\in\mathbb{B}^{d}$,
$\lim_{n_{i}\to\infty}\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu_{n_{i}}(v)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu(v)=:\widetilde{f}$.
By the compactness of $\mathbb{B}^{d}$, $f_{n_{i}}$ convergence uniformly to $\widetilde{f}$. Hence $f=\widetilde{f}\in\mathcal{F}_{\sigma_{k}}(1)$ and $\mathcal{F}_{\sigma_{k}}(1)$ is closed in $L^{\infty}(\mathbb{B}^{d})$. ∎
The proof of Proposition 2.2 actually shows the approximation rate $\mathcal{O}(N^{-1/2})$ for the subclass $\mathcal{F}_{\sigma_{k}}(N,1)$. This rate can be improved if we take into account the smoothness of the activation function.
Lemma 2.3 ([Bach, 2017; Ma et al., 2022; Siegel and Xu, 2022]).
Let $2\leq p<\infty$, then there exists an $M\geq 1$ depending only on $k,p,d$ such that,
$\sup_{f\in\mathcal{F}_{\sigma_{k}}(1)}\inf_{f_{N}\in\mathcal{F}_{\sigma_{k}}(N,M)}\|f-f_{N}\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{pk+1}{pd}}.$
For $k\in\{0,1\}$ and $d>2k+1$, it holds that
$\sup_{f\in\mathcal{F}_{\sigma}(1)}\inf_{f_{N}\in\mathcal{F}_{\sigma}(N,M)}\|f-f_{N}\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{2k+1}{2d}}.$
Proof.
Using similar argument as Siegel and Xu [2022, Lemma 1], one can show that the parameterization map $\mathcal{P}_{\sigma_{k}}:\mathbb{S}^{d}\to L^{p}(\mathbb{B}^{d})$ given by $\mathcal{P}_{\sigma_{k}}(v)=\sigma_{k}((x^{\intercal},1)v)$ is of smoothness class $\alpha=k+1/p$ in the sense that, for each coordinate chart $(V,\phi)$ and $\xi\in L^{q}(\mathbb{B}^{d})$ where $1/p+1/q=1$, the function $f_{\xi}(u)=\langle\xi,\mathcal{P}_{\sigma_{k}}(\phi^{-1}(u))\rangle$ satisfies $\|f_{\xi}\|_{C^{r,\beta}(\phi(V))}\leq C\|\xi\|_{L^{q}(\mathbb{B}^{d})}$ for some constant $C<\infty$. By Siegel and Xu [2022, Theorem 2], if the parameterization map $\mathcal{P}_{\sigma_{k}}$ is of smoothness class $\alpha$, then there exists an $M\geq 1$ such that for any $f$ in the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$ in $L^{p}(\mathbb{B}^{d})$, we have
$\inf_{f_{N}\in\mathcal{F}_{\sigma_{k}}(N,M)}\|f-f_{N}\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{\alpha}{d}}.$
By Proposition 2.2, we can apply this bound to $f\in\mathcal{F}_{\sigma_{k}}(1)$.
The uniform bound for $k=0$ is obtained by [Ma et al., 2022, Theorem 4]. For $k=1$, the uniform bound is proved by Bach [2017, Proposition 1], who showed that approximating $f\in\mathcal{F}_{\sigma}(1)$ by finitely many neurons is essentially equivalent to the approximation of a zonoid by zonotopes [Matoušek, 1996]. The condition $d\geq 4$ is needed in order to use [Matoušek, 1996, Theorem 1]. We note that the proof also implies that one can take $M=3/2$. ∎
One can show that the rate $\mathcal{O}(N^{-\frac{1}{2}-\frac{2k+1}{2d}})$ in Lemma 2.3 is sharp for $p=2$ [Siegel and Xu, 2022]. Since the same rate holds in the supremum norm when $k=0,1$, we expect that it also holds for all $k\in\mathbb{N}_{0}$ and $2\leq p\leq\infty$. It seems that one needs to use techniques of geometric discrepancy theory to prove such kind of results [Bourgain et al., 1989; Matoušek, 1996].
Combining Theorem 2.1 and Lemma 2.3, we can derive approximation rate of Hölder class $\mathcal{H}^{\alpha}$ for shallow neural network $\mathcal{F}_{\sigma_{k}}(N,M)$. Recall that we use the notation $a\lor b:=\max\{a,b\}$.
Corollary 2.4.
Let $k\in\mathbb{N}_{0}$, $d\in\mathbb{N}$, $2\leq p<\infty$ and $\alpha>0$.
  1. (1)
    If $\alpha>(d+2k+1)/2$ or $\alpha=(d+2k+1)/2$ is an even integer, then there exists a constant $M$ depending on $k,p,d,\alpha$ such that
    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{pk+1}{pd}}.$
  2. (2)
    If $\alpha=(d+2k+1)/2$ is not an even integer, then there exists $M\asymp\sqrt{\log N}$ such that
    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{pk+1}{pd}}\sqrt{\log N}.$
  3. (3)
    If $\alpha<(d+2k+1)/2$, then
    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{\alpha}{d}\frac{d+2k+2/p}{d+2k+1}}\lor M^{-\frac{2\alpha}{d+2k+1-2\alpha}}.$
  4. (4)
    If $k\in\{0,1\}$ and $d>2k+1$, the bounds for $p=2$ also hold in the supremum norm. In particular, for $\alpha<(d+2k+1)/2$, we have
    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{\alpha}{d}}\lor M^{-\frac{2\alpha}{d+2k+1-2\alpha}}.$
    Thus, the rate $\mathcal{O}(N^{-\alpha/d})$ holds when $M\gtrsim N^{-(d+2k+1-2\alpha)/(2d)}$.
Proof.
We only present the proof for part (3), since other parts can be derived similarly. If $\alpha<(d+2k+1)/2$, then by Theorem 2.1, for any $h\in\mathcal{H}^{\alpha}$, there exists $g\in\mathcal{F}_{\sigma_{k}}(K)$ such that $\|h-g\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}$. Let $M_{0}$ be the constant in Lemma 2.3, then there exists $f\in\mathcal{F}_{\sigma_{k}}(N,KM_{0})$ such that $\|g-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim KN^{-\frac{1}{2}-\frac{pk+1}{pd}}$. If $M\geq N^{\frac{d+2k+1-2\alpha}{2d}\frac{d+2k+2/p}{d+2k+1}}$, we choose $K=N^{\frac{d+2k+1-2\alpha}{2d}\frac{d+2k+2/p}{d+2k+1}}/M_{0}$ then $f\in\mathcal{F}_{\sigma_{k}}(N,KM_{0})\subseteq\mathcal{F}_{\sigma_{k}}(N,M)$ and
$\|h-f\|_{L^{p}(\mathbb{B}^{d})}$ $\leq\|h-g\|_{L^{p}(\mathbb{B}^{d})}+\|g-f\|_{L^{p}(\mathbb{B}^{d})}$
$\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}+KN^{-\frac{d+2k+2/p}{2d}}$
$\lesssim N^{-\frac{\alpha}{d}\frac{d+2k+2/p}{d+2k+1}}.$
If $M\leq N^{\frac{d+2k+1-2\alpha}{2d}\frac{d+2k+2/p}{d+2k+1}}$, we choose $K=M/M_{0}$, then
$\|h-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}+KN^{-\frac{d+2k+2/p}{2d}}\lesssim M^{-\frac{2\alpha}{d+2k+1-2\alpha}}.$
Combining the two bounds gives the desired result. By Lemma 2.3, if $k=0,1$ and $d>2k+1$, the bounds for $p=2$ also hold in the supremum norm. ∎
We make some comments on the approximation rate for $\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$. As shown by Pinkus [1999, Corollary 6.10], the rate $\mathcal{O}(N^{-\alpha/d})$ in the $L^{2}$ norm is already known for $\alpha=1,2,\dots,(d+2k+1)/2$. For ReLU activation, the recent paper [Mao and Zhou, 2023] obtained the rate $\mathcal{O}(N^{-\frac{\alpha}{d}\frac{d+2}{d+4}})$ in the supremum norm. Corollary 2.4 shows that the rate $\mathcal{O}(N^{-\alpha/d})$ also holds in the supremum norm for ReLU and Heaviside activations. And more importantly, we also provide an explicit control on the network weights to ensure that such rate can be achieved, which is useful for estimating generalization errors (see Section 4). It is well-known that the optimal approximation rate for $\mathcal{H}^{\alpha}$ is $\mathcal{O}(N^{-\alpha/d})$, if we approximate $h\in\mathcal{H}^{\alpha}$ by a function class with $N$ parameters and the parameters are continuously depend on the target function $h$ [DeVore et al., 1989]. However, this result is not exactly applicable to neural networks, because we do not have guarantee that the parameters in the network depend continuously on the target function (in fact, this is not true for some constructions [Yarotsky, 2018; Lu et al., 2021; Yang et al., 2022]). Nevertheless, one can still prove that the rate $\mathcal{O}(N^{-\alpha/d})$ is optimal for shallow ReLUk neural networks by arguments based on pseudo-dimension as done in [Yarotsky, 2017; Lu et al., 2021; Yang et al., 2022].
We describe the idea of proving approximation lower bound through pseudo-dimension by reviewing the result of Maiorov and Ratsaby [1999] (see also [Achour et al., 2022]). Recall that the pseudo-dimension ${\rm Pdim}\,(\mathcal{F})$ of a real-valued function class $\mathcal{F}$ defined on $\mathbb{B}^{d}$ is the largest integer $n$ for which there exist points $x_{1},\dots,x_{n}\in\mathbb{B}^{d}$ and constants $c_{1},\dots,c_{n}\in\mathbb{R}$ such that
$\left|\{\,{\rm sgn}\,(f(x_{1})-c_{1}),\dots,\,{\rm sgn}\,(f(x_{n})-c_{n}):f\in\mathcal{F}\}\right|=2^{n}.$ (2.2)
Maiorov and Ratsaby [1999] introduced a nonlinear $n$-width defined as
$\rho_{n}(\mathcal{H}^{\alpha})=\inf_{\mathcal{F}_{n}}\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{n}}\|h-f\|_{L^{p}(\mathbb{B}^{d})},$
where $\mathcal{F}_{n}$ runs over all the classes in $L^{p}(\mathbb{B}^{d})$ with ${\rm Pdim}\,(\mathcal{F}_{n})\leq n$. They constructed a well-separated subclass of $\mathcal{H}^{\alpha}$ such that if a function class $\mathcal{F}$ can approximate this subclass with small error, then ${\rm Pdim}\,(\mathcal{F})$ should be large. In other words, the approximation error of any class $\mathcal{F}_{n}$ with ${\rm Pdim}\,(\mathcal{F}_{n})\leq n$ can be lower bounded. Consequently, they proved that
$\rho_{n}(\mathcal{H}^{\alpha})\gtrsim n^{-\alpha/d}.$
By [Bartlett et al., 2019], we can upper bound the pseudo-dimension of shallow ReLUk neural networks as $n:=\,{\rm Pdim}\,(\mathcal{F}_{\sigma_{k}}(N,M))\lesssim N\log N$. Hence,
$\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f_{N}\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f_{N}\|_{L^{p}(\mathbb{B}^{d})}\geq\rho_{n}(\mathcal{H}^{\alpha})\gtrsim(N\log N)^{-\alpha/d},$
which shows that the rate $\mathcal{O}(N^{-\alpha/d})$ in Corollary 2.4 is optimal in the $L^{2}$ norm (ignoring logarithmic factors). This also implies the optimality of Theorem 2.1 (otherwise, the proof of Corollary 2.4 would give a rate better than $\mathcal{O}(N^{-\alpha/d})$).
3 Proof of Theorem 2.1
Following the idea of Bach [2017], we first transfer the problem to approximation on spheres. Let us begin with a brief review of harmonic analysis on spheres [Dai and Xu, 2013]. For $n\in\mathbb{N}_{0}$, the spherical harmonic space $\mathbb{Y}_{n}$ of degree $n$ is the linear space that contains the restrictions of real harmonic homogeneous polynomials of degree $n$ on $\mathbb{R}^{d+1}$ to the sphere $\mathbb{S}^{d}$. The dimension of $\mathbb{Y}_{n}$ is $N(d,n):=\frac{2n+d-1}{n}\binom{n+d-2}{d-1}$ if $n\neq 0$ and $N(d,n):=1$ if $n=0$. Spherical harmonics are eigenfunctions of the Laplace-Beltrami operator:
$\Delta Y_{n}=-n(n+d-1)Y_{n},\quad Y_{n}\in\mathbb{Y}_{n},$
where, in the coordinates $u=(u_{1},\dots,u_{d+1})\in\mathbb{S}^{d}$,
$\Delta=\sum_{i=1}^{d}\frac{\partial^{2}}{\partial u_{i}^{2}}-\sum_{i=1}^{d}\sum_{j=1}^{d}u_{i}u_{j}\frac{\partial^{2}}{\partial u_{i}\partial u_{j}}-d\sum_{i=1}^{d}u_{i}\frac{\partial}{\partial u_{i}}.$
Spherical harmonics of different degrees are orthogonal with respect to the inner product $\langle f,g\rangle=\int_{\mathbb{S}^{d}}f(u)g(u)d\tau_{d}(u)$, where $\tau_{d}$ is the surface area measure of $\mathbb{S}^{d}$ (normalized by the surface area $\omega_{d}:=2\pi^{(d+1)/2}/\Gamma((d+1)/2)$ so that $\tau_{d}(\mathbb{S}^{d})=1$).
Let $\mathcal{P}_{n}:L^{2}(\mathbb{S}^{d})\to\mathbb{Y}_{n}$ denote the orthogonal projection operator. For any orthonormal basis $\{Y_{nj}:1\leq j\leq N(d,n)\}$ of $\mathbb{Y}_{n}$, the addition formula [Dai and Xu, 2013, Theorem 1.2.6] shows
$\sum_{j=1}^{N(d,n)}Y_{nj}(u)Y_{nj}(v)=N(d,n)P_{n}(u^{\intercal}v),\quad u,v\in\mathbb{S}^{d},$ (3.1)
where $P_{n}$ is the Gegenbauer polynomial
$P_{n}(t):=\frac{(-1)^{n}}{2^{n}}\frac{\Gamma(d/2)}{\Gamma(n+d/2)}(1-t^{2})^{(2-d)/2}\left(\frac{d}{dt}\right)^{n}(1-t^{2})^{n+(d-2)/2},\quad t\in[-1,1],$
with normalization $P_{n}(1)=1$. Applying the Cauchy–Schwarz inequality to (3.1), we get $|P_{n}(t)|\leq 1$. For $n\neq 0$, $P_{n}(t)$ is odd (even) if $n$ is odd (even). Note that. for $d=1$ and $n\neq 0$, $N(d,n)=2$ and $P_{n}(t)$ is the Chebyshev polynomial such that $P_{n}(\cos\theta)=\cos(n\theta)$. We can write the projection $\mathcal{P}_{n}$ as
$\mathcal{P}_{n}f(u)=N(d,n)\int_{\mathbb{S}^{d}}f(v)P_{n}(u^{\intercal}v)d\tau_{d}(v).$
This motivates the following definition of a convolution operator on the sphere.
Definition 3.1 (Convolution).
Let $\varrho$ be the probability distribution with density $c_{d}(1-t^{2})^{(d-2)/2}$ on [1,1][-1,1], where the constant $c_{d}=(\int_{-1}^{1}(1-t^{2})^{(d-2)/2}dt)^{-1}=\omega_{d-1}/\omega_{d}$. For $f\in L^{1}(\mathbb{S}^{d})$ and $g\in L^{1}_{\varrho}([-1,1])$, define
$(f*g)(u):=\int_{\mathbb{S}^{d}}f(v)g(u^{\intercal}v)d\tau_{d}(v),\quad u\in\mathbb{S}^{d}.$
The convolution on the sphere satisfies Young’s inequality [Dai and Xu, 2013, Theorem 2.1.2]: for $p,q,r\geq 1$ with $p^{-1}=q^{-1}+r^{-1}-1$, it holds
$\|f*g\|_{L^{p}(\mathbb{S}^{d})}\leq\|f\|_{L^{q}(\mathbb{S}^{d})}\|g\|_{L^{r}_{\varrho}([-1,1])},$
where the norm is the uniform norm when $r=\infty$. Observe that the projection $\mathcal{P}_{n}f=f*(N(d,n)P_{n})$ is a convolution operator with $\|N(d,n)P_{n}\|_{L^{\infty}([-1,1])}\leq N(d,n)$. Furthermore, for $g\in L^{1}_{\varrho}([-1,1])$, let $\widehat{g}(n)$ denote the Fourier coefficient of $g$ with respect to the Gegenbauer polynomials,
$\widehat{g}(n):=\frac{\omega_{d-1}}{\omega_{d}}\int_{-1}^{1}g(t)P_{n}(t)(1-t^{2})^{(d-2)/2}dt.$
By the Funk-Hecke formula, one can show that [Dai and Xu, 2013, Theorem 2.1.3]
$\mathcal{P}_{n}(f*g)=\widehat{g}(n)\mathcal{P}_{n}f,\quad f\in L^{1}(\mathbb{S}^{d}),n\in\mathbb{N}_{0}.$ (3.2)
This identity is analogous to the Fourier transform of ordinary convolution.
One of the key steps in our proof of Theorem 2.1 is the observation that functions of the form $f(u)=\int_{\mathbb{S}^{d}}\phi(v)\sigma_{k}(u^{\intercal}v)d\tau_{d}(v)$ are convolutions $\phi*\sigma_{k}$ with the activation function $\sigma_{k}\in L^{\infty}([-1,1])$. Bach [2017, Appendix D.2] has computed the Fourier coefficients $\widehat{\sigma_{k}}(n)$ explicitly. We summarize the result in the following.
Proposition 3.2.
For $k\in\mathbb{N}_{0}$, $\widehat{\sigma_{k}}(n)=0$ if and only if $n\geq k+1$ and $n\equiv k\bmod 2$. If $n=0$,
$\widehat{\sigma_{k}}(0)=\frac{\omega_{d-1}}{\omega_{d}}\frac{\Gamma(d/2)\Gamma((k+1)/2)}{2\Gamma((k+d+1)/2)}.$
If $n\geq k+1$ and $n+1\equiv k\bmod 2$,
$\widehat{\sigma_{k}}(n)=\frac{\omega_{d-1}}{\omega_{d}}\frac{k!(-1)^{(n-k-1)/2}}{2^{n}}\frac{\Gamma(d/2)\Gamma(n-k)}{\Gamma((n-k+1)/2)\Gamma((n+d+k+1)/2)}.$
By the Stirling formula $\Gamma(x)=\sqrt{2\pi}x^{x-1/2}e^{-x}(1+\mathcal{O}(x^{-1}))$, we have $\widehat{\sigma_{k}}(n)\asymp n^{-(d+2k+1)/2}$ for $n\in\mathbb{N}$ satisfying $\widehat{\sigma_{k}}(n)\neq 0$.
Next, we introduce the smoothness of functions on the sphere. For $0\leq\theta\leq\pi$, the translation operator $T_{\theta}$, also called spherical mean operator, is defined by
$T_{\theta}f(u):=\int_{\mathbb{S}^{\perp}_{u}}f(u\cos\theta+v\sin\theta)d\tau_{d-1}(v),\quad u\in\mathbb{S}^{d},f\in L^{1}(\mathbb{S}^{d}),$
where $\mathbb{S}^{\perp}_{u}:=\{v\in\mathbb{S}^{d}:u^{\intercal}v=0\}$ is the equator in $\mathbb{S}^{d}$ with respect to $u$ (hence $\mathbb{S}^{\perp}_{u}$ is isomorphic to the sphere $\mathbb{S}^{d-1}$). We note that the translation operator satisfies $\mathcal{P}_{n}(T_{\theta}f)=P_{n}(\cos\theta)\mathcal{P}_{n}(f)$. For $\alpha>0$ and $0<\theta<\pi$, we define the $\alpha$-th order difference operator
$\Delta_{\theta}^{\alpha}:=(I-T_{\theta})^{\alpha/2}=\sum_{j=0}(-1)^{j}\binom{\alpha/2}{j}T_{\theta}^{j},$
where $\binom{\alpha}{j}=\frac{\alpha(\alpha-1)\cdots(\alpha-j+1)}{j!}$, in a distributional sense by $\mathcal{P}_{n}(\Delta_{\theta}^{\alpha}f)=(1-P_{n}(\cos\theta))^{\alpha/2}\mathcal{P}_{n}f$, $n\in\mathbb{N}_{0}$. For $f\in L^{p}(\mathbb{S}^{d})$ and $1\leq p<\infty$ or $f\in C(\mathbb{S}^{d})$ and $p=\infty$, the $\alpha$-th order modulus of smoothness is defined by
$\omega_{\alpha}(f,t)_{p}:=\sup_{0<\theta\leq t}\|\Delta_{\theta}^{\alpha}f\|_{L^{p}(\mathbb{S}^{d})},\quad 0<t<\pi.$
For even integers $\alpha=2s$, one can also use combinations of $T_{j\theta}$ and obtain [Rustamov, 1993; Ditzian, 2006]
$\omega_{2s}(f,t)_{p}\asymp\sup_{0<\theta\leq t}\left\|\sum_{j=0}^{2s}(-1)^{j}\binom{2s}{j}T_{j\theta}f\right\|_{L^{p}(\mathbb{S}^{d})},\quad s\in\mathbb{N}.$ (3.3)
Another way to characterize the smoothness is through the $K$-functional. We first introduce the fractional Sobolev space induced by the Laplace-Beltrami operator. We say a function $f\in L^{p}(\mathbb{S}^{d})$ belong to the Sobolev space $\mathcal{W}^{\alpha,p}(\mathbb{S}^{d})$ if there exists a function in $L^{p}(\mathbb{S}^{d})$, which will be denoted by $(-\Delta)^{\alpha/2}f$, such that
$\mathcal{P}_{n}((-\Delta)^{\alpha/2}f)=(n(n+d-1))^{\alpha/2}\mathcal{P}_{n}f,\quad n\in\mathbb{N}_{0},$
where we assume $f,(-\Delta)^{\alpha/2}f\in C(\mathbb{S}^{d})$ for $p=\infty$. Then we can define the $\alpha$-th $K$-functional of $f\in L^{p}(\mathbb{S}^{d})$ as
$K_{\alpha}(f,t)_{p}:=\inf_{g\in\mathcal{W}^{\alpha,p}(\mathbb{S}^{d})}\left\{\|f-g\|_{L^{p}(\mathbb{S}^{d})}+t^{\alpha}\|(-\Delta)^{\alpha/2}g\|_{L^{p}(\mathbb{S}^{d})}\right\}.$
It can be shown [Dai and Xu, 2013, Theorem 10.4.1] that the moduli of smoothness and the $K$-functional are equivalent:
$\omega_{\alpha}(f,t)_{p}\asymp K_{\alpha}(f,t)_{p}.$ (3.4)
To prove Theorem 2.1, we denote the function class
$\mathcal{G}_{\sigma_{k}}(M):=\left\{g\in L^{\infty}(\mathbb{S}^{d}):g(u)=\int_{\mathbb{S}^{d}}\sigma_{k}(u^{\intercal}v)d\mu(v),\|\mu\|\leq M\right\},$
as the corresponding function class of $\mathcal{F}_{\sigma_{k}}(M)$ on $\mathbb{S}^{d}$. Abusing the notation, we will also denote $\gamma(g)=\inf_{\mu}\|\mu\|$ as the variation norm of $g\in\mathcal{G}_{\sigma_{k}}(M)$. The next proposition transfer our approximation problem on the unit ball $\mathbb{B}^{d}$ to approximation on the sphere $\mathbb{S}^{d}$.
Proposition 3.3.
Let $k\in\mathbb{N}_{0}$, $d\in\mathbb{N}$ and $\alpha=r+\beta$ where $r\in\mathbb{N}_{0}$ and $\beta\in(0,1]$. Denote $\Omega:=\{(u_{1},\dots,u_{d+1})^{\intercal}\in\mathbb{S}^{d}:u_{d+1}\geq 1/\sqrt{2}\}$ and define an operator $S_{k}:L^{\infty}(\Omega)\to L^{\infty}(\mathbb{B}^{d})$ by
$S_{k}g(x):=(\|x\|_{2}^{2}+1)^{k/2}g\left(\frac{1}{\sqrt{\|x\|^{2}+1}}\begin{pmatrix}x\\ 1\end{pmatrix}\right),\quad x\in\mathbb{B}^{d}.$
The operator $S_{k}$ satisfies: (1) If $g\in\mathcal{G}_{\sigma_{k}}(M)$, then $S_{k}g\in\mathcal{F}_{\sigma_{k}}(M)$. (2) For any $h\in\mathcal{H}^{\alpha}$, there exists $\widetilde{h}\in C(\mathbb{S}^{d})$ such that $S_{k}\widetilde{h}=h$, $\|\widetilde{h}\|_{L^{\infty}(\mathbb{S}^{d})}\leq C$ and $\omega_{2s^{*}}(\widetilde{h},t)_{\infty}\leq Ct^{\alpha}$, where $s^{*}\in\mathbb{N}$ is the smallest integer such that $\alpha\leq 2s^{*}$ and $C$ is a constant independent of $h$. Furthermore, $\widetilde{h}$ can be chosen to be odd or even.
Proof.
(1) If $g(u)=\int_{\mathbb{S}^{d}}\sigma_{k}(u^{\intercal}v)d\mu(v)$ for some $\|\mu\|\leq M$, then for $x\in\mathbb{B}^{d}$,
$S_{k}g(x)$ $=(\|x\|_{2}^{2}+1)^{k/2}\int_{\mathbb{S}^{d}}\sigma_{k}\left((\|x\|_{2}^{2}+1)^{-1/2}(x^{\intercal},1)v\right)d\mu(v)$
$=\int_{\mathbb{S}^{d}}\sigma_{k}\left((x^{\intercal},1)v\right)d\mu(v).$
Hence, $S_{k}g\in\mathcal{F}_{\sigma_{k}}(M)$ by definition.
(2) Given $h\in\mathcal{H}^{\alpha}$, for any $u=(u_{1},\dots,u_{d+1})^{\intercal}\in\Omega$, we define $\widetilde{h}(u):=u_{d+1}^{k}h(u_{d+1}^{-1}u^{\prime})$, where $u^{\prime}=(u_{1},\dots,u_{d})^{\intercal}$. It is easy to check that $S_{k}\widetilde{h}=h$. Note that $h$ is completely determined by the function values of $\widetilde{h}$ on $\Omega$. Observe that the smoothness of $\widetilde{h}$ on $\Omega$ can be controlled by the smoothness of $h$. We can extend $\widetilde{h}$ to $\mathbb{R}^{d+1}$ so that $\|\widetilde{h}\|_{C^{r,\beta}(\mathbb{R}^{d+1})}\leq C_{0}$ for some constant $C_{0}$ independent of $h$, by using (refined version of) Whitney’s extension theorem [Fefferman, 2006, 2009; Fefferman and Israel, 2020]. It remains to show that $\omega_{2s^{*}}(\widetilde{h},t)_{\infty}\lesssim t^{\alpha}$.
By the equivalence (3.3),
$\omega_{2s^{*}}(\widetilde{h},t)_{p}$ $\lesssim\sup_{0<\theta\leq t}\sup_{u\in\mathbb{S}^{d}}\left|\sum_{j=0}^{2s^{*}}(-1)^{j}\binom{2s^{*}}{j}\int_{\mathbb{S}^{\perp}_{u}}\widetilde{h}(u\cos j\theta+v\sin j\theta)d\tau_{d-1}(v)\right|$
$\leq\sup_{0<\theta\leq t}\sup_{u\in\mathbb{S}^{d}}\sup_{v\in\mathbb{S}^{\perp}_{u}}\left|\sum_{j=0}^{2s^{*}}(-1)^{j}\binom{2s^{*}}{j}\widetilde{h}(u\cos j\theta+v\sin j\theta)\right|$
$=:\sup_{0<\theta\leq t}\sup_{u\in\mathbb{S}^{d}}\sup_{v\in\mathbb{S}^{\perp}_{u}}|H(u,v,\theta)|.$
Next, we estimate $\sup_{0<\theta\leq t}|H(u,v,\theta)|$ for small $t>0$ and fixed $u,v$. One can check that the function $f(\cdot):=\widetilde{h}(u\cos(\cdot)+v\sin(\cdot))$ is in $C^{r,\beta}([0,t_{0}])$ for small $t_{0}>0$, and $\|f\|_{C^{r,\beta}([0,t_{0}])}\lesssim\|\widetilde{h}\|_{C^{r,\beta}(\mathbb{R}^{d+1})}$. Let $\widetilde{\Delta}_{\theta}f(\cdot):=f(\cdot+\theta)-f(\cdot)$ be the difference operator and $\widetilde{\Delta}_{\theta}^{n+1}:=\widetilde{\Delta}_{\theta}\widetilde{\Delta}_{\theta}^{n}$ for $n\in\mathbb{N}$. The binomial theorem shows $H(u,v,\theta)=\widetilde{\Delta}_{\theta}^{2s^{*}}f(0)$. Then, classical theory of moduli of smoothness [DeVore and Lorentz, 1993, Chapter 2.6–2.9] implies
$\sup_{0<\theta\leq t}|H(u,v,\theta)|\leq\sup_{0<\theta\leq t}|\widetilde{\Delta}_{\theta}^{2s^{*}}f(0)|\lesssim\|f\|_{C^{r,\beta}([0,t_{0}])}t^{\alpha}.$
Consequently, we get the desired bound $\omega_{2s^{*}}(\widetilde{h},t)_{\infty}\lesssim t^{\alpha}$.
Finally, in order to ensure that $\widetilde{h}$ is odd or even, we can multiply $\widetilde{h}$ by an infinitely differentiable function, which is equal to one on $\Omega$ and zero for $u_{d+1}\leq 1/(2\sqrt{2})$, and extend $\widetilde{h}$ to be odd or even. These operations do not decrease the smoothness of $\widetilde{h}$. ∎
By Proposition 3.3, for any $h\in\mathcal{H}^{\alpha}$ and $g\in\mathcal{G}_{\sigma_{k}}(M)$, we have
$\|h-S_{k}g\|_{L^{\infty}(\mathbb{B}^{d})}=\|S_{k}\widetilde{h}-S_{k}g\|_{L^{\infty}(\mathbb{B}^{d})}\leq 2^{k/2}\|\widetilde{h}-g\|_{L^{\infty}(\mathbb{S}^{d})},$
for some $\widetilde{h}\in C(\mathbb{S}^{d})$. Since $S_{k}g\in\mathcal{F}_{\sigma_{k}}(M)$, we can derive approximation bounds for $\mathcal{F}_{\sigma_{k}}(M)$ by studying the approximation capacity of $\mathcal{G}_{\sigma_{k}}(M)$. Now, we are ready to prove Theorem 2.1.
Proof of Theorem 2.1.
By Proposition 3.3, for any $h\in\mathcal{H}^{\alpha}$, there exists $\widetilde{h}\in C(\mathbb{S}^{d})$ such that $S_{k}\widetilde{h}=h$, $\|\widetilde{h}\|_{L^{\infty}(\mathbb{S}^{d})}\leq C$ and $\omega_{2s^{*}}(\widetilde{h},t)_{\infty}\leq Ct^{\alpha}$, where $s^{*}\in\mathbb{N}$ is the smallest integer such that $\alpha\leq 2s^{*}$. We choose $\widetilde{h}$ to be odd (even) if $k$ is even (odd). Using $\omega_{s}(\widetilde{h},t)_{2}\leq 2^{s-2s^{*}+2}\omega_{2s^{*}}(\widetilde{h},t)_{2}$ for $s>2s^{*}$ [Dai and Xu, 2013, Proposition 10.1.2] and the Marchaud inequality [Ditzian, 2006, Eq.(9.6)]
$\omega_{s}(\widetilde{h},t)_{2}\lesssim t^{s}\left(\int_{t}^{1}\frac{\omega_{2s^{*}}(\widetilde{h},\theta)_{2}^{2}}{\theta^{2s+1}}d\theta\right)^{1/2},\quad s<2s^{*},$
we have
$\omega_{s}(\widetilde{h},t)_{2}\lesssim\begin{cases}t^{s},&\mbox{if }s<\alpha,\\ t^{\alpha},&\mbox{if }s=\alpha=2s^{*},\\ t^{\alpha}\sqrt{\log(1/t)},&\mbox{if }s=\alpha\neq 2s^{*},\\ t^{\alpha},&\mbox{if }s>\alpha.\end{cases}$ (3.5)
We study how well $g\in\mathcal{G}_{\sigma_{k}}(M)$ approximates $\widetilde{h}$. It turns out that it is enough to consider a subset of $\mathcal{G}_{\sigma_{k}}(M)$ that contains functions of the form
$g(u)=\int_{\mathbb{S}^{d}}\phi(v)\sigma_{k}(u^{\intercal}v)d\tau_{d}(v),\quad u\in\mathbb{S}^{d},$
for some $\phi\in L^{2}(\mathbb{S}^{d})$. Note that $\gamma(g)\leq\inf_{\phi}\|\phi\|_{L^{1}(\mathbb{S}^{d})}\leq\inf_{\phi}\|\phi\|_{L^{2}(\mathbb{S}^{d})}$, where the infimum is taken over all $\phi\in L^{2}(\mathbb{S}^{d})$ satisfy the integral representation of $g$. Observing that $g=\phi*\sigma_{k}$ is a convolution, by identity (3.2), $\mathcal{P}_{n}g=\widehat{\sigma_{k}}(n)\mathcal{P}_{n}\phi$. Hence, we have the Fourier decomposition
$g=\sum_{n=0}^{\infty}\mathcal{P}_{n}g=\sum_{n=0}^{\infty}\widehat{\sigma_{k}}(n)\mathcal{P}_{n}\phi,$
which converges in $L^{2}(\mathbb{S}^{d})$. This implies that $g\in\mathcal{F}_{\sigma_{k}}(M)$ if $\mathcal{P}_{n}g=0$ for any $n\in\mathbb{N}_{0}$ satisfying $\widehat{\sigma_{k}}(n)=0$ and
$\gamma(g)^{2}\leq\sum_{\widehat{\sigma_{k}}(n)\neq 0}\widehat{\sigma_{k}}(n)^{-2}\|\mathcal{P}_{n}g\|^{2}_{L^{2}(\mathbb{S}^{d})}\leq M^{2}.$
By Proposition 3.2, we know that $\widehat{\sigma_{k}}(n)=0$ if and only if $n\geq k+1$ and $n\equiv k\bmod 2$. For $\widehat{\sigma_{k}}(n)\neq 0$, we have $\widehat{\sigma_{k}}(n)\asymp n^{-(d+2k+1)/2}$.
We consider the convolutions $g_{m}:=\widetilde{h}*L_{m}=\int_{\mathbb{S}^{d}}\widetilde{h}(u)L_{m}(u^{\intercal}v)d\tau_{d}(v)$ with
$L_{m}(t):=\sum_{n=0}^{\infty}\eta\left(\frac{n}{m}\right)N(d,n)P_{n}(t),\quad m\in\mathbb{N},$
where $\eta$ is a $C^{\infty}$-function on $0,\infty)$ such that $\eta(t)=1$ for $0\leq t\leq 1$ and $\eta(t)=0$ for $t\geq 2$. Since $\eta$ is supported on $[0,2]$, the summation can be terminated at n=2m1n=2m-1, so that $g_{m}$ is a polynomial of degree at most 2m12m-1. Since $\widetilde{h}$ is odd (even) if $k$ is even (odd), $\mathcal{P}_{n}g_{m}=\eta(n/m)\mathcal{P}_{n}\widetilde{h}=0$ for any $n\equiv k\bmod 2$. Furthermore, Dai and Xu [2013, Theorem 10.3.2] shows that
$K_{s}(\widetilde{h},m^{-1})_{p}\asymp\|\widetilde{h}-g_{m}\|_{L^{p}(\mathbb{S}^{d})}+m^{-s}\|(-\Delta)^{s/2}g_{m}\|_{L^{p}(\mathbb{S}^{d})}.$
By the equivalence (3.4) and $\omega_{2s^{*}}(\widetilde{h},m^{-1})_{\infty}\lesssim m^{-\alpha}$, we can bound the approximation error as
$\|\widetilde{h}-g_{m}\|_{L^{\infty}(\mathbb{S}^{d})}\lesssim m^{-\alpha}.$
By the estimate (3.5), we get
$\|(-\Delta)^{s/2}g_{m}\|_{L^{2}(\mathbb{S}^{d})}\lesssim\begin{cases}1,&\mbox{if }s<\alpha,\\ 1,&\mbox{if }s=\alpha=2s^{*},\\ \sqrt{\log m},&\mbox{if }s=\alpha\mbox{ and }\alpha\neq 2s^{*},\\ m^{s-\alpha},&\mbox{if }\alpha<s\leq 2s^{*}.\end{cases}$
Using $\mathcal{P}_{n}((-\Delta)^{s/2}g_{m})=(n(n+d-1))^{s/2}\mathcal{P}_{n}g_{m}$, we can estimate the norm $\gamma(g_{m})$ as follows
$\gamma(g_{m})^{2}$ $\leq\sum_{\widehat{\sigma_{k}}(n)\neq 0}\widehat{\sigma_{k}}(n)^{-2}\|\mathcal{P}_{n}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})}$
$\lesssim\widehat{\sigma_{k}}(0)^{-2}\|\mathcal{P}_{0}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})}+\sum_{n=1}^{2m-1}n^{d+2k+1}n^{-2s}\|\mathcal{P}_{n}((-\Delta)^{s/2}g_{m})\|^{2}_{L^{2}(\mathbb{S}^{d})}$
$\lesssim 1+\sum_{n=1}^{2m-1}n^{d+2k+1-2s}\|\mathcal{P}_{n}((-\Delta)^{s/2}g_{m})\|^{2}_{L^{2}(\mathbb{S}^{d})}$
$\lesssim 1+\|(-\Delta)^{s/2}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})},$
where we choose $s=(d+2k+1)/2$ in the last inequality. There are three cases:
Case I: $\alpha>(d+2k+1)/2$ or $\alpha=(d+2k+1)/2$ is an even integer. In this case, $s<\alpha$ or $s=\alpha=2s^{*}$. Thus,
$\gamma(g_{m})^{2}\leq\sum_{\widehat{\sigma_{k}}(n)\neq 0}\widehat{\sigma_{k}}(n)^{-2}\|\mathcal{P}_{n}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})}\lesssim 1+\|(-\Delta)^{s/2}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})}\lesssim 1.$
Since $\mathcal{P}_{n}g_{m}=\eta(n/m)\mathcal{P}_{n}\widetilde{h}=\mathcal{P}_{n}\widetilde{h}$ for $n\leq m$, we have
$\gamma(\widetilde{h})^{2}$ $\leq\lim_{m\to\infty}\sum_{n\leq m,\widehat{\sigma_{k}}(n)\neq 0}\widehat{\sigma_{k}}(n)^{-2}\|\mathcal{P}_{n}\widetilde{h}\|^{2}_{L^{2}(\mathbb{S}^{d})}$
$\leq\lim_{m\to\infty}\sum_{\widehat{\sigma_{k}}(n)\neq 0}\widehat{\sigma_{k}}(n)^{-2}\|\mathcal{P}_{n}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})}\lesssim 1.$
This shows that $\widetilde{h}\subseteq\mathcal{G}_{\sigma_{k}}(M)$ for some constant $M$. Hence, $h=S_{k}\widetilde{h}\in\mathcal{F}_{\sigma_{k}}(M)$ by Proposition 3.3.
Case II: $\alpha=(d+2k+1)/2$ is not an even integer. We have $s=\alpha\neq 2s^{*}$ and
$\gamma(g_{m})^{2}\lesssim 1+\|(-\Delta)^{s/2}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})}\lesssim\log m.$
This shows that $g_{m}\in\mathcal{G}_{\sigma_{k}}(M)$ with $M\lesssim\sqrt{\log m}$. Therefore,
$\|\widetilde{h}-g_{m}\|_{L^{\infty}(\mathbb{S}^{d})}\lesssim m^{-\alpha}\lesssim\exp(-\alpha M^{2}).$
By Proposition 3.3, $f:=S_{k}g_{m}\in\mathcal{F}_{\sigma_{k}}(M)$ and
$\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\leq 2^{k/2}\|\widetilde{h}-g_{m}\|_{L^{\infty}(\mathbb{S}^{d})}\lesssim\exp(-\alpha M^{2}).$
Case III: $\alpha<(d+2k+1)/2$. In this case, $s>\alpha$ and
$\gamma(g_{m})^{2}\lesssim 1+\|(-\Delta)^{s/2}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})}\lesssim m^{d+2k+1-2\alpha}.$
This shows that $g_{m}\in\mathcal{G}_{\sigma_{k}}(M)$ with $M\lesssim m^{(d+2k+1-2\alpha)/2}$. Therefore,
$\|\widetilde{h}-g_{m}\|_{L^{\infty}(\mathbb{S}^{d})}\lesssim m^{-\alpha}\lesssim M^{-\frac{2\alpha}{d+2k+1-2\alpha}}.$
By Proposition 3.3, $f:=S_{k}g_{m}\in\mathcal{F}_{\sigma_{k}}(M)$ and
$\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\leq 2^{k/2}\|\widetilde{h}-g_{m}\|_{L^{\infty}(\mathbb{S}^{d})}\lesssim M^{-\frac{2\alpha}{d+2k+1-2\alpha}},$
which finishes the proof. ∎
Remark 3.4.
Since we are only able to estimate the smoothness $\omega_{2s^{*}}(\widetilde{h},t)_{\infty}$ for even integer $2s^{*}$, we have an extra logarithmic factor for the bound $\omega_{s}(\widetilde{h},t)_{2}\lesssim t^{\alpha}\sqrt{\log(1/t)}$ in (3.5) when $s=\alpha\neq 2s^{*}$, due to the Marchaud inequality. Consequently, we can only obtain exponential convergence rate when $\alpha=(d+2k+1)/2$ is not an even integer. We conjecture the bound $\omega_{\alpha}(\widetilde{h},t)_{2}\lesssim t^{\alpha}$ holds for all $\alpha>0$. If this is the case, then the proof of Theorem 2.1 implies $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma_{k}}(M)$ for some constant $M$ when $\alpha\geq(d+2k+1)/2$.
4 Nonparametric regression
In this section, we apply our approximation results to nonparametric regression using neural networks. For simplicity, we will only consider ReLU activation function ($k=1$), which is the most popular activation in deep learning. To use the uniform approximation bound in Corollary 2.4, we also require that the dimension $d\geq 4$.
We consider the classical problem of learning a $d$-variate function $h\in\mathcal{H}$ from its noisy samples, where we will assume $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$ or $\mathcal{H}=\mathcal{F}_{\sigma}(1)$. Note that, due to Theorem 2.1, the results for $\mathcal{F}_{\sigma}(1)$ can be applied to $\mathcal{H}^{\alpha}$ with $\alpha>(d+3)/2$ by scaling the variation norm. Suppose we have a data set of $n\geq 2$ samples $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}\subseteq\mathbb{B}^{d}\times\mathbb{R}$ which are independently and identically generated from the regression model
$Y_{i}=h(X_{i})+\eta_{i},\quad X_{i}\sim\mu,\quad\eta_{i}\sim\mathcal{N}(0,V^{2}),\quad i=1,\dots,n,\quad h\in\mathcal{H},$ (4.1)
where $\mu$ is the marginal distribution of the covariates $X_{i}$ supported on $\mathbb{B}^{d}$, and $\eta_{i}$ is an i.i.d. Gaussian noise independent of $X_{i}$ (we will treat the variance $V^{2}$ as a fixed constant). We are interested in the empirical risk minimizer (ERM)
$f_{n}^{*}\in\operatorname*{argmin}_{f\in\mathcal{F}_{n}}\mathcal{L}_{n}(f):=\operatorname*{argmin}_{f\in\mathcal{F}_{n}}\frac{1}{n}\sum_{i=1}^{n}|f(X_{i})-Y_{i}|^{2},$ (4.2)
where $\mathcal{F}_{n}$ is a function class parameterized by neural networks. For simplicity, we assume here and in the sequel that the minimum above indeed exists. When this is not the case, our results also hold for any estimator which minimizes the empirical risk up to a small error term. The performance of the estimation is measured by the expected risk
$\mathcal{L}(f):=\mathbb{E}_{(X,Y)}[(f(X)-Y)^{2}]=\mathbb{E}_{X\sim\mu}[(f(X)-h(X))^{2}]+V^{2}.$
It is equivalent to evaluate the estimator by the excess risk
$\|f-h\|_{L^{2}(\mu)}^{2}=\mathcal{L}(f)-\mathcal{L}(h).$
In the statistical analysis of learning algorithms, we often require that the hypothesis class is uniformly bounded. We define the truncation operator $\mathcal{T}_{B}$ with level $B>0$ for real-valued functions $f$ as
$\mathcal{T}_{B}f(x):=\begin{cases}f(x)&\quad\mbox{if }|f(x)|\leq B,\\ \,{\rm sgn}\,(f(x))B&\quad\mbox{if }|f(x)|>B.\end{cases}$
Since we always assume the regression function $h$ is bounded, truncating the output of the estimator $f_{n}^{*}$ appropriately dose not increase the excess risk. We will estimate the convergence rate of $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}$, where $B_{n}\lesssim\log n$, as the number of samples $n\to\infty$.
4.1 Shallow neural networks
The rate of convergence of neural network regression estimates has been analyzed by many papers [McCaffrey and Gallant, 1994; Kohler and Krzyżak, 2005; Bauer and Kohler, 2019; Schmidt-Hieber, 2020; Nakada and Imaizumi, 2020; Kohler and Langer, 2021]. It is well-known that the optimal minimax rate of convergence for learning a regression function $h\in\mathcal{H}^{\alpha}$ is $n^{-2\alpha/(d+2\alpha)}$ [Stone, 1982]. This optimal rate has been established (up to logarithmic factors) for two-hidden-layers neural networks with certain squashing activation functions [Kohler and Krzyżak, 2005] and for deep ReLU neural networks [Schmidt-Hieber, 2020; Kohler and Langer, 2021]. For shallow networks, McCaffrey and Gallant [1994] proved a rate of $n^{-2\alpha/(2\alpha+d+5)+\epsilon}$ for a certain cosine squasher activation function. However, to the best of our knowledge, it is unknown whether shallow neural networks can achieve the optimal rate. In this section, we provide an affirmative answer to this question by proving that shallow ReLU neural networks can achieve the optimal rate for $\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$.
We will use the following lemma to analyze the convergence rate. It decomposes the error of the ERM into generalization error and approximation error, and bounds the generalization error by the covering number of the hypothesis class $\mathcal{F}_{n}$.
Lemma 4.1 (Kohler and Langer [2021]).
Let $f_{n}^{*}$ be the estimator (4.2) and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$. Then,
$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}$
$\leq$ $\frac{c_{2}(\log n)^{2}\sup_{X_{1:n}\in(\mathbb{B}^{d})^{n}}\log(\mathcal{N}(n^{-1}B_{n}^{-1},\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})+1)}{n}+2\inf_{f\in\mathcal{F}_{n}}\|f-h\|_{L^{2}(\mu)}^{2},$
for $n>1$ and some constant $c_{2}>0$ (independent of $n$ and $f_{n}^{*}$), where $X_{1:n}=(X_{1},\dots,X_{n})$ denotes a sequence of sample points in $\mathbb{B}^{d}$ and $\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})$ denotes the $\epsilon$-covering number of the function class $\mathcal{T}_{B_{n}}\mathcal{F}_{n}:=\{\mathcal{T}_{B_{n}}f,f\in\mathcal{F}_{n}\}$ in the metric $\|f-g\|_{L^{1}(X_{1:n})}=\frac{1}{n}\sum_{i=1}^{n}|f(X_{i})-g(X_{i})|$.
For shallow neural network model $\mathcal{F}_{n}=\mathcal{F}_{\sigma}(N_{n},M_{n})$, Lemma 2.3 and Corollary 2.4 provide bounds for the approximation errors. The covering number of $\mathcal{F}_{n}$ can be estimated by using the Lipschitz continuity of the activation function, since the weights in the network are bounded. Choosing $N_{n},M_{n}$ appropriately to balance the approximation and generalization errors, we can derive convergence rates for the ERM.
Theorem 4.2.
Let $d\geq 4$, $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\mathcal{F}_{\sigma}(N_{n},M_{n})$ and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$.
  1. (1)
    If $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, we choose
    $N_{n}\asymp n^{\frac{d}{d+2\alpha}},\quad M_{n}\asymp n^{\frac{d+3-2\alpha}{2d+4\alpha}},$
    then
    $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}(\log n)^{3}.$
  2. (2)
    If $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, we choose
    $N_{n}\asymp n^{\frac{d}{2d+3}},\quad M_{n}=M,$
    for some sufficiently large constant $M\geq 1$, then
    $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{d+3}{2d+3}}(\log n)^{3}.$
Proof.
To apply the bound in Lemma 4.1, we need to estimate the covering number $\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})$. It is easy to check that $\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})\leq\mathcal{N}(\epsilon,\mathcal{F}_{n},\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})$. Hence, it is enough to provide a bound for $\mathcal{N}(\epsilon,\mathcal{F}_{n},\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})$.
Recall that any $f\in\mathcal{F}_{n}$ can be represented in the form $f(x)=\sum_{i=1}^{N_{n}}a_{i}\sigma((x^{\intercal},1)v_{i})$ with $v_{i}\in\mathbb{S}^{d}$ and $\sum_{i=1}^{N_{n}}|a_{i}|\leq M_{n}$. It is easy to see that there exists $\epsilon/(2\sqrt{2}M_{n})$-cover $V_{\epsilon}$ of $\mathbb{S}^{d}$ such that its cardinality $|V_{\epsilon}|\lesssim(1+M_{n}/\epsilon)^{d}$. Similarly, there exists $\epsilon/(2\sqrt{2})$-cover $A_{\epsilon}$ of the set $A=\{a=(a_{1},\dots,a_{N_{n}}):\|a\|_{1}\leq M_{n}\}$ in the metric $\|\cdot\|_{1}$ such that $|A_{\epsilon}|\lesssim(1+M_{n}/\epsilon)^{N_{n}}$. Hence, for any $v_{i}\in\mathbb{S}^{d}$ and $a=(a_{1},\dots,a_{N_{n}})\in A$, there exists $\widetilde{v}_{i}\in V_{\epsilon}$ and $\widetilde{a}=(\widetilde{a}_{1},\dots,\widetilde{a}_{N_{n}})\in A_{\epsilon}$ satisfying $\|v_{i}-\widetilde{v}_{i}\|_{2}\leq\epsilon/(2\sqrt{2}M_{n})$ and $\|a-\widetilde{a}\|_{1}\leq\epsilon/(2\sqrt{2})$. Therefore, letting $\widetilde{f}(x)=\sum_{i=1}^{N_{n}}\widetilde{a}_{i}\sigma((x^{\intercal},1)\widetilde{v}_{i})$, we have
$|f(x)-\widetilde{f}(x)|$ $\leq\|a-\widetilde{a}\|_{1}\sup_{1\leq i\leq N_{n}}|\sigma((x^{\intercal},1)v_{i})|+\|\widetilde{a}\|_{1}\sup_{1\leq i\leq N_{n}}|\sigma((x^{\intercal},1)v_{i})-\sigma((x^{\intercal},1)\widetilde{v}_{i})|$
$\leq\epsilon/2+\sqrt{2}M_{n}\sup_{1\leq i\leq N_{n}}\|v_{i}-\widetilde{v}_{i}\|_{2}\leq\epsilon,$
where we use the Lipschitz continuity of ReLU in the second inequality. This shows that
$\log\mathcal{N}(\epsilon,\mathcal{F}_{n},\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})\lesssim N_{n}\log|V_{\epsilon}|+\log|A_{\epsilon}|\lesssim N_{n}\log(1+M_{n}/\epsilon).$
Applying Lemma 4.1 and Corollary 2.4, if $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, then
$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim\frac{(\log n)^{2}N_{n}\log(nM_{n}B_{n})}{n}+N_{n}^{-\frac{2\alpha}{d}}\lor M_{n}^{-\frac{4\alpha}{d+3-2\alpha}}.$
By choosing $N_{n}\asymp n^{d/(d+2\alpha)}$ and $M_{n}\asymp N_{n}^{(d+3-2\alpha)/(2d)}$, we get $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-2\alpha/(d+2\alpha)}(\log n)^{3}$. Similarly, by Lemma 4.1 and Lemma 2.3, if $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, there exists $M_{n}=M$ such that
$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim\frac{(\log n)^{2}N_{n}\log(nM_{n}B_{n})}{n}+N_{n}^{-\frac{d+3}{d}}.$
We choose $N_{n}\asymp n^{d/(2d+3)}$, then $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-(d+3)/(2d+3)}(\log n)^{3}$. ∎
Remark 4.3.
The Gaussian noise assumption on the model (4.1) can be weaken for Lemma 4.1 and hence for Theorem 4.2. We refer the reader to Kohler and Langer [2021, Appendix B, Lemma 18] for more details. Theorem 4.2 can be easily generalized to shallow ReLUk neural networks for $k\geq 1$ by using the same proof technique, if we assume that $\mu$ has a bounded density (so that we can use Corollary 2.4 to bound the approximation error in the $L^{2}(\mu)$ norm). For example, one can show that, if $h\in\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$, then we can choose $\mathcal{F}_{n}=\mathcal{F}_{\sigma_{k}}(N_{n},M_{n})$ with $N_{n}\asymp n^{\frac{d}{d+2\alpha}}$ and $M_{n}\asymp n^{\frac{d+2k+1-2\alpha}{2d+4\alpha}}$, such that $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}(\log n)^{3}$.
Theorem 4.2 shows that least square minimization using shallow ReLU neural networks can achieve the optimal rate $n^{-\frac{2\alpha}{d+2\alpha}}$ for learning functions in $\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$. For the function class $\mathcal{F}_{\sigma}(1)$, the rate $n^{-\frac{d+3}{2d+3}}$ is also minimax optimal as proven by Parhi and Nowak [2023, Lemma 25] (they studied a slightly different function class, but their result also holds for $\mathcal{F}_{\sigma}(1)$). Specifically, Siegel and Xu [2022, Theorem 4 and Theorem 8] give a sharp estimate for the metric entropy
$\log\mathcal{N}(\epsilon,\mathcal{F}_{\sigma}(1),\|\cdot\|_{L^{2}(\mathbb{B}^{d})})\asymp\epsilon^{-\frac{2d}{d+3}}.$
Combining this estimate with the classical result of Yang and Barron (see [Yang and Barron, 1999, Proposition 1] and [Wainwright, 2019, Chapter 15]), we get
$\inf_{\widehat{f}_{n}}\sup_{h\in\mathcal{F}_{\sigma}(1)}\mathbb{E}_{\mathcal{D}_{n}}\|\widehat{f}_{n}-h\|_{L^{2}(\mathbb{B}^{d})}^{2}\asymp n^{-\frac{d+3}{2d+3}},$
where the infimum is over all functions of the samples $\mathcal{D}_{n}$, which are generated from the model (4.1).
4.2 Deep neural networks and over-parameterization
There is a direct way to generalize the analysis in the last section to deep neural networks: we can implement shallow neural networks by sparse multi-layer neural networks with the same order of parameters, and estimate the approximation and generalization performance of the constructed networks. Since the optimal convergence rates of deep neural networks have already been established in [Schmidt-Hieber, 2020; Kohler and Langer, 2021], we do not pursue in this direction. Instead, we study the convergence rates of over-parameterized neural networks by using the idea discussed in [Jiao et al., 2023]. The reason for studying such networks is that, in modern applications of deep learning, the number of parameters in the networks is often much larger than the number of samples. However, in the convergence analysis of [Schmidt-Hieber, 2020; Kohler and Langer, 2021], the network that achieves the optimal rate is under-parameterized (see also the choice of $N_{n}$ in Theorem 4.2). Hence, these analysis can not explain the empirical performance of deep learning models used in practice.
Following Jiao et al. [2023], we consider deep neural networks with norm constraints on weight matrices. For $W,L\in\mathbb{N}$, we denote by $\mathcal{NN}(W,L)$ the set of functions that can be parameterized by ReLU neural networks in the form
$f^{(0)}(x)$ $=x\in\mathbb{R}^{d},$ (4.3)
$f^{(\ell+1)}(x)$ $=\sigma(A^{(\ell)}f^{(\ell)}(x)+b^{(\ell)}),\quad\ell=0,\dots,L-1,$
$f(x)$ $=A^{(L)}f^{(L)}(x)+b^{(L)},$
where $A^{(\ell)}\in\mathbb{R}^{N_{\ell+1}\times N_{\ell}}$, $b^{(\ell)}\in\mathbb{R}^{N_{\ell+1}}$ with $N_{0}=d,N_{L+1}=1$ and $\max\{N_{1},\dots,N_{L}\}=W$. The numbers $W$ and $L$ are called the width and depth of the neural network, respectively. Let us use the notation $f_{\theta}$ to emphasize that the neural network function is parameterized by $\theta=((A^{(0)},b^{(0)}),\dots,(A^{(L)},b^{(L)}))$. We can define a norm constraint on the weight matrices as follows
$\kappa(\theta):=\|(A^{(L)},b^{(L)})\|\prod_{\ell=0}^{L-1}\max\left\{\|(A^{(\ell)},b^{(\ell)})\|,1\right\},$
where we use $\|A\|:=\sup_{\|x\|_{\infty}\leq 1}\|Ax\|_{\infty}$ to denote the operator norm (induced by the $\ell^{\infty}$ norm) of a matrix $A=(a_{i,j})\in\mathbb{R}^{m\times n}$. It is well-known that $\|A\|$ is the maximum $1$-norm of the rows of $A$:
$\|A\|=\max_{1\leq i\leq m}\sum_{j=1}^{n}|a_{i,j}|.$
The motivation for such a definition of $\kappa(\theta)$ is discussed in [Jiao et al., 2023]. For $M\geq 0$, we denote by $\mathcal{NN}(W,L,M)$ as the set of functions $f_{\theta}\in\mathcal{NN}$ that satisfy $\kappa(\theta)\leq M$. It is shown by [Jiao et al., 2023, Proposition 2.5] that, if $W_{1}\leq W_{2},L_{1}\leq L_{2},M_{1}\leq M_{2}$, then $\mathcal{NN}(W_{1},L_{1},M_{1})\subseteq\mathcal{NN}(W_{2},L_{2},M_{2})$. (Strictly speaking, Jiao et al. [2023] use the convention that the bias $b^{(L)}=0$ in the last layer. But the results can be easily generalized to the case $b^{(L)}\neq 0$, see [Yang, 2022, Section 2.1] for details.)
To derive approximation bounds for deep neural networks, we consider the relationship of $\mathcal{F}_{\sigma}(N,M)$ and $\mathcal{NN}(N,1,M)$. The next proposition shows the function classes $\mathcal{NN}(N,1,M)$ and $\mathcal{F}_{\sigma}(N,M)$ have essentially the same approximation power.
Proposition 4.4.
For any $N\in\mathbb{N}$ and $M>0$, we have $\mathcal{F}_{\sigma}(N,M)\subseteq\mathcal{NN}(N,1,\sqrt{d+1}M)$ and $\mathcal{NN}(N,1,M)\subseteq\mathcal{F}_{\sigma}(N+1,M)$.
Proof.
Each function $f(x)=\sum_{i=1}^{N}a_{i}\sigma((x^{\intercal},1)v_{i})$ in $\mathcal{F}_{\sigma}(N,M)$ can be parameterized in the form (4.3) with $W=N,L=1$ and
$(A^{(0)},b^{(0)})=(v_{1},\dots,v_{N})^{\intercal},\quad(A^{(1)},b^{(1)})=(a_{1},\dots,a_{N},0).$
Since $v_{i}\in\mathbb{S}^{d}$, it is easy to see that $\kappa(\theta)\leq\sqrt{d+1}M$. Hence, $\mathcal{F}_{\sigma}(N,M)\subseteq\mathcal{NN}(N,1,\sqrt{d+1}M)$.
On the contrary, let $f_{\theta}\in\mathcal{NN}(N,1,M)$ be a function parameterized in the form (4.3) with $(A^{(0)},b^{(0)})=(a^{(0)}_{1},\dots,a^{(0)}_{N})^{\intercal}$ and $(A^{(1)},b^{(1)})=(a^{(1)}_{1},\dots,a^{(1)}_{N},b^{(1)})$, where $a^{(0)}_{i}\in\mathbb{R}^{d+1}$ and $a^{(1)}_{i},b^{(1)}\in\mathbb{R}$. Then, $f_{\theta}$ can be represented as
$f_{\theta}(x)=\sum_{i=1}^{N}a^{(1)}_{i}\|a^{(0)}_{i}\|_{2}\sigma\left((x^{\intercal},1)\frac{a^{(0)}_{i}}{\|a^{(0)}_{i}\|_{2}}\right)+b^{(1)}\sigma(1),$
where we assume $\|a^{(0)}_{i}\|_{2}\neq 0$ without loss of generality. Since
$\gamma(f_{\theta})\leq\sum_{i=1}^{N}|a^{(1)}_{i}|\|a^{(0)}_{i}\|_{2}+|b^{(1)}|\leq\|(A^{(0)},b^{(0)})\|\sum_{i=1}^{N}|a^{(1)}_{i}|+|b^{(1)}|\leq\kappa(\theta),$
we conclude that $\mathcal{NN}(N,1,M)\subseteq\mathcal{F}_{\sigma}(N+1,M)$. ∎
As a corollary of Theorem 2.1 and Lemma 2.3, we get the following approximation bounds for deep neural networks.
Corollary 4.5.
Let $d\geq 4$.
  1. (1)
    For $\mathcal{H}^{\alpha}$ with $0<\alpha<(d+3)/2$, we have
    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{NN}(W,L,M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim W^{-\frac{\alpha}{d}}\lor M^{-\frac{2\alpha}{d+3-2\alpha}}.$
  2. (2)
    For $\mathcal{F}_{\sigma}(1)$, there exists a constant $M\geq 1$ such that
    $\sup_{h\in\mathcal{F}_{\sigma}(1)}\inf_{f\in\mathcal{NN}(W,L,M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim W^{-\frac{1}{2}-\frac{3}{2d}}.$
Proof.
Part (1) is a direct consequence of Corollary 2.4 and $\mathcal{F}_{\sigma}(W,M)\subseteq\mathcal{NN}(W,L,\sqrt{d+1}M)$. Part (2) follows from Lemma 2.3. ∎
In the first part of Corollary 4.5, if we allow the width $W$ to be arbitrary large, say $W\gtrsim M^{2d/(d+3-2\alpha)}$, then we can bound the approximation error by the size of weights. Hence, this result can be applied to over-parameterized neural networks. (Note that, in Theorem 4.2, we use a different regime of the bound.) In part (2), the size of weights is bounded by a constant. We will show that this constant can be used to control the generalization error. Since the approximation error is bounded by $W$ and is independent of $M$, we do not have trade-off in the error decomposition of ERM and only need to choose $W$ sufficiently large to reduce the approximation error. Hence, it can also be applied to over-parameterized neural networks.
The approximation rate $M^{-\frac{2\alpha}{d+3-2\alpha}}$ for $\mathcal{H}^{\alpha}$ in Corollary 4.5 improves the rate $M^{-\frac{\alpha}{d+1}}$ proved by Jiao et al. [2023]. Using the upper bound for Rademacher complexity of $\mathcal{NN}(W,L,M)$ (see Lemma 4.6), Jiao et al. [2023] also gave an approximation lower bound $(M\sqrt{L})^{-\frac{2\alpha}{d-2\alpha}}$. For fixed depth $L$, our upper bound is very close to this lower bound. We conjecture that the rate in Corollary 4.5 is optimal with respect to $M$ (for fixed depth $L$). The discussion of optimality in the last part of Section 2 implies that the conjecture is true for shallow neural networks.
To control the generalization performance of over-parameterized neural networks, we need to have size-independent sample complexity bounds for such networks. Several methods have been applied to obtain such kind of bounds in recent works [Neyshabur et al., 2015, 2018; Bartlett et al., 2017; Golowich et al., 2020]. Here, we will use the result of Golowich et al. [2020], which estimates the Rademacher complexity of deep neural networks [Bartlett and Mendelson, 2002]. For a set $S\subseteq\mathbb{R}^{n}$, let us denote its Rademacher complexity by
$\mathcal{R}_{n}(S):=\mathbb{E}_{\xi_{1:n}}\left[\sup_{(s_{1},\dots,s_{n})\in S}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}s_{i}\right],$
where $\xi_{1:n}=(\xi_{1},\dots,\xi_{n})$ is a sequence of i.i.d. Rademacher random variables. The following lemma is from Golowich et al. [2020, Theorem 3.2] and Jiao et al. [2023, Lemma 2.3].
Lemma 4.6 (Golowich et al. [2020]).
For any $x_{1},\dots,x_{n}\in[-1,1]^{d}$, let $S:=\{(f(x_{1}),\dots,f(x_{n})):f\in\mathcal{NN}(W,L,M)\}\subseteq\mathbb{R}^{n}$, then
$\mathcal{R}_{n}(S)\leq\frac{M\sqrt{2(L+2+\log(d+1))}}{\sqrt{n}}.$
Now, we can estimate the convergence rates of the ERM based on over-parameterized neural networks. As usual, we decompose the excess risk of the ERM into approximation error and generalization error, and bound them by Corollary 4.5 and Lemma 4.6, respectively. Note that the convergence rates in the following theorem are very close to the optimal rates in Theorem 4.2.
Theorem 4.7.
Let $d\geq 4$ and $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\{\mathcal{T}_{B_{n}}f:f\in\mathcal{NN}(W_{n},L,M_{n})\}$, where $L\in\mathbb{N}$ is a fixed constant, $1\leq B_{n}\lesssim\log n$ in case (1) and $\sqrt{2}\leq B_{n}\lesssim\log n$ in case (2).
  1. (1)
    If $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, we choose
    $M_{n}\asymp n^{\frac{1}{2}-\frac{2\alpha}{d+3+2\alpha}},\quad W_{n}\gtrsim n^{\frac{d}{d+3+2\alpha}},$
    then
    $\mathbb{E}_{\mathcal{D}_{n}}\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+3+2\alpha}}\log n.$
  2. (2)
    If $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, then there exist a constant $M\geq 1$ such that for $M_{n}=M$ and any $W_{n}\gtrsim n^{d/(2d+6)}$, we have
    $\mathbb{E}_{\mathcal{D}_{n}}\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-1/2}\log n.$
Proof.
The proof is essentially the same as Jiao et al. [2023, Theorem 4.1]. Observe that, for any $f\in\mathcal{F}_{n}$,
$\|f_{n}^{*}-h\|^{2}_{L^{2}(\mu)}=\mathcal{L}(f_{n}^{*})-\mathcal{L}(h)$
$=$ $\left[\mathcal{L}(f_{n}^{*})-\mathcal{L}_{n}(f_{n}^{*})\right]+\left[\mathcal{L}_{n}(f_{n}^{*})-\mathcal{L}_{n}(f)\right]+\left[\mathcal{L}_{n}(f)-\mathcal{L}(f)\right]+\left[\mathcal{L}(f)-\mathcal{L}(h)\right]$
$\leq$ $\left[\mathcal{L}(f_{n}^{*})-\mathcal{L}_{n}(f_{n}^{*})\right]+\left[\mathcal{L}_{n}(f)-\mathcal{L}(f)\right]+\|f-h\|^{2}_{L^{2}(\mu)}.$
Using $\mathbb{E}_{\mathcal{D}_{n}}[\mathcal{L}_{n}(f)]=\mathcal{L}(f)$ and taking the infimum over $f\in\mathcal{F}_{n}$, we get
$\mathbb{E}_{\mathcal{D}_{n}}\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\leq\inf_{f\in\mathcal{F}_{n}}\|f-h\|^{2}_{L^{2}(\mu)}+\mathbb{E}_{\mathcal{D}_{n}}\left[\mathcal{L}(f_{n}^{*})-\mathcal{L}_{n}(f_{n}^{*})\right].$ (4.4)
Let us denote the collections of sample points and noises by $X_{1:n}=(X_{1},\dots,X_{n})$ and $\eta_{1:n}=(\eta_{1},\dots,\eta_{n})$. We can bound the generalization error as follows
$\mathbb{E}_{\mathcal{D}_{n}}\left[\mathcal{L}(f_{n}^{*})-\mathcal{L}_{n}(f_{n}^{*})\right]$
$=$ $\mathbb{E}_{\mathcal{D}_{n}}\left[\|f_{n}^{*}-h\|^{2}_{L^{2}(\mu)}+V^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}(f_{n}^{*}(X_{i})-h(X_{i}))^{2}-2\eta_{i}(f_{n}^{*}(X_{i})-h(X_{i}))+\eta_{i}^{2}\right)\right]$
$=$ $\mathbb{E}_{X_{1:n}}\left[\|f_{n}^{*}-h\|^{2}_{L^{2}(\mu)}-\frac{1}{n}\sum_{i=1}^{n}(f_{n}^{*}(X_{i})-h(X_{i}))^{2}\right]+2\mathbb{E}_{\mathcal{D}_{n}}\left[\frac{1}{n}\sum_{i=1}^{n}\eta_{i}(f_{n}^{*}(X_{i})-h(X_{i}))\right]$
$\leq$ $\mathbb{E}_{X_{1:n}}\left[\sup_{\phi\in\Phi_{n}}\left(\mathbb{E}_{X}[\phi^{2}(X)]-\frac{1}{n}\sum_{i=1}^{n}\phi^{2}(X_{i})\right)\right]+2\mathbb{E}_{X_{1:n}}\mathbb{E}_{\eta_{1:n}}\left[\sup_{\phi\in\Phi_{n}}\frac{1}{n}\sum_{i=1}^{n}\eta_{i}\phi(X_{i})\right],$ (4.5)
where we denote $\Phi_{n}:=\{f-h:f\in\mathcal{F}_{n}\}$. By a standard symmetrization argument (see Wainwright [2019, Theorem 4.10]), we can bound the first term in (4.5) by the Rademacher complexity:
$\mathbb{E}_{X_{1:n}}\left[\sup_{\phi\in\Phi_{n}}\left(\mathbb{E}_{X}[\phi^{2}(X)]-\frac{1}{n}\sum_{i=1}^{n}\phi^{2}(X_{i})\right)\right]\leq 2\mathbb{E}_{X_{1:n}}\left[\mathcal{R}_{n}(\Phi_{n}^{2}(X_{1:n}))\right],$
where $\Phi_{n}^{2}(X_{1:n}):=\{(\phi^{2}(X_{1}),\dots,\phi^{2}(X_{n}))\in\mathbb{R}^{n}:\phi\in\Phi_{n}\}\subseteq\mathbb{R}^{n}$ is the set of function values on the sample points. Recall that we assume $B_{n}\geq 1$ in case (1) and $B_{n}\geq\sqrt{2}$ in case (2). Hence, we always have $\|h\|_{L^{\infty}(\mathbb{B}^{d})}\leq\sqrt{2}$ and $\|\phi\|_{L^{\infty}(\mathbb{B}^{d})}\leq 2B_{n}$ for any $\phi\in\Phi_{n}$. By the structural properties of Rademacher complexity [Bartlett and Mendelson, 2002, Theorem 12],
$\mathbb{E}_{X_{1:n}}\left[\mathcal{R}_{n}(\Phi_{n}^{2}(X_{1:n}))\right]$ $\leq 8B_{n}\mathbb{E}_{X_{1:n}}\left[\mathcal{R}_{n}(\Phi_{n}(X_{1:n}))\right]$
$\leq 8B_{n}\left(\mathbb{E}_{X_{1:n}}\left[\mathcal{R}_{n}(\mathcal{F}_{n}(X_{1:n}))\right]+\frac{\|h\|_{L^{\infty}(\mathbb{B}^{d})}}{\sqrt{n}}\right)$
$\lesssim\frac{M_{n}\log n}{\sqrt{n}},$
where we apply Lemma 4.6 in the last inequality. Note that the second term in (4.5) is a Gaussian complexity. We can also bound it by the Rademacher complexity [Bartlett and Mendelson, 2002, Lemma 4]:
$\mathbb{E}_{X_{1:n}}\mathbb{E}_{\eta_{1:n}}\left[\sup_{\phi\in\Phi_{n}}\frac{1}{n}\sum_{i=1}^{n}\eta_{i}\phi(X_{i})\right]\lesssim\mathbb{E}_{X_{1:n}}\left[\mathcal{R}_{n}(\Phi_{n}(X_{1:n}))\right]\log n\lesssim\frac{M_{n}\log n}{\sqrt{n}}.$
In summary, we conclude that
$\mathbb{E}_{\mathcal{D}_{n}}\left[\mathcal{L}(f_{n}^{*})-\mathcal{L}_{n}(f_{n}^{*})\right]\lesssim\frac{M_{n}\log n}{\sqrt{n}}.$ (4.6)
If $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, by Corollary 4.5, we have
$\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{n}}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim W_{n}^{-\frac{\alpha}{d}}\lor M_{n}^{-\frac{2\alpha}{d+3-2\alpha}}.$
Combining with (4.4) and (4.6), if we choose $M_{n}\asymp n^{\frac{1}{2}-\frac{2\alpha}{d+3+2\alpha}}$ and $W_{n}\gtrsim n^{\frac{d}{d+3+2\alpha}}$, then
$\mathbb{E}_{\mathcal{D}_{n}}\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim W_{n}^{-\frac{2\alpha}{d}}\lor M_{n}^{-\frac{4\alpha}{d+3-2\alpha}}+\frac{M_{n}\log n}{\sqrt{n}}\lesssim n^{-\frac{2\alpha}{d+3+2\alpha}}\log n.$
Similarly, if $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, by Corollary 4.5, then there exist a constant $M\geq 1$ such that, if $M_{n}=M$,
$\mathbb{E}_{\mathcal{D}_{n}}\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim W_{n}^{-\frac{d+3}{d}}+\frac{M\log n}{\sqrt{n}}.$
Thus, for any $W_{n}\gtrsim n^{d/(2d+6)}$, we get $\mathbb{E}_{\mathcal{D}_{n}}\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-1/2}\log n$, ∎
4.3 Convolutional neural networks
In contrast to the vast amount of theoretical studies on fully connected neural networks, there are only a few papers analyzing the performance of convolutional neural networks [Zhou, 2020a, b; Fang et al., 2020; Petersen and Voigtlaender, 2020; Lin et al., 2022]. The recent work [Lin et al., 2022] showed the universal consistency of CNNs for nonparametric regression. In this section, we show how to use our approximation results to analyze the convergence rates of CNNs.
Following [Zhou, 2020b], we introduce a sparse convolutional structure on deep neural networks. Let $s\geq 2$ be a fixed integer, which is used to control the filter length. Given a sequence $w=(w_{i})_{i\in\mathbb{Z}}$ on $\mathbb{Z}$ supported on $\{0,1,\dots,s\}$, the convolution of the filter $w$ with another sequence $x=(x_{i})_{i\in\mathbb{Z}}$ supported on $\{1,\dots,d\}$ is a sequence wxw*x given by
$(w*x)_{i}:=\sum_{j\in\mathbb{Z}}w_{i-j}x_{j}=\sum_{j=1}^{d}w_{i-j}x_{j},\quad i\in\mathbb{Z}.$
Regarding $x$ as a vector of $\mathbb{R}^{d}$, this convolution induces a $(d+s)\times d$ Toeplitz type convolutional matrix
$A^{w}:=(w_{i-j})_{1\leq i\leq d+s,1\leq j\leq d}.$
[…]
$b^{(\ell)}=\left(b^{(\ell)}_{1},\dots,b^{(\ell)}_{s},b^{(\ell)}_{s+1},\dots,b^{(\ell)}_{s+1},b^{(\ell)}_{N_{\ell}-s+1},\dots,b^{(\ell)}_{N_{\ell}}\right)^{\intercal},\quad 0\leq\ell\leq L-2,$ (4.7)
with $N_{\ell}-2s$ repeated components in the middle. By definition, it is easy to see that $\mathcal{CNN}(s,L)\subseteq\mathcal{NN}(d+Ls,L)$. The assumption on the special form (4.7) of biases is used to reduce the free parameters in the network. As in [Zhou, 2020b], one can compute that the number of free parameters in $\mathcal{CNN}(s,L)$ is (5s+2)L+2d2s(5s+2)L+2d-2s, which grows linearly on $L$.
The next proposition shows that all functions in $\mathcal{NN}(N,1)$ can be implemented by CNNs.
Proposition 4.8 (Zhou [2020b]).
If $N,L\in\mathbb{N}$ satisfy $L\geq\lfloor\frac{Nd}{s-1}+1\rfloor$, then $\mathcal{NN}(N,1)\subseteq\mathcal{CNN}(s,L)$.
Proof.
This result is proved in Zhou [2020b, Proof of Theorem 2]. We only give a sketch of the construction for completeness. Any function $f\in\mathcal{NN}(N,1)$ can be written as $f(x)=\sum_{i=1}^{N}c_{i}\sigma(a_{i}^{\intercal}x+b_{i})+c_{0}$, where $a_{i}\in\mathbb{R}^{d}$ and $b_{i},c_{i}\in\mathbb{R}$. Define a sequence $v$ supported on $\{0,\dots,Nd-1\}$ by stacking the vectors $a_{1},\dots,a_{N}$ (with components reversed) by
$(v_{Nd-1},\dots,v_{0})=(a_{N}^{\intercal},\dots,a_{1}^{\intercal}).$
Applying Zhou [2020b, Theorem 3] to the sequence $v$, we can construct filters $\{w^{(\ell)}\}_{\ell=0}^{L-1}$ supported on $\{0,1,\dots,s\}$ such that $v=w^{(L-1)}*w^{(L-2)}*\dots*w^{(0)}$, which implies $A^{w^{(L-1)}}\cdots A^{w^{(0)}}=A^{v}\in\mathbb{R}^{(d+Ls)\times d}$. Note that, by definition, for $i=1,\dots,N$, the $id$-th row of $A^{v}$ is exactly $a_{i}^{\intercal}$. Then, for $\ell=0,\dots,L-2$, we can choose $b^{(\ell)}$ satisfying (4.7) such that $f^{(\ell+1)}(x)=A^{w^{(\ell)}}\cdots A^{w^{(0)}}x+B^{(\ell)}$, where $B^{(\ell)}>0$ is a sufficiently large constant that makes the components of $f^{(\ell+1)}(x)$ positive for all $x\in\mathbb{B}^{d}$. Finally, we can construct $b^{(L-1)}$ such that $f^{(L)}_{k}(x)=\sigma(a_{i}^{\intercal}x+b_{i})$ for $i=1,\dots,N$ and $k=id$, which implies $f\in\mathcal{CNN}(s,L)$. ∎
Note that Proposition 4.8 shows each shallow neural network can be represent by a CNN, with the same order of number of parameters. As a corollary, we obtain approximation rates for CNNs.
Corollary 4.9.
Let $s\geq 2$ and $d\geq 4$ be integers.
  1. (1)
    For $\mathcal{H}^{\alpha}$ with $0<\alpha<(d+3)/2$, we have
    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{CNN}(s,L)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim L^{-\frac{\alpha}{d}}.$
  2. (2)
    For $\mathcal{F}_{\sigma}(1)$, we have
    $\sup_{h\in\mathcal{F}_{\sigma}(1)}\inf_{f\in\mathcal{CNN}(s,L)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim L^{-\frac{1}{2}-\frac{3}{2d}}.$
Proof.
For any $N\in\mathbb{N}$, we take $L=\lfloor\frac{Nd}{s-1}+1\rfloor\asymp N$, then $\mathcal{F}_{\sigma}(N,M)\subseteq\mathcal{NN}(N,1)\subseteq\mathcal{CNN}(s,L)$ for any $M>0$, by Proposition 4.8. (1) follows from Corollary 2.4 and (2) is from Lemma 2.3. ∎
Since the number of parameters in $\mathcal{CNN}(s,L)$ is approximately $L$, the rate $\mathcal{O}(L^{-\alpha/d})$ in part (1) of Corollary 4.9 is the same as the rate in [Yarotsky, 2017] for fully connected neural networks. However, Yarotsky [2018]; Lu et al. [2021] showed that this rate can be improved to $\mathcal{O}(L^{-2\alpha/d})$ for fully connected neural networks by using the bit extraction technique [Bartlett et al., 2019]. It would be interesting to see whether this rate also holds for $\mathcal{CNN}(s,L)$.
As in Theorem 4.2, we use Lemma 4.1 to decompose the error and bound the approximation error by Corollary 4.9. But we need to use pseudo-dimension to bound the covering number, because we do not have estimate for the size of weights in CNNs.
Theorem 4.10.
Let $d\geq 4$ and $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\mathcal{CNN}(s,L_{n})$, where $s\geq 2$ is a fixed integer, and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$.
  1. (1)
    If $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, we choose
    $L_{n}\asymp n^{\frac{d}{2d+2\alpha}},$
    then
    $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{\alpha}{d+\alpha}}(\log n)^{4}.$
  2. (2)
    If $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, we choose
    $L_{n}\asymp n^{\frac{d}{3d+3}},$
    then
    $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{d+3}{3d+3}}(\log n)^{4}.$
Proof.
To apply Lemma 4.1, we need to estimate the covering number $\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})$. The classical result of Haussler [1992, Theorem 6] showed that the covering number can be bounded by pseudo-dimension:
$\log\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})\lesssim\,{\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})\log(B_{n}/\epsilon),$
where ${\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})$ is the pseudo-dimension of the function class $\mathcal{T}_{B_{n}}\mathcal{F}_{n}$, see (2.2). For ReLU neural networks, Bartlett et al. [2019] gave the following estimate of the pseudo-dimension:
${\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})\lesssim L_{n}p(s,L_{n})\log(q(s,L_{n}))\lesssim L_{n}^{2}\log L_{n},$
where $p(s,L_{n})=(5s+2)L_{n}+2d-2s\lesssim L_{n}$ and $q(s,L_{n})\leq L_{n}(d+sL_{n})\lesssim L_{n}^{2}$ are the numbers of parameters and neurons of the network $\mathcal{CNN}(s,L_{n})$, respectively. Therefore,
$\log\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})\lesssim L_{n}^{2}\log(L_{n})\log(B_{n}/\epsilon).$
Applying Lemma 4.1 and Corollary 4.9, if $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, then
$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim\frac{L_{n}^{2}\log(L_{n})(\log n)^{3}}{n}+L_{n}^{-\frac{2\alpha}{d}}.$
We choose $L_{n}\asymp n^{d/(2d+2\alpha)}$, then $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\alpha/(d+\alpha)}(\log n)^{4}$. Similarly, if $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, then
$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim\frac{L_{n}^{2}\log(L_{n})(\log n)^{3}}{n}+L_{n}^{-\frac{d+3}{d}}.$
We choose $L_{n}\asymp n^{d/(3d+3)}$, then $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-(d+3)/(3d+3)}(\log n)^{4}$. ∎
5 Conclusion
This paper has established approximation bounds for shallow ReLUk neural networks. We showed how to use these bounds to derive approximation rates for (deep or shallow) neural networks with constraints on the weights and convolutional neural networks. We also applied the approximation results to study the convergence rates of nonparametric regression using neural networks. In particular, we established the optimal convergence rates for shallow neural networks and showed that over-parameterized neural networks can achieve nearly optimal rates.
There are a few interesting questions we would like to propose for future research. First, for shallow neural networks approximation, the optimal rate $\mathcal{O}(N^{-\alpha/d})$ in the supremum norm (in Corollary 2.4) is established only for ReLU and Heaviside activations. Can this result be extended to all ReLUk activations? Second, it is unclear whether over-parameterized neural networks can achieve the optimal rate for learning functions in $\mathcal{H}^{\alpha}$. It seems that refined generalization error analysis is needed. Finally, it would be interesting to extend the theory developed in this paper to general activation functions.
Acknowledgments
The work described in this paper was partially supported by InnoHK initiative, The Government of the HKSAR, Laboratory for AI-Powered Financial Technologies, the Research Grants Council of Hong Kong [Projects No. CityU 11308121, 11306220, 11308020 and No. N_CityU102/20] and National Natural Science Foundation of China [Projects No. 11461161006].

### 09 Jan 2024 v3

@@
 1 Introduction
-Neural networks generate very popular function classes used in machine learning algorithms [Anthony and Bartlett, 2009; LeCun et al., 2015].
+Neural networks generate very popular function classes used in machine learning algorithms [2, 28].
 The fundamental building blocks of neural networks are ridge functions (also called neurons) of the form $x\in\mathbb{R}^{d}\mapsto\rho((x^{\intercal},1)v)$, where $\rho:\mathbb{R}\to\mathbb{R}$ is a continuous activation function and $v\in\mathbb{R}^{d+1}$ is a trainable parameter.
@@
 $f(x)=\sum_{i=1}^{N}a_{i}\rho((x^{\intercal},1)v_{i}),$ (1.1)
-is universal in the sense that it can approximate any continuous function on any compact set with desired accuracy when the number of neurons $N$ is sufficiently large [Cybenko, 1989; Hornik, 1991; Pinkus, 1999].
-The approximation and statistical properties of neural networks with different architectures have also been widely studied in the literature [Yarotsky, 2017; Zhou, 2020b; Bartlett et al., 2019; Schmidt-Hieber, 2020], especially when $\rho$ is a sigmoidal activation or $\rho$ is the ReLUk function $\sigma_{k}(t)=\max\{0,t\}^{k}$, the $k$-power of the rectified linear unit (ReLU) with $k\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
+is universal in the sense that it can approximate any continuous functions on any compact set with desired accuracy when the number of neurons $N$ is sufficiently large [11, 23, 48].
+The approximation and statistical properties of neural networks with different architectures have also been widely studied in the literature [64, 67, 6, 52], especially when $\rho$ is a sigmoidal activation or $\rho$ is the ReLUk function $\sigma_{k}(t)=\max\{0,t\}^{k}$, the $k$-power of the rectified linear unit (ReLU) with $k\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
 The main focus of this paper is rates of approximation by neural networks.
-For classical smooth function classes, such as Hölder functions, Mhaskar [1996] (see also [Pinkus, 1999, Theorem 6.8]) presented the optimal approximation rates for shallow neural networks, when the activation function $\rho\in C^{\infty}(\Omega)$ is not a polynomial on some open interval $\Omega$ (ReLUk does not satisfy this condition).
-Similar results have been established for deep ReLU neural networks in recent works [Yarotsky, 2017, 2018; Shen et al., 2020; Lu et al., 2021].
+For classical smooth function classes, such as Hölder functions, Mhaskar [39] (see also [48, Theorem 6.8]) presented approximation rates for shallow neural networks, when the activation function $\rho\in C^{\infty}(\Omega)$ is not a polynomial on some open interval $\Omega$ (ReLUk does not satisfy this condition).
+It is known that the rates obtained by Mhaskar are optimal if the network weights are required to be continuous functions of the target function.
+Recently, optimal rates of approximation have also been established for deep ReLU neural networks [64, 65, 54, 31], even without the continuity requirement on the network weights.
 All these approximation rates are obtained by using the idea that one can construct neural networks to approximate polynomials efficiently.
-There is another line of works [Barron, 1993; Makovoz, 1996; Klusowski and Barron, 2018; Siegel and Xu, 2020, 2022] studying the approximation rates for functions of certain integral forms (such as (1.2)) by using a random sampling argument due to Maurey [Pisier, 1981].
-In particular, Barron [1993] derived dimension independent approximation rate for sigmoid type activations and functions $h$, whose Fourier transform $\widehat{h}$ satisfies $\int_{\mathbb{R}^{d}}|\omega||\widehat{h}(\omega)|d\omega<\infty$.
-This result has been improved and generalized to ReLU in recent articles [Klusowski and Barron, 2018; Siegel and Xu, 2020, 2022].
-In this paper, we give a connection of these two lines of approximation results for neural networks (i.e.
-the constructive approximation of smooth functions and the random approximation of integral representation).
+There is another line of works [4, 34, 25, 56, 57] studying the approximation rates for functions of certain integral forms (such as (1.2)) by using a random sampling argument due to Maurey [49].
+In particular, Barron [4] derived dimension independent approximation rates for sigmoid type activations and functions $h$, whose Fourier transform $\widehat{h}$ satisfies $\int_{\mathbb{R}^{d}}|\omega||\widehat{h}(\omega)|d\omega<\infty$.
+This result has been improved and generalized to ReLU activation in recent articles [25, 56, 57].
+In this paper, we continue the study of these two lines of approximation theories for neural networks (i.e.
+the constructive approximation of smooth functions and the random approximation of integral representations).
+Our main result shows how well integral representations corresponding to ReLUk neural networks can approximate smooth functions.
+By combining this result with the random approximation theory of integral forms, we are able to establish the optimal rates of approximation for shallow ReLUk neural networks.
 Specifically, we consider the following function class defined on the unit ball $\mathbb{B}^{d}$ of $\mathbb{R}^{d}$ induced by vectors on unit sphere $\mathbb{S}^{d}$ of $\mathbb{R}^{d+1}$ as
 $\mathcal{F}_{\sigma_{k}}(M):=\left\{f(x)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu(v):\|\mu\|\leq M,x\in\mathbb{B}^{d}\right\},$ (1.2)
-which can be regarded as a shallow ReLUk neural network with infinite width [Bach, 2017].
-The variation norm restriction $\|\mu\|\leq M$ gives a constraint on the size of the weights in the network.
+which can be regarded as a shallow ReLUk neural network with infinite width [3].
+The restriction on the total variation $\|\mu\|:=|\mu|(\mathbb{S}^{d})\leq M$ gives a constraint on the size of the weights in the network.
 We study how well $\mathcal{F}_{\sigma_{k}}(M)$ approximates the unit ball of Hölder class $\mathcal{H}^{\alpha}$ with smoothness index $\alpha>0$ as $M\to\infty$.
@@
 And, for less smooth functions, we can characterize the approximation error by the variation norm.
-Furthermore, combining our result with the random approximation bounds from Bach [2017]; Siegel and Xu [2022], we are able to prove that shallow ReLUk neural network of the form (1.1) achieves the optimal approximation rate $\mathcal{O}(N^{-\alpha/d})$ for $\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$, which generalizes the result of [Mhaskar, 1996] to ReLUk.
+Furthermore, combining our result with the random approximation bounds from [3, 57, 55], we are able to prove that shallow ReLUk neural network of the form (1.1) achieves the optimal approximation rate $\mathcal{O}(N^{-\alpha/d})$ for $\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$, which generalizes the result of Mhaskar [39] to ReLUk activation.
 In addition to shallow neural networks, we can also apply our results to derive approximation bounds for multi-layer neural networks and convolutional neural networks (CNNs) when $k=1$ (ReLU activation $\sigma:=\sigma_{1}$).
-These approximation bounds can then be used to study the performances of machine learning algorithms based on neural networks [Anthony and Bartlett, 2009].
+These approximation bounds can then be used to study the performances of machine learning algorithms based on neural networks [2].
 Here, we illustrate the idea by studying the nonparametric regression problem.
-The goal is to learn a function $h$ from a hypothesis space $\mathcal{H}$ from its noisy samples
+The goal of this problem is to learn a function $h$ from a hypothesis space $\mathcal{H}$ from its noisy samples
 $Y_{i}=h(X_{i})+\eta_{i},\quad i=1,\dots,n,$
@@
 The error can be decomposed into two components: approximation error and generalization error (also called estimation error).
-For neural network models $\mathcal{F}_{n}$, our results provide bounds for the approximation errors, while the generalization errors can be bounded by the complexity of the models [Shalev-Shwartz and Ben-David, 2014; Mohri et al., 2018].
+For neural network models $\mathcal{F}_{n}$, our results provide bounds for the approximation errors, while the generalization errors can be bounded by the complexity of the models [53, 40].
 We study the cases $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$ or $\mathcal{H}=\mathcal{F}_{\sigma}(1)$ for three ReLU neural network models: shallow neural network, over-parameterized neural network, and CNN.
@@
   1. (1)
-    Shallow ReLU neural network $\mathcal{F}_{\sigma}(N,M)$, where $N$ is the number of neurons and $M$ is a bound for the variation norm that measures the size of the weights.
-We prove optimal approximation rates (in terms of $N$) for this model.
-It is also shown that this model can achieve the optimal convergence rates for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$, which complements the recent results for deep neural networks [Schmidt-Hieber, 2020; Kohler and Langer, 2021].
+    Shallow ReLU neural network $\mathcal{F}_{\sigma}(N,M)$, where NN is the number of neurons and $M$ is a bound for the variation norm that measures the size of the weights.
+We prove optimal approximation rates (in terms of NN) for this model.
+It is also shown that this model can achieve the optimal convergence rates for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$, which complements the recent results for deep neural networks [52, 27].
   2. (2)
-    Over-parameterized (deep or shallow) ReLU neural network $\mathcal{NN}(W,L,M)$ studied in [Jiao et al., 2023], where $W,L$ are the width and depth respectively, and $M$ is a constraint on the weight matrices.
-For fixed depth $L$, the generalization error for this model can be controlled by $M$ [Jiao et al., 2023].
+    Over-parameterized (deep or shallow) ReLU neural network $\mathcal{NN}(W,L,M)$ studied in [24], where $W,L$ are the width and depth respectively, and $M$ is a constraint on the weight matrices.
+For fixed depth $L$, the generalization error for this model can be controlled by $M$ [24].
 When $\mathcal{H}=\mathcal{H}^{\alpha}$, we characterize the approximation error by $M$, and allow the width $W$ to be arbitrary large so that the model can be over-parameterized (the number of parameters is larger than the number of samples).
@@
 Our result shows that this model can achieve nearly optimal convergence rates for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$.
-Both the approximation and convergence rates improve the results of [Jiao et al., 2023].
+Both the approximation and convergence rates improve the results of [24].
   3. (3)
-    Sparse convolutional ReLU neural network $\mathcal{CNN}(s,L)$ introduced by Zhou [2020b], where $L$ is the depth and $s\geq 2$ is a fixed integer that controls the filter length.
-This model is shown to be universal for approximation [Zhou, 2020b] and universal consistent for regression [Lin et al., 2022].
-We improve the approximation bound in [Zhou, 2020b] and give new convergence rates of this model for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$.
+    Sparse convolutional ReLU neural network $\mathcal{CNN}(s,L)$ introduced by [67], where $L$ is the depth and $s\geq 2$ is a fixed integer that controls the filter length.
+This model is shown to be universal for approximation [67] and universal consistent for regression [30].
+We improve the approximation bound in [67] and give new convergence rates of this model for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$.
 The approximation rates and convergence rates of nonparametric regression for these models are summarized in Table 1.1, where we use the notation $a\lor b:=\max\{a,b\}$.
@@
 Section 3 gives a proof of our main theorem.
-In Section 4, we apply our approximation results to study there neural network models and derive convergence rates for nonparametric regression using these models.
+In Section 4, we apply our approximation results to study these neural network models and derive convergence rates for nonparametric regression using these models.
 Section 5 concludes this paper with a discussion on possible future directions of research.
@@
 For simplicity, we will also denote the ReLU function by $\sigma:=\sigma_{1}$.
-The variation norm $\gamma(f)$ of $f$ is the infimum of $\|\mu\|$ over all decompositions of $f$ as (2.1) [Bach, 2017].
+The variation norm $\gamma(f)$ of $f$ is the infimum of $\|\mu\|$ over all decompositions of $f$ as (2.1) [3].
 By the compactness of $\mathbb{S}^{d}$, the infimum can be attained by some signed measure $\mu$.
@@
 The variation spaces corresponding to shallow neural networks have been studied by many researchers.
-We refer the reader to [Savarese et al., 2019; Ongie et al., 2020; Bartolucci et al., 2023; Parhi and Nowak, 2022, 2023; Siegel and Xu, 2020, 2022, 2023] for several other definitions and characterizations of these spaces.
+We refer the reader to [51, 44, 8, 45, 46, 55, 56, 57, 58] for several other definitions and characterizations of these spaces.
 We will also need the notion of classical smoothness of functions on Euclidean space.
@@
 $\|f\|_{C^{r,\beta}(\mathbb{R}^{d})}:=\max\left\{\|f\|_{C^{r}(\mathbb{R}^{d})},\max_{\|s\|_{1}=r}|\partial^{s}f|_{C^{0,\beta}(\mathbb{R}^{d})}\right\},$
-where $s=(s_{1},\dots,s_{d})\in\mathbb{N}_{0}^{d}$ is an multi-index and
-$\|f\|_{C^{r}(\mathbb{R}^{d})}:=\max_{\|s\|_{1}\leq r}\|\partial^{s}f\|_{L^{\infty}(\mathbb{R}^{d})},\quad|f|_{C^{0,\beta}(\mathbb{R}^{d})}:=\sup_{x\neq y\in\mathbb{R}^{d}}\frac{|f(x)-f(y)|}{\|x-y\|_{2}^{\beta}},$
-where we use $\|\cdot\|_{L^{\infty}}$ to denote the supremum norm, since we mainly consider continuous functions.
+where $s=(s_{1},\dots,s_{d})\in\mathbb{N}_{0}^{d}$ is a multi-index and
+$\|f\|_{C^{r}(\mathbb{R}^{d})}:=\max_{\|s\|_{1}\leq r}\|\partial^{s}f\|_{L^{\infty}(\mathbb{R}^{d})},\quad|f|_{C^{0,\beta}(\mathbb{R}^{d})}:=\sup_{x\neq y\in\mathbb{R}^{d}}\frac{|f(x)-f(y)|}{\|x-y\|_{2}^{\beta}}.$
+Here we use $\|\cdot\|_{L^{\infty}}$ to denote the supremum norm, since we mainly consider continuous functions.
 We write $C^{r,\beta}(\mathbb{B}^{d})$ for the Banach space of all restrictions to $\mathbb{B}^{d}$ of functions in $C^{r,\beta}(\mathbb{R}^{d})$.
 The norm is given by $\|f\|_{C^{r,\beta}(\mathbb{B}^{d})}=\inf\{\|g\|_{C^{r,\beta}(\mathbb{R}^{d})}:g\in C^{r,\beta}(\mathbb{R}^{d})\mbox{ and }g=f\mbox{ on }\mathbb{B}^{d}\}$.
-For convenience, we will denote the unit ball of $C^{r,\beta}(\mathbb{B}^{d})$ by $\mathcal{H}^{\alpha}:=\{f\in C^{r,\beta}(\mathbb{B}^{d}):\|f\|_{C^{r,\beta}(\mathbb{B}^{d})}\leq 1\}$.
-Note that, for $\alpha=1$, $\mathcal{H}^{\alpha}$ is Lipschitz continuous function class.
-Due to the universality of shallow neural networks [Pinkus, 1999], $\mathcal{F}_{\sigma_{k}}(M)$ can approximate any continuous functions on $\mathbb{B}^{d}$ if $M$ is sufficiently large.
+For convenience, we will denote the unit ball of $C^{r,\beta}(\mathbb{B}^{d})$ by
+$\mathcal{H}^{\alpha}:=\left\{f\in C^{r,\beta}(\mathbb{B}^{d}):\|f\|_{C^{r,\beta}(\mathbb{B}^{d})}\leq 1\right\}.$
+Note that, for $\alpha=1$, $\mathcal{H}^{\alpha}$ is a class of Lipschitz continuous functions.
+Due to the universality of shallow neural networks [48], $\mathcal{F}_{\sigma_{k}}(M)$ can approximate any continuous functions on $\mathbb{B}^{d}$ if $M$ is sufficiently large.
 Our main theorem estimates the rate of this approximation for Hölder class.
@@
 The proof of Theorem 2.1 is deferred to the next section.
-Our proof uses similar ideas as Bach [2017, Proposition 3], which obtained the same approximation rate for $\alpha=1$.
+Our proof uses similar ideas as [3, Proposition 3], which obtained the same approximation rate for $\alpha=1$ (with an additional logarithmic factor).
 The conclusion is more complicated for the critical value $\alpha=(d+2k+1)/2$.
@@
 $\mathcal{F}_{\sigma_{k}}(N,M):=\left\{f(x)=\sum_{i=1}^{N}a_{i}\sigma_{k}((x^{\intercal},1)v_{i}):v_{i}\in\mathbb{S}^{d},\sum_{i=1}^{N}|a_{i}|\leq M\right\},$
-where we restrict the measure $\mu$ to be a discrete one supported on at most $N$ points.
+where we restrict the measure $\mu$ to be a discrete one supported on at most NN points.
 The next proposition shows that any function in $\mathcal{F}_{\sigma_{k}}(M)$ is the limit of functions in $\mathcal{F}_{\sigma_{k}}(N,M)$ as $N\to\infty$.
-Proposition 2.2.
-Let $\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$ denote the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$ in $L^{\infty}(\mathbb{B}^{d})$.
-If $k\in\mathbb{N}$, then $\mathcal{F}_{\sigma_{k}}(1)=\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$.
-If $k=0$, then $\mathcal{F}_{\sigma_{0}}(1)\subseteq\widetilde{\mathcal{F}}_{\sigma_{0}}(1)$.
+Proposition $2.$
+For $k\in\mathbb{N}_{0}$, $\mathcal{F}_{\sigma_{k}}(1)$ is the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$ in $L^{\infty}(\mathbb{B}^{d})$.
 Proof.
-We first show that $\mathcal{F}_{\sigma_{k}}(1)\subseteq\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$ for $k\in\mathbb{N}_{0}$.
+Let us denote the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$ in $L^{\infty}(\mathbb{B}^{d})$ by $\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$.
+We first show that $\mathcal{F}_{\sigma_{k}}(1)\subseteq\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$.
 For any $f\in\mathcal{F}_{\sigma_{k}}(1)$ with the integral form $f(x)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu(v)$, we can decompose $f$ as
-$f(x)$ $=\|\mu_{+}\|\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)\frac{d\mu_{+}(v)}{\|\mu_{+}\|}-\|\mu_{-}\|\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)\frac{d\mu_{+}(v)}{\|\mu_{-}\|}$
+$f(x)$ $=\|\mu_{+}\|\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)\frac{d\mu_{+}(v)}{\|\mu_{+}\|}-\|\mu_{-}\|\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)\frac{d\mu_{-}(v)}{\|\mu_{-}\|}$
 $=:\|\mu_{+}\|f_{+}(x)-\|\mu_{-}\|f_{-}(x),$
@@
 We are going to approximate $f$ by uniform laws of large numbers.
-Let $\{v_{i}\}_{i=1}^{N}$ be $N$ i.i.d.
+Let $\{v_{i}\}_{i=1}^{N}$ be NN i.i.d.
 samples from $\mu$.
-By symmetrization argument (see Wainwright [2019, Theorem 4.10] for example), we can bound the expected approximation error by Rademacher complexity [Bartlett and Mendelson, 2002]:
+By symmetrization argument (see [60, Theorem 4.10] for example), we can bound the expected approximation error by Rademacher complexity [7]:
 $\mathbb{E}\left[\sup_{x\in\mathbb{B}^{d}}\left|f(x)-\frac{1}{N}\sum_{i=1}^{N}\sigma_{k}((x^{\intercal},1)v_{i})\right|\right]\leq 2\mathbb{E}\left[\sup_{x\in\mathbb{B}^{d}}\left|\frac{1}{N}\sum_{i=1}^{N}\epsilon_{i}\sigma_{k}((x^{\intercal},1)v_{i})\right|\right]=:\mathcal{E}_{k}(N),$
@@
 For $k\in\mathbb{N}$, the Lipschitz constant of $\sigma_{k}$ on $\sqrt{2},\sqrt{2}]$ is $k2^{(k-1)/2}$.
-By the contraction property of Rademacher complexity [Ledoux and Talagrand, 1991, Corollary 3.17],
+By the contraction property of Rademacher complexity [29, Corollary 3.17],
 $\mathcal{E}_{k}(N)$ $\leq\frac{k2^{(k+1)/2}}{N}\mathbb{E}\left[\sup_{x\in\mathbb{B}^{d}}\left|\sum_{i=1}^{N}\epsilon_{i}(x^{\intercal},1)v_{i}\right|\right]\leq\frac{k2^{k/2+1}}{N}\mathbb{E}\left[\left\|\sum_{i=1}^{N}\epsilon_{i}v_{i}\right\|_{2}\right]$
 $\leq\frac{k2^{k/2+1}}{N}\sqrt{\mathbb{E}\left[\left\|\sum_{i=1}^{N}\epsilon_{i}v_{i}\right\|_{2}^{2}\right]}=\frac{k2^{k/2+1}}{N}\sqrt{\mathbb{E}\left[\sum_{i=1}^{N}\left\|v_{i}\right\|_{2}^{2}\right]}=\frac{k2^{k/2+1}}{\sqrt{N}}.$
-For $k=0$, the VC dimension of the function class $\{f_{x}(v)=\sigma_{0}((x^{\intercal},1)v):x\in\mathbb{B}^{d}\}$ is at most $d$ [Wainwright, 2019, Proposition 4.20].
-Thus, we have the bound $\mathcal{E}_{0}(N)\lesssim\sqrt{d/N}$ by Wainwright [2019, Example 5.24].
+For $k=0$, the VC dimension of the function class $\{f_{x}(v)=\sigma_{0}((x^{\intercal},1)v):x\in\mathbb{B}^{d}\}$ is at most $d$ [60, Proposition 4.20].
+Thus, we have the bound $\mathcal{E}_{0}(N)\lesssim\sqrt{d/N}$ by [60, Example 5.24].
 Hence, $f$ is in the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$.
-Next, we show that $\widetilde{\mathcal{F}}_{\sigma_{k}}(1)\subseteq\mathcal{F}_{\sigma_{k}}(1)$ for $k\in\mathbb{N}$.
+Next, we show that $\widetilde{\mathcal{F}}_{\sigma_{k}}(1)\subseteq\mathcal{F}_{\sigma_{k}}(1)$ for $k\in\mathbb{N}_{0}$.
 Since $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)\subseteq\mathcal{F}_{\sigma_{k}}(1)$, we only need to show that $\mathcal{F}_{\sigma_{k}}(1)$ is closed in $L^{\infty}(\mathbb{B}^{d})$.
 Let $f_{n}(x)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu_{n}(v)$, where $\|\mu_{n}\|\leq 1$, be a convergent sequence with limit $f\in L^{\infty}(\mathbb{B}^{d})$.
-By the compactness of $\mathbb{S}^{d}$ and Prokhorov’s theorem, there exists a weakly convergent subsequence $\mu_{n_{i}}\to\mu$.
+It remains to show that $f\in\mathcal{F}_{\sigma_{k}}(1)$.
+For $k\in\mathbb{N}$, by the compactness of $\mathbb{S}^{d}$ and Prokhorov’s theorem, there exists a weakly convergent subsequence $\mu_{n_{i}}\to\mu$.
 In particular, $\|\mu\|\leq 1$ and for any $x\in\mathbb{B}^{d}$,
 $\lim_{n_{i}\to\infty}\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu_{n_{i}}(v)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu(v)=:\widetilde{f}$.
-By the compactness of $\mathbb{B}^{d}$, $f_{n_{i}}$ convergence uniformly to $\widetilde{f}$.
-Hence $f=\widetilde{f}\in\mathcal{F}_{\sigma_{k}}(1)$ and $\mathcal{F}_{\sigma_{k}}(1)$ is closed in $L^{\infty}(\mathbb{B}^{d})$.
+By the compactness of $\mathbb{B}^{d}$, $f_{n_{i}}$ converges uniformly to $\widetilde{f}$. Hence $f=\widetilde{f}\in\mathcal{F}_{\sigma_{k}}(1)$.
+For $k=0$, we use the idea from [58, Lemma 3].
+We can view $f_{n}$ as a Bochner integral $\int_{\mathbb{D}}i_{\mathbb{D}\to L^{2}(\mathbb{B}^{d})}d\mu_{n}$ of the inclusion map $i_{\mathbb{D}\to L^{2}(\mathbb{B}^{d})}$, where $\mathbb{D}:=\{g_{v}(x)=\sigma_{0}((x^{\intercal},1)v):v\in\mathbb{S}^{d}\}$.
+Notice that the set $\mathbb{D}\subseteq L^{2}(\mathbb{B}^{d})$ is compact, because the mapping $v\mapsto g_{v}$ is continuous.
+By Prokhorov’s theorem, there exists a weakly convergent subsequence $\mu_{n_{i}}\to\mu$.
+Let us denote $\widetilde{f}=\int_{\mathbb{D}}i_{\mathbb{D}\to L^{2}(\mathbb{B}^{d})}d\mu$, then $\widetilde{f}\in\mathcal{F}_{\sigma_{k}}(1)$ by viewing the Bochner integral as an integral over $\mathbb{S}^{d}$.
+If we choose a countable dense sequence $\{g_{j}\}_{j=1}^{\infty}$ of $L^{2}(\mathbb{B}^{d})$, then the weak convergence implies that
+$\lim_{n_{i}\to\infty}\langle g_{j},f_{n_{i}}\rangle_{L^{2}(\mathbb{B}^{d})}=\left\langle g_{j},\widetilde{f}\right\rangle_{L^{2}(\mathbb{B}^{d})},$
+for all $j$.
+The strong convergence $f_{n_{i}}\to f$ in $L^{\infty}(\mathbb{B}^{d})$ implies that the same equality for $f$ replacing $\widetilde{f}$.
+Therefore, $\langle g_{j},f\rangle_{L^{2}(\mathbb{B}^{d})}=\langle g_{j},\widetilde{f}\rangle_{L^{2}(\mathbb{B}^{d})}$ for all $j$, which shows $f=\widetilde{f}\in\mathcal{F}_{\sigma_{k}}(1)$.
 ∎
@@
 This rate can be improved if we take into account the smoothness of the activation function.
-Lemma 2.3 ([Bach, 2017; Ma et al., 2022; Siegel and Xu, 2022]).
-Let $2\leq p<\infty$, then there exists an $M\geq 1$ depending only on $k,p,d$ such that,
-$\sup_{f\in\mathcal{F}_{\sigma_{k}}(1)}\inf_{f_{N}\in\mathcal{F}_{\sigma_{k}}(N,M)}\|f-f_{N}\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{pk+1}{pd}}.$
-For $k\in\{0,1\}$ and $d>2k+1$, it holds that
-$\sup_{f\in\mathcal{F}_{\sigma}(1)}\inf_{f_{N}\in\mathcal{F}_{\sigma}(N,M)}\|f-f_{N}\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{2k+1}{2d}}.$
-Proof.
-Using similar argument as Siegel and Xu [2022, Lemma 1], one can show that the parameterization map $\mathcal{P}_{\sigma_{k}}:\mathbb{S}^{d}\to L^{p}(\mathbb{B}^{d})$ given by $\mathcal{P}_{\sigma_{k}}(v)=\sigma_{k}((x^{\intercal},1)v)$ is of smoothness class $\alpha=k+1/p$ in the sense that, for each coordinate chart $(V,\phi)$ and $\xi\in L^{q}(\mathbb{B}^{d})$ where $1/p+1/q=1$, the function $f_{\xi}(u)=\langle\xi,\mathcal{P}_{\sigma_{k}}(\phi^{-1}(u))\rangle$ satisfies $\|f_{\xi}\|_{C^{r,\beta}(\phi(V))}\leq C\|\xi\|_{L^{q}(\mathbb{B}^{d})}$ for some constant $C<\infty$.
-By Siegel and Xu [2022, Theorem 2], if the parameterization map $\mathcal{P}_{\sigma_{k}}$ is of smoothness class $\alpha$, then there exists an $M\geq 1$ such that for any $f$ in the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$ in $L^{p}(\mathbb{B}^{d})$, we have
-$\inf_{f_{N}\in\mathcal{F}_{\sigma_{k}}(N,M)}\|f-f_{N}\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{\alpha}{d}}.$
-By Proposition 2.2, we can apply this bound to $f\in\mathcal{F}_{\sigma_{k}}(1)$.
-The uniform bound for $k=0$ is obtained by [Ma et al., 2022, Theorem 4].
-For $k=1$, the uniform bound is proved by Bach [2017, Proposition 1], who showed that approximating $f\in\mathcal{F}_{\sigma}(1)$ by finitely many neurons is essentially equivalent to the approximation of a zonoid by zonotopes [Matoušek, 1996].
-The condition $d\geq 4$ is needed in order to use [Matoušek, 1996, Theorem 1].
-We note that the proof also implies that one can take $M=3/2$.
-∎
-One can show that the rate $\mathcal{O}(N^{-\frac{1}{2}-\frac{2k+1}{2d}})$ in Lemma 2.3 is sharp for $p=2$ [Siegel and Xu, 2022].
-Since the same rate holds in the supremum norm when $k=0,1$, we expect that it also holds for all $k\in\mathbb{N}_{0}$ and $2\leq p\leq\infty$.
-It seems that one needs to use techniques of geometric discrepancy theory to prove such kind of results [Bourgain et al., 1989; Matoušek, 1996].
-Combining Theorem 2.1 and Lemma 2.3, we can derive approximation rate of Hölder class $\mathcal{H}^{\alpha}$ for shallow neural network $\mathcal{F}_{\sigma_{k}}(N,M)$.
+For ReLU activation, Bach [3, Proposition 1] showed that approximating $f\in\mathcal{F}_{\sigma}(1)$ by neural networks with finitely many neurons is essentially equivalent to the approximation of a zonoid by zonotopes [10, 37].
+Using this equivalence, he obtained the rate $\mathcal{O}(N^{-\frac{1}{2}-\frac{3}{2d}})$ for ReLU neural networks.
+Similar idea was applied to the Heaviside activation in [32, Theorem 4], which proved the rate $\mathcal{O}(N^{-\frac{1}{2}-\frac{1}{2d}})$ for such an activation function.
+For ReLUk neural networks, the general approximation rate $\mathcal{O}(N^{-\frac{1}{2}-\frac{2k+1}{2d}})$ was established in $L^{2}$ norm by [57], which also showed that this rate is sharp.
+The recent work [55] further proved that this rate indeed holds in the uniform norm.
+We summarize their results in the following lemma.
+Lemma 2.3 ([55]).
+For $k\in\mathbb{N}_{0}$ and $d\in\mathbb{N}$, it holds that
+$\sup_{f\in\mathcal{F}_{\sigma_{k}}(1)}\inf_{f_{N}\in\mathcal{F}_{\sigma_{k}}(N,1)}\|f-f_{N}\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{2k+1}{2d}}.$
+Combining Theorem 2.1 and Lemma 2.3, we can derive the rate of approximation by shallow neural network $\mathcal{F}_{\sigma_{k}}(N,M)$ for Hölder class $\mathcal{H}^{\alpha}$.
 Recall that we use the notation $a\lor b:=\max\{a,b\}$.
 Corollary 2.4.
-Let $k\in\mathbb{N}_{0}$, $d\in\mathbb{N}$, $2\leq p<\infty$ and $\alpha>0$.
+Let $k\in\mathbb{N}_{0}$, $d\in\mathbb{N}$ and $\alpha>0$.
   1. (1)
-    If $\alpha>(d+2k+1)/2$ or $\alpha=(d+2k+1)/2$ is an even integer, then there exists a constant $M$ depending on $k,p,d,\alpha$ such that
-    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{pk+1}{pd}}.$
+    If $\alpha>(d+2k+1)/2$ or $\alpha=(d+2k+1)/2$ is an even integer, then there exists a constant $M$ depending on $k,d,\alpha$ such that
+    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{2k+1}{2d}}.$
   2. (2)
     If $\alpha=(d+2k+1)/2$ is not an even integer, then there exists $M\asymp\sqrt{\log N}$ such that
-    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{pk+1}{pd}}\sqrt{\log N}.$
+    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{2k+1}{2d}}\sqrt{\log N}.$
   3. (3)
     If $\alpha<(d+2k+1)/2$, then
-    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{\alpha}{d}\frac{d+2k+2/p}{d+2k+1}}\lor M^{-\frac{2\alpha}{d+2k+1-2\alpha}}.$
-  4. (4)
-    If $k\in\{0,1\}$ and $d>2k+1$, the bounds for $p=2$ also hold in the supremum norm. In particular, for $\alpha<(d+2k+1)/2$, we have
     $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{\alpha}{d}}\lor M^{-\frac{2\alpha}{d+2k+1-2\alpha}}.$
-    Thus, the rate $\mathcal{O}(N^{-\alpha/d})$ holds when $M\gtrsim N^{-(d+2k+1-2\alpha)/(2d)}$.
+    Thus, the rate $\mathcal{O}(N^{-\alpha/d})$ holds when $M\gtrsim N^{(d+2k+1-2\alpha)/(2d)}$.
 Proof.
@@
 If $\alpha<(d+2k+1)/2$, then by Theorem 2.1, for any $h\in\mathcal{H}^{\alpha}$, there exists $g\in\mathcal{F}_{\sigma_{k}}(K)$ such that $\|h-g\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}$.
-Let $M_{0}$ be the constant in Lemma 2.3, then there exists $f\in\mathcal{F}_{\sigma_{k}}(N,KM_{0})$ such that $\|g-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim KN^{-\frac{1}{2}-\frac{pk+1}{pd}}$.
-If $M\geq N^{\frac{d+2k+1-2\alpha}{2d}\frac{d+2k+2/p}{d+2k+1}}$, we choose $K=N^{\frac{d+2k+1-2\alpha}{2d}\frac{d+2k+2/p}{d+2k+1}}/M_{0}$ then $f\in\mathcal{F}_{\sigma_{k}}(N,KM_{0})\subseteq\mathcal{F}_{\sigma_{k}}(N,M)$ and
-$\|h-f\|_{L^{p}(\mathbb{B}^{d})}$ $\leq\|h-g\|_{L^{p}(\mathbb{B}^{d})}+\|g-f\|_{L^{p}(\mathbb{B}^{d})}$
-$\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}+KN^{-\frac{d+2k+2/p}{2d}}$
-$\lesssim N^{-\frac{\alpha}{d}\frac{d+2k+2/p}{d+2k+1}}.$
-If $M\leq N^{\frac{d+2k+1-2\alpha}{2d}\frac{d+2k+2/p}{d+2k+1}}$, we choose $K=M/M_{0}$, then
-$\|h-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}+KN^{-\frac{d+2k+2/p}{2d}}\lesssim M^{-\frac{2\alpha}{d+2k+1-2\alpha}}.$
-Combining the two bounds gives the desired result. By Lemma 2.3, if $k=0,1$ and $d>2k+1$, the bounds for $p=2$ also hold in the supremum norm. ∎
+By Lemma 2.3, then there exists $f\in\mathcal{F}_{\sigma_{k}}(N,K)$ such that $\|g-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim KN^{-\frac{1}{2}-\frac{2k+1}{2d}}$.
+If $M\geq N^{\frac{d+2k+1-2\alpha}{2d}}$, we choose $K=N^{\frac{d+2k+1-2\alpha}{2d}}$ then $f\in\mathcal{F}_{\sigma_{k}}(N,K)\subseteq\mathcal{F}_{\sigma_{k}}(N,M)$ and
+$\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}$ $\leq\|h-g\|_{L^{\infty}(\mathbb{B}^{d})}+\|g-f\|_{L^{\infty}(\mathbb{B}^{d})}$
+$\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}+KN^{-\frac{d+2k+1}{2d}}\lesssim N^{-\frac{\alpha}{d}}.$
+If $M\leq N^{\frac{d+2k+1-2\alpha}{2d}}$, we choose $K=M$, then
+$\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}+KN^{-\frac{d+2k+1}{2d}}\lesssim M^{-\frac{2\alpha}{d+2k+1-2\alpha}}.$
+Combining the two bounds gives the desired result. ∎
 We make some comments on the approximation rate for $\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$.
-As shown by Pinkus [1999, Corollary 6.10], the rate $\mathcal{O}(N^{-\alpha/d})$ in the $L^{2}$ norm is already known for $\alpha=1,2,\dots,(d+2k+1)/2$.
-For ReLU activation, the recent paper [Mao and Zhou, 2023] obtained the rate $\mathcal{O}(N^{-\frac{\alpha}{d}\frac{d+2}{d+4}})$ in the supremum norm.
-Corollary 2.4 shows that the rate $\mathcal{O}(N^{-\alpha/d})$ also holds in the supremum norm for ReLU and Heaviside activations.
-And more importantly, we also provide an explicit control on the network weights to ensure that such rate can be achieved, which is useful for estimating generalization errors (see Section 4).
-It is well-known that the optimal approximation rate for $\mathcal{H}^{\alpha}$ is $\mathcal{O}(N^{-\alpha/d})$, if we approximate $h\in\mathcal{H}^{\alpha}$ by a function class with $N$ parameters and the parameters are continuously depend on the target function $h$ [DeVore et al., 1989].
-However, this result is not exactly applicable to neural networks, because we do not have guarantee that the parameters in the network depend continuously on the target function (in fact, this is not true for some constructions [Yarotsky, 2018; Lu et al., 2021; Yang et al., 2022]).
-Nevertheless, one can still prove that the rate $\mathcal{O}(N^{-\alpha/d})$ is optimal for shallow ReLUk neural networks by arguments based on pseudo-dimension as done in [Yarotsky, 2017; Lu et al., 2021; Yang et al., 2022].
-We describe the idea of proving approximation lower bound through pseudo-dimension by reviewing the result of Maiorov and Ratsaby [1999] (see also [Achour et al., 2022]).
+As shown by [48, Corollary 6.10], the rate $\mathcal{O}(N^{-\alpha/d})$ in the $L^{2}$ norm is already known for $\alpha=1,2,\dots,(d+2k+1)/2$.
+For ReLU activation, the recent paper [36] obtained the rate $\mathcal{O}(N^{-\frac{\alpha}{d}\frac{d+2}{d+4}})$ in the supremum norm.
+Corollary 2.4 shows that the rate $\mathcal{O}(N^{-\alpha/d})$ holds in the supremum norm for all ReLUk activations.
+And more importantly, we also provide an explicit control on the network weights to ensure that this rate can be achieved, which is useful for estimating generalization errors (see Section 4.2).
+It is well-known that the optimal approximation rate for $\mathcal{H}^{\alpha}$ is $\mathcal{O}(N^{-\alpha/d})$, if we approximate $h\in\mathcal{H}^{\alpha}$ by a function class with NN parameters and the parameters are continuously dependent on the target function $h$ [13].
+However, this result is not directly applicable to neural networks, because we do not have guarantee that the parameters in the network depend continuously on the target function (in fact, this is not true for some constructions [65, 31, 63]).
+Nevertheless, one can still prove that the rate $\mathcal{O}(N^{-\alpha/d})$ is optimal for shallow ReLUk neural networks by arguments based on pseudo-dimension as done in [64, 31, 63].
+We describe the idea of proving approximation lower bounds through pseudo-dimension by reviewing the result of Maiorov and Ratsaby [33] (see also [1]).
 Recall that the pseudo-dimension ${\rm Pdim}\,(\mathcal{F})$ of a real-valued function class $\mathcal{F}$ defined on $\mathbb{B}^{d}$ is the largest integer $n$ for which there exist points $x_{1},\dots,x_{n}\in\mathbb{B}^{d}$ and constants $c_{1},\dots,c_{n}\in\mathbb{R}$ such that
@@
 $\rho_{n}(\mathcal{H}^{\alpha})=\inf_{\mathcal{F}_{n}}\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{n}}\|h-f\|_{L^{p}(\mathbb{B}^{d})},$
-where $\mathcal{F}_{n}$ runs over all the classes in $L^{p}(\mathbb{B}^{d})$ with ${\rm Pdim}\,(\mathcal{F}_{n})\leq n$.
+where $p\in[1,\infty]$ and $\mathcal{F}_{n}$ runs over all the classes in $L^{p}(\mathbb{B}^{d})$ with ${\rm Pdim}\,(\mathcal{F}_{n})\leq n$.
 They constructed a well-separated subclass of $\mathcal{H}^{\alpha}$ such that if a function class $\mathcal{F}$ can approximate this subclass with small error, then ${\rm Pdim}\,(\mathcal{F})$ should be large.
@@
 $\rho_{n}(\mathcal{H}^{\alpha})\gtrsim n^{-\alpha/d}.$
-By [Bartlett et al., 2019], we can upper bound the pseudo-dimension of shallow ReLUk neural networks as $n:=\,{\rm Pdim}\,(\mathcal{F}_{\sigma_{k}}(N,M))\lesssim N\log N$. Hence,
+By [6], we can upper bound the pseudo-dimension of shallow ReLUk neural networks as $n:=\,{\rm Pdim}\,(\mathcal{F}_{\sigma_{k}}(N,M))\lesssim N\log N$. Hence,
 $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f_{N}\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f_{N}\|_{L^{p}(\mathbb{B}^{d})}\geq\rho_{n}(\mathcal{H}^{\alpha})\gtrsim(N\log N)^{-\alpha/d},$
-which shows that the rate $\mathcal{O}(N^{-\alpha/d})$ in Corollary 2.4 is optimal in the $L^{2}$ norm (ignoring logarithmic factors).
+which shows that the rate $\mathcal{O}(N^{-\alpha/d})$ in Corollary 2.4 is optimal in the $L^{p}$ norm (ignoring logarithmic factors).
 This also implies the optimality of Theorem 2.1 (otherwise, the proof of Corollary 2.4 would give a rate better than $\mathcal{O}(N^{-\alpha/d})$).
 3 Proof of Theorem 2.1
-Following the idea of Bach [2017], we first transfer the problem to approximation on spheres.
-Let us begin with a brief review of harmonic analysis on spheres [Dai and Xu, 2013].
+Following the idea of [3], we first transfer the problem to approximation on spheres.
+Let us begin with a brief review of harmonic analysis on spheres [12].
 For $n\in\mathbb{N}_{0}$, the spherical harmonic space $\mathbb{Y}_{n}$ of degree $n$ is the linear space that contains the restrictions of real harmonic homogeneous polynomials of degree $n$ on $\mathbb{R}^{d+1}$ to the sphere $\mathbb{S}^{d}$.
@@
 Let $\mathcal{P}_{n}:L^{2}(\mathbb{S}^{d})\to\mathbb{Y}_{n}$ denote the orthogonal projection operator.
-For any orthonormal basis $\{Y_{nj}:1\leq j\leq N(d,n)\}$ of $\mathbb{Y}_{n}$, the addition formula [Dai and Xu, 2013, Theorem 1.2.6] shows
+For any orthonormal basis $\{Y_{nj}:1\leq j\leq N(d,n)\}$ of $\mathbb{Y}_{n}$, the addition formula [12, Theorem 1.2.6] shows
 $\sum_{j=1}^{N(d,n)}Y_{nj}(u)Y_{nj}(v)=N(d,n)P_{n}(u^{\intercal}v),\quad u,v\in\mathbb{S}^{d},$ (3.1)
@@
 For $n\neq 0$, $P_{n}(t)$ is odd (even) if $n$ is odd (even).
-Note that.
-for $d=1$ and $n\neq 0$, $N(d,n)=2$ and $P_{n}(t)$ is the Chebyshev polynomial such that $P_{n}(\cos\theta)=\cos(n\theta)$.
+Note that, for $d=1$ and $n\neq 0$, $N(d,n)=2$ and $P_{n}(t)$ is the Chebyshev polynomial such that $P_{n}(\cos\theta)=\cos(n\theta)$.
 We can write the projection $\mathcal{P}_{n}$ as
@@
 Definition 3.1 (Convolution).
-Let $\varrho$ be the probability distribution with density $c_{d}(1-t^{2})^{(d-2)/2}$ on [1,1][-1,1], where the constant $c_{d}=(\int_{-1}^{1}(1-t^{2})^{(d-2)/2}dt)^{-1}=\omega_{d-1}/\omega_{d}$.
+Let $\varrho$ be the probability distribution with density $c_{d}(1-t^{2})^{(d-2)/2}$ on [1,1][-1,1], with the constant $c_{d}=(\int_{-1}^{1}(1-t^{2})^{(d-2)/2}dt)^{-1}=\omega_{d-1}/\omega_{d}$.
 For $f\in L^{1}(\mathbb{S}^{d})$ and $g\in L^{1}_{\varrho}([-1,1])$, define
 $(f*g)(u):=\int_{\mathbb{S}^{d}}f(v)g(u^{\intercal}v)d\tau_{d}(v),\quad u\in\mathbb{S}^{d}.$
-The convolution on the sphere satisfies Young’s inequality [Dai and Xu, 2013, Theorem 2.1.2]: for $p,q,r\geq 1$ with $p^{-1}=q^{-1}+r^{-1}-1$, it holds
+The convolution on the sphere satisfies Young’s inequality [12, Theorem 2.1.2]: for $p,q,r\geq 1$ with $p^{-1}=q^{-1}+r^{-1}-1$, it holds
 $\|f*g\|_{L^{p}(\mathbb{S}^{d})}\leq\|f\|_{L^{q}(\mathbb{S}^{d})}\|g\|_{L^{r}_{\varrho}([-1,1])},$
-where the norm is the uniform norm when $r=\infty$.
+where the norm is the uniform one when $r=\infty$.
 Observe that the projection $\mathcal{P}_{n}f=f*(N(d,n)P_{n})$ is a convolution operator with $\|N(d,n)P_{n}\|_{L^{\infty}([-1,1])}\leq N(d,n)$.
@@
 $\widehat{g}(n):=\frac{\omega_{d-1}}{\omega_{d}}\int_{-1}^{1}g(t)P_{n}(t)(1-t^{2})^{(d-2)/2}dt.$
-By the Funk-Hecke formula, one can show that [Dai and Xu, 2013, Theorem 2.1.3]
+By the Funk-Hecke formula, one can show that [12, Theorem 2.1.3]
 $\mathcal{P}_{n}(f*g)=\widehat{g}(n)\mathcal{P}_{n}f,\quad f\in L^{1}(\mathbb{S}^{d}),n\in\mathbb{N}_{0}.$ (3.2)
@@
 One of the key steps in our proof of Theorem 2.1 is the observation that functions of the form $f(u)=\int_{\mathbb{S}^{d}}\phi(v)\sigma_{k}(u^{\intercal}v)d\tau_{d}(v)$ are convolutions $\phi*\sigma_{k}$ with the activation function $\sigma_{k}\in L^{\infty}([-1,1])$.
-Bach [2017, Appendix D.2] has computed the Fourier coefficients $\widehat{\sigma_{k}}(n)$ explicitly.
+[3, Appendix D.2] has computed the Fourier coefficients $\widehat{\sigma_{k}}(n)$ explicitly.
 We summarize the result in the following.
@@
 $\omega_{\alpha}(f,t)_{p}:=\sup_{0<\theta\leq t}\|\Delta_{\theta}^{\alpha}f\|_{L^{p}(\mathbb{S}^{d})},\quad 0<t<\pi.$
-For even integers $\alpha=2s$, one can also use combinations of $T_{j\theta}$ and obtain [Rustamov, 1993; Ditzian, 2006]
+For even integers $\alpha=2s$, one can also use combinations of $T_{j\theta}$ and obtain [50, 15]
 $\omega_{2s}(f,t)_{p}\asymp\sup_{0<\theta\leq t}\left\|\sum_{j=0}^{2s}(-1)^{j}\binom{2s}{j}T_{j\theta}f\right\|_{L^{p}(\mathbb{S}^{d})},\quad s\in\mathbb{N}.$ (3.3)
-Another way to characterize the smoothness is through the $K$-functional.
+Another way to characterize the smoothness is through the $K$-functionals.
 We first introduce the fractional Sobolev space induced by the Laplace-Beltrami operator.
@@
 where we assume $f,(-\Delta)^{\alpha/2}f\in C(\mathbb{S}^{d})$ for $p=\infty$. Then we can define the $\alpha$-th $K$-functional of $f\in L^{p}(\mathbb{S}^{d})$ as
-$K_{\alpha}(f,t)_{p}:=\inf_{g\in\mathcal{W}^{\alpha,p}(\mathbb{S}^{d})}\left\{\|f-g\|_{L^{p}(\mathbb{S}^{d})}+t^{\alpha}\|(-\Delta)^{\alpha/2}g\|_{L^{p}(\mathbb{S}^{d})}\right\}.$
-It can be shown [Dai and Xu, 2013, Theorem 10.4.1] that the moduli of smoothness and the $K$-functional are equivalent:
+$K_{\alpha}(f,t)_{p}:=\inf_{g\in\mathcal{W}^{\alpha,p}(\mathbb{S}^{d})}\left\{\|f-g\|_{L^{p}(\mathbb{S}^{d})}+t^{\alpha}\|(-\Delta)^{\alpha/2}g\|_{L^{p}(\mathbb{S}^{d})}\right\},\quad t>0.$
+It can be shown [12, Theorem 10.4.1] that the moduli of smoothness and the $K$-functional are equivalent:
 $\omega_{\alpha}(f,t)_{p}\asymp K_{\alpha}(f,t)_{p}.$ (3.4)
@@
 Abusing the notation, we will also denote $\gamma(g)=\inf_{\mu}\|\mu\|$ as the variation norm of $g\in\mathcal{G}_{\sigma_{k}}(M)$.
-The next proposition transfer our approximation problem on the unit ball $\mathbb{B}^{d}$ to approximation on the sphere $\mathbb{S}^{d}$.
-Proposition 3.3.
+The next proposition transfers our approximation problem on the unit ball $\mathbb{B}^{d}$ to that on the sphere $\mathbb{S}^{d}$.
+Proposition $3.$
 Let $k\in\mathbb{N}_{0}$, $d\in\mathbb{N}$ and $\alpha=r+\beta$ where $r\in\mathbb{N}_{0}$ and $\beta\in(0,1]$.
@@
 Observe that the smoothness of $\widetilde{h}$ on $\Omega$ can be controlled by the smoothness of $h$.
-We can extend $\widetilde{h}$ to $\mathbb{R}^{d+1}$ so that $\|\widetilde{h}\|_{C^{r,\beta}(\mathbb{R}^{d+1})}\leq C_{0}$ for some constant $C_{0}$ independent of $h$, by using (refined version of) Whitney’s extension theorem [Fefferman, 2006, 2009; Fefferman and Israel, 2020].
+We can extend $\widetilde{h}$ to $\mathbb{R}^{d+1}$ so that $\|\widetilde{h}\|_{C^{r,\beta}(\mathbb{R}^{d+1})}\leq C_{0}$ for some constant $C_{0}$ independent of $h$, by using (refined version of) Whitney’s extension theorem [17, 18, 19].
 It remains to show that $\omega_{2s^{*}}(\widetilde{h},t)_{\infty}\lesssim t^{\alpha}$.
 By the equivalence (3.3),
-$\omega_{2s^{*}}(\widetilde{h},t)_{p}$ $\lesssim\sup_{0<\theta\leq t}\sup_{u\in\mathbb{S}^{d}}\left|\sum_{j=0}^{2s^{*}}(-1)^{j}\binom{2s^{*}}{j}\int_{\mathbb{S}^{\perp}_{u}}\widetilde{h}(u\cos j\theta+v\sin j\theta)d\tau_{d-1}(v)\right|$
+$\omega_{2s^{*}}(\widetilde{h},t)_{\infty}$ $\lesssim\sup_{0<\theta\leq t}\sup_{u\in\mathbb{S}^{d}}\left|\sum_{j=0}^{2s^{*}}(-1)^{j}\binom{2s^{*}}{j}\int_{\mathbb{S}^{\perp}_{u}}\widetilde{h}(u\cos j\theta+v\sin j\theta)d\tau_{d-1}(v)\right|$
 $\leq\sup_{0<\theta\leq t}\sup_{u\in\mathbb{S}^{d}}\sup_{v\in\mathbb{S}^{\perp}_{u}}\left|\sum_{j=0}^{2s^{*}}(-1)^{j}\binom{2s^{*}}{j}\widetilde{h}(u\cos j\theta+v\sin j\theta)\right|$
@@
 The binomial theorem shows $H(u,v,\theta)=\widetilde{\Delta}_{\theta}^{2s^{*}}f(0)$.
-Then, classical theory of moduli of smoothness [DeVore and Lorentz, 1993, Chapter 2.6–2.9] implies
+Then, the classical theory of moduli of smoothness [14, Chapter 2.6–2.9] implies
 $\sup_{0<\theta\leq t}|H(u,v,\theta)|\leq\sup_{0<\theta\leq t}|\widetilde{\Delta}_{\theta}^{2s^{*}}f(0)|\lesssim\|f\|_{C^{r,\beta}([0,t_{0}])}t^{\alpha}.$
@@
 We choose $\widetilde{h}$ to be odd (even) if $k$ is even (odd).
-Using $\omega_{s}(\widetilde{h},t)_{2}\leq 2^{s-2s^{*}+2}\omega_{2s^{*}}(\widetilde{h},t)_{2}$ for $s>2s^{*}$ [Dai and Xu, 2013, Proposition 10.1.2] and the Marchaud inequality [Ditzian, 2006, Eq.(9.6)]
+Using $\omega_{s}(\widetilde{h},t)_{2}\leq 2^{s-2s^{*}+2}\omega_{2s^{*}}(\widetilde{h},t)_{2}$ for $s>2s^{*}$ [12, Proposition 10.1.2] and the Marchaud inequality [15, Eq.(9.6)]
 $\omega_{s}(\widetilde{h},t)_{2}\lesssim t^{s}\left(\int_{t}^{1}\frac{\omega_{2s^{*}}(\widetilde{h},\theta)_{2}^{2}}{\theta^{2s+1}}d\theta\right)^{1/2},\quad s<2s^{*},$
@@
 $g=\sum_{n=0}^{\infty}\mathcal{P}_{n}g=\sum_{n=0}^{\infty}\widehat{\sigma_{k}}(n)\mathcal{P}_{n}\phi,$
-which converges in $L^{2}(\mathbb{S}^{d})$. This implies that $g\in\mathcal{F}_{\sigma_{k}}(M)$ if $\mathcal{P}_{n}g=0$ for any $n\in\mathbb{N}_{0}$ satisfying $\widehat{\sigma_{k}}(n)=0$ and
+which converges in $L^{2}(\mathbb{S}^{d})$.
+This implies that $g\in\mathcal{G}_{\sigma_{k}}(M)$ if $g$ is continuous, $\mathcal{P}_{n}g=0$ for any $n\in\mathbb{N}_{0}$ satisfying $\widehat{\sigma_{k}}(n)=0$ and
 $\gamma(g)^{2}\leq\sum_{\widehat{\sigma_{k}}(n)\neq 0}\widehat{\sigma_{k}}(n)^{-2}\|\mathcal{P}_{n}g\|^{2}_{L^{2}(\mathbb{S}^{d})}\leq M^{2}.$
@@
 Since $\widetilde{h}$ is odd (even) if $k$ is even (odd), $\mathcal{P}_{n}g_{m}=\eta(n/m)\mathcal{P}_{n}\widetilde{h}=0$ for any $n\equiv k\bmod 2$.
-Furthermore, Dai and Xu [2013, Theorem 10.3.2] shows that
-$K_{s}(\widetilde{h},m^{-1})_{p}\asymp\|\widetilde{h}-g_{m}\|_{L^{p}(\mathbb{S}^{d})}+m^{-s}\|(-\Delta)^{s/2}g_{m}\|_{L^{p}(\mathbb{S}^{d})}.$
-By the equivalence (3.4) and $\omega_{2s^{*}}(\widetilde{h},m^{-1})_{\infty}\lesssim m^{-\alpha}$, we can bound the approximation error as
+Furthermore, [12, Theorem 10.3.2] shows that
+$K_{s}(\widetilde{h},m^{-1})_{p}\asymp\|\widetilde{h}-g_{m}\|_{L^{p}(\mathbb{S}^{d})}+m^{-s}\|(-\Delta)^{s/2}g_{m}\|_{L^{p}(\mathbb{S}^{d})}.$ (3.6)
+By the equivalence (3.4) and $\omega_{2s^{*}}(\widetilde{h},m^{-1})_{\infty}\lesssim m^{-\alpha}$, the equivalence (3.6) for $p=\infty$ implies that we can bound the approximation error as
 $\|\widetilde{h}-g_{m}\|_{L^{\infty}(\mathbb{S}^{d})}\lesssim m^{-\alpha}.$
-By the estimate (3.5), we get
+Applying the estimate (3.5) to the equivalence (3.6) with $p=2$, we get
 $\|(-\Delta)^{s/2}g_{m}\|_{L^{2}(\mathbb{S}^{d})}\lesssim\begin{cases}1,&\mbox{if }s<\alpha,\\ 1,&\mbox{if }s=\alpha=2s^{*},\\ \sqrt{\log m},&\mbox{if }s=\alpha\mbox{ and }\alpha\neq 2s^{*},\\ m^{s-\alpha},&\mbox{if }\alpha<s\leq 2s^{*}.\end{cases}$
@@
 $\lesssim 1+\|(-\Delta)^{s/2}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})},$
-where we choose $s=(d+2k+1)/2$ in the last inequality. There are three cases:
+where we choose $s=(d+2k+1)/2$ in the last inequality. We continue the proof in three different cases.
 Case I: $\alpha>(d+2k+1)/2$ or $\alpha=(d+2k+1)/2$ is an even integer. In this case, $s<\alpha$ or $s=\alpha=2s^{*}$. Thus,
@@
 $\gamma(\widetilde{h})^{2}$ $\leq\lim_{m\to\infty}\sum_{n\leq m,\widehat{\sigma_{k}}(n)\neq 0}\widehat{\sigma_{k}}(n)^{-2}\|\mathcal{P}_{n}\widetilde{h}\|^{2}_{L^{2}(\mathbb{S}^{d})}$
-$\leq\lim_{m\to\infty}\sum_{\widehat{\sigma_{k}}(n)\neq 0}\widehat{\sigma_{k}}(n)^{-2}\|\mathcal{P}_{n}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})}\lesssim 1.$
-This shows that $\widetilde{h}\subseteq\mathcal{G}_{\sigma_{k}}(M)$ for some constant $M$. Hence, $h=S_{k}\widetilde{h}\in\mathcal{F}_{\sigma_{k}}(M)$ by Proposition 3.3.
+$\leq\lim_{m\to\infty}\sum_{\widehat{\sigma_{k}}(n)\neq 0}\widehat{\sigma_{k}}(n)^{-2}\|\mathcal{P}_{n}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})}$
+$\lesssim 1.$
+This shows that $\widetilde{h}\subseteq\mathcal{G}_{\sigma_{k}}(M)$ for some constant $M$. Hence, $h=S_{k}\widetilde{h}\in\mathcal{F}_{\sigma_{k}}(M)$ by Proposition $3.$
 Case II: $\alpha=(d+2k+1)/2$ is not an even integer. We have $s=\alpha\neq 2s^{*}$ and
@@
 Consequently, we can only obtain exponential convergence rate when $\alpha=(d+2k+1)/2$ is not an even integer.
-We conjecture the bound $\omega_{\alpha}(\widetilde{h},t)_{2}\lesssim t^{\alpha}$ holds for all $\alpha>0$.
+We conjecture the bound $\omega_{s}(\widetilde{h},t)_{2}\lesssim t^{\alpha}$ holds for all $s\geq\alpha$.
 If this is the case, then the proof of Theorem 2.1 implies $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma_{k}}(M)$ for some constant $M$ when $\alpha\geq(d+2k+1)/2$.
@@
 For simplicity, we will only consider ReLU activation function ($k=1$), which is the most popular activation in deep learning.
-To use the uniform approximation bound in Corollary 2.4, we also require that the dimension $d\geq 4$.
-We consider the classical problem of learning a $d$-variate function $h\in\mathcal{H}$ from its noisy samples, where we will assume $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$ or $\mathcal{H}=\mathcal{F}_{\sigma}(1)$.
+We study the classical problem of learning a $d$-variate function $h\in\mathcal{H}$ from its noisy samples, where we will assume $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$ or $\mathcal{H}=\mathcal{F}_{\sigma}(1)$.
 Note that, due to Theorem 2.1, the results for $\mathcal{F}_{\sigma}(1)$ can be applied to $\mathcal{H}^{\alpha}$ with $\alpha>(d+3)/2$ by scaling the variation norm.
@@
 For simplicity, we assume here and in the sequel that the minimum above indeed exists.
-When this is not the case, our results also hold for any estimator which minimizes the empirical risk up to a small error term.
 The performance of the estimation is measured by the expected risk
 $\mathcal{L}(f):=\mathbb{E}_{(X,Y)}[(f(X)-Y)^{2}]=\mathbb{E}_{X\sim\mu}[(f(X)-h(X))^{2}]+V^{2}.$
-It is equivalent to evaluate the estimator by the excess risk
+It is equivalent to evaluating the estimator by the excess risk
 $\|f-h\|_{L^{2}(\mu)}^{2}=\mathcal{L}(f)-\mathcal{L}(h).$
@@
 4.1 Shallow neural networks
-The rate of convergence of neural network regression estimates has been analyzed by many papers [McCaffrey and Gallant, 1994; Kohler and Krzyżak, 2005; Bauer and Kohler, 2019; Schmidt-Hieber, 2020; Nakada and Imaizumi, 2020; Kohler and Langer, 2021].
-It is well-known that the optimal minimax rate of convergence for learning a regression function $h\in\mathcal{H}^{\alpha}$ is $n^{-2\alpha/(d+2\alpha)}$ [Stone, 1982].
-This optimal rate has been established (up to logarithmic factors) for two-hidden-layers neural networks with certain squashing activation functions [Kohler and Krzyżak, 2005] and for deep ReLU neural networks [Schmidt-Hieber, 2020; Kohler and Langer, 2021].
-For shallow networks, McCaffrey and Gallant [1994] proved a rate of $n^{-2\alpha/(2\alpha+d+5)+\epsilon}$ for a certain cosine squasher activation function.
+The rate of convergence of neural network regression estimates has been analyzed by many papers [38, 26, 9, 52, 41, 27].
+It is well-known that the optimal minimax rate of convergence for learning a regression function $h\in\mathcal{H}^{\alpha}$ is $n^{-2\alpha/(d+2\alpha)}$ [59].
+This optimal rate has been established (up to logarithmic factors) for two-hidden-layers neural networks with certain squashing activation functions [26] and for deep ReLU neural networks [52, 27].
+For shallow networks, [38] proved a rate of $n^{-2\alpha/(2\alpha+d+5)+\epsilon}$ with $\epsilon>0$ for a certain cosine squasher activation function.
 However, to the best of our knowledge, it is unknown whether shallow neural networks can achieve the optimal rate.
@@
 It decomposes the error of the ERM into generalization error and approximation error, and bounds the generalization error by the covering number of the hypothesis class $\mathcal{F}_{n}$.
-Lemma 4.1 (Kohler and Langer [2021]).
+Lemma 4.1 ([27]).
 Let $f_{n}^{*}$ be the estimator (4.2) and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$. Then,
@@
 For shallow neural network model $\mathcal{F}_{n}=\mathcal{F}_{\sigma}(N_{n},M_{n})$, Lemma 2.3 and Corollary 2.4 provide bounds for the approximation errors.
-The covering number of $\mathcal{F}_{n}$ can be estimated by using the Lipschitz continuity of the activation function, since the weights in the network are bounded.
+The covering number of the function class $\mathcal{T}_{B_{n}}\mathcal{F}_{n}$ can be estimated by using the pseudo-dimension of $\mathcal{T}_{B_{n}}\mathcal{F}_{n}$ [22].
 Choosing $N_{n},M_{n}$ appropriately to balance the approximation and generalization errors, we can derive convergence rates for the ERM.
 Theorem 4.2.
-Let $d\geq 4$, $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\mathcal{F}_{\sigma}(N_{n},M_{n})$ and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$.
+Let $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\mathcal{F}_{\sigma}(N_{n},M_{n})$ and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$.
   1. (1)
     If $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, we choose
-    $N_{n}\asymp n^{\frac{d}{d+2\alpha}},\quad M_{n}\asymp n^{\frac{d+3-2\alpha}{2d+4\alpha}},$
+    $N_{n}\asymp n^{\frac{d}{d+2\alpha}},\quad M_{n}\gtrsim n^{\frac{d+3-2\alpha}{2d+4\alpha}},$
     then
-    $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}(\log n)^{3}.$
+    $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}(\log n)^{4}.$
   2. (2)
     If $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, we choose
-    $N_{n}\asymp n^{\frac{d}{2d+3}},\quad M_{n}=M,$
-    for some sufficiently large constant $M\geq 1$, then
-    $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{d+3}{2d+3}}(\log n)^{3}.$
+    $N_{n}\asymp n^{\frac{d}{2d+3}},\quad M_{n}\geq 1,$
+    then
+    $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{d+3}{2d+3}}(\log n)^{4}.$
 Proof.
 To apply the bound in Lemma 4.1, we need to estimate the covering number $\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})$.
-It is easy to check that $\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})\leq\mathcal{N}(\epsilon,\mathcal{F}_{n},\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})$.
-Hence, it is enough to provide a bound for $\mathcal{N}(\epsilon,\mathcal{F}_{n},\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})$.
-Recall that any $f\in\mathcal{F}_{n}$ can be represented in the form $f(x)=\sum_{i=1}^{N_{n}}a_{i}\sigma((x^{\intercal},1)v_{i})$ with $v_{i}\in\mathbb{S}^{d}$ and $\sum_{i=1}^{N_{n}}|a_{i}|\leq M_{n}$.
-It is easy to see that there exists $\epsilon/(2\sqrt{2}M_{n})$-cover $V_{\epsilon}$ of $\mathbb{S}^{d}$ such that its cardinality $|V_{\epsilon}|\lesssim(1+M_{n}/\epsilon)^{d}$.
-Similarly, there exists $\epsilon/(2\sqrt{2})$-cover $A_{\epsilon}$ of the set $A=\{a=(a_{1},\dots,a_{N_{n}}):\|a\|_{1}\leq M_{n}\}$ in the metric $\|\cdot\|_{1}$ such that $|A_{\epsilon}|\lesssim(1+M_{n}/\epsilon)^{N_{n}}$.
-Hence, for any $v_{i}\in\mathbb{S}^{d}$ and $a=(a_{1},\dots,a_{N_{n}})\in A$, there exists $\widetilde{v}_{i}\in V_{\epsilon}$ and $\widetilde{a}=(\widetilde{a}_{1},\dots,\widetilde{a}_{N_{n}})\in A_{\epsilon}$ satisfying $\|v_{i}-\widetilde{v}_{i}\|_{2}\leq\epsilon/(2\sqrt{2}M_{n})$ and $\|a-\widetilde{a}\|_{1}\leq\epsilon/(2\sqrt{2})$.
-Therefore, letting $\widetilde{f}(x)=\sum_{i=1}^{N_{n}}\widetilde{a}_{i}\sigma((x^{\intercal},1)\widetilde{v}_{i})$, we have
-$|f(x)-\widetilde{f}(x)|$ $\leq\|a-\widetilde{a}\|_{1}\sup_{1\leq i\leq N_{n}}|\sigma((x^{\intercal},1)v_{i})|+\|\widetilde{a}\|_{1}\sup_{1\leq i\leq N_{n}}|\sigma((x^{\intercal},1)v_{i})-\sigma((x^{\intercal},1)\widetilde{v}_{i})|$
-$\leq\epsilon/2+\sqrt{2}M_{n}\sup_{1\leq i\leq N_{n}}\|v_{i}-\widetilde{v}_{i}\|_{2}\leq\epsilon,$
-where we use the Lipschitz continuity of ReLU in the second inequality. This shows that
-$\log\mathcal{N}(\epsilon,\mathcal{F}_{n},\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})\lesssim N_{n}\log|V_{\epsilon}|+\log|A_{\epsilon}|\lesssim N_{n}\log(1+M_{n}/\epsilon).$
+The classical result of [22, Theorem 6] showed that the covering number can be bounded by pseudo-dimension:
+$\log\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})\lesssim\,{\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})\log(B_{n}/\epsilon),$ (4.3)
+where ${\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})$ is the pseudo-dimension of the function class $\mathcal{T}_{B_{n}}\mathcal{F}_{n}$, see (2.2). For ReLU neural networks, [6] showed that
+${\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})\lesssim N_{n}\log N_{n}.$
+Consequently, we have
+$\log\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})\lesssim N_{n}\log(N_{n})\log(B_{n}/\epsilon).$
 Applying Lemma 4.1 and Corollary 2.4, if $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, then
-$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim\frac{(\log n)^{2}N_{n}\log(nM_{n}B_{n})}{n}+N_{n}^{-\frac{2\alpha}{d}}\lor M_{n}^{-\frac{4\alpha}{d+3-2\alpha}}.$
-By choosing $N_{n}\asymp n^{d/(d+2\alpha)}$ and $M_{n}\asymp N_{n}^{(d+3-2\alpha)/(2d)}$, we get $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-2\alpha/(d+2\alpha)}(\log n)^{3}$.
-Similarly, by Lemma 4.1 and Lemma 2.3, if $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, there exists $M_{n}=M$ such that
-$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim\frac{(\log n)^{2}N_{n}\log(nM_{n}B_{n})}{n}+N_{n}^{-\frac{d+3}{d}}.$
-We choose $N_{n}\asymp n^{d/(2d+3)}$, then $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-(d+3)/(2d+3)}(\log n)^{3}$. ∎
+$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim\frac{(\log n)^{2}N_{n}\log(N_{n})\log(nB_{n}^{2})}{n}+N_{n}^{-\frac{2\alpha}{d}}\lor M_{n}^{-\frac{4\alpha}{d+3-2\alpha}}.$
+By choosing $N_{n}\asymp n^{d/(d+2\alpha)}$ and $M_{n}\gtrsim N_{n}^{(d+3-2\alpha)/(2d)}$, we get $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-2\alpha/(d+2\alpha)}(\log n)^{4}$.
+Similarly, by Lemma 4.1 and Lemma 2.3, if $\mathcal{H}=\mathcal{F}_{\sigma}(1)$ and $M_{n}\geq 1$, then
+$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim\frac{(\log n)^{2}N_{n}\log(N_{n})\log(nB_{n}^{2})}{n}+N_{n}^{-\frac{d+3}{d}}.$
+We choose $N_{n}\asymp n^{d/(2d+3)}$, then $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-(d+3)/(2d+3)}(\log n)^{4}$. ∎
 Remark 4.3.
 The Gaussian noise assumption on the model (4.1) can be weaken for Lemma 4.1 and hence for Theorem 4.2.
-We refer the reader to Kohler and Langer [2021, Appendix B, Lemma 18] for more details.
-Theorem 4.2 can be easily generalized to shallow ReLUk neural networks for $k\geq 1$ by using the same proof technique, if we assume that $\mu$ has a bounded density (so that we can use Corollary 2.4 to bound the approximation error in the $L^{2}(\mu)$ norm).
-For example, one can show that, if $h\in\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$, then we can choose $\mathcal{F}_{n}=\mathcal{F}_{\sigma_{k}}(N_{n},M_{n})$ with $N_{n}\asymp n^{\frac{d}{d+2\alpha}}$ and $M_{n}\asymp n^{\frac{d+2k+1-2\alpha}{2d+4\alpha}}$, such that $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}(\log n)^{3}$.
+We refer the reader to [27, Appendix B, Lemma 18] for more details.
+Theorem 4.2 can be easily generalized to shallow ReLUk neural networks for $k\geq 1$ by using the same proof technique.
+For example, one can show that, if $h\in\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$, then we can choose $\mathcal{F}_{n}=\mathcal{F}_{\sigma_{k}}(N_{n},M_{n})$ with $N_{n}\asymp n^{\frac{d}{d+2\alpha}}$ and $M_{n}\gtrsim n^{\frac{d+2k+1-2\alpha}{2d+4\alpha}}$, such that $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}(\log n)^{4}$.
 Theorem 4.2 shows that least square minimization using shallow ReLU neural networks can achieve the optimal rate $n^{-\frac{2\alpha}{d+2\alpha}}$ for learning functions in $\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$.
-For the function class $\mathcal{F}_{\sigma}(1)$, the rate $n^{-\frac{d+3}{2d+3}}$ is also minimax optimal as proven by Parhi and Nowak [2023, Lemma 25] (they studied a slightly different function class, but their result also holds for $\mathcal{F}_{\sigma}(1)$).
-Specifically, Siegel and Xu [2022, Theorem 4 and Theorem 8] give a sharp estimate for the metric entropy
+For the function class $\mathcal{F}_{\sigma}(1)$, the rate $n^{-\frac{d+3}{2d+3}}$ is also minimax optimal as proven by [46, Lemma 25] (they studied a slightly different function class, but their result also holds for $\mathcal{F}_{\sigma}(1)$).
+Specifically, [57, Theorem 4 and Theorem 8] give a sharp estimate for the metric entropy
 $\log\mathcal{N}(\epsilon,\mathcal{F}_{\sigma}(1),\|\cdot\|_{L^{2}(\mathbb{B}^{d})})\asymp\epsilon^{-\frac{2d}{d+3}}.$
-Combining this estimate with the classical result of Yang and Barron (see [Yang and Barron, 1999, Proposition 1] and [Wainwright, 2019, Chapter 15]), we get
+Combining this estimate with the classical result of Yang and Barron (see [61, Proposition 1] and [60, Chapter 15]), we get
 $\inf_{\widehat{f}_{n}}\sup_{h\in\mathcal{F}_{\sigma}(1)}\mathbb{E}_{\mathcal{D}_{n}}\|\widehat{f}_{n}-h\|_{L^{2}(\mathbb{B}^{d})}^{2}\asymp n^{-\frac{d+3}{2d+3}},$
-where the infimum is over all functions of the samples $\mathcal{D}_{n}$, which are generated from the model (4.1).
+where the infimum taken is over all estimators based on the samples $\mathcal{D}_{n}$, which are generated from the model (4.1).
 4.2 Deep neural networks and over-parameterization
 There is a direct way to generalize the analysis in the last section to deep neural networks: we can implement shallow neural networks by sparse multi-layer neural networks with the same order of parameters, and estimate the approximation and generalization performance of the constructed networks.
-Since the optimal convergence rates of deep neural networks have already been established in [Schmidt-Hieber, 2020; Kohler and Langer, 2021], we do not pursue in this direction.
-Instead, we study the convergence rates of over-parameterized neural networks by using the idea discussed in [Jiao et al., 2023].
+Since the optimal convergence rates of deep neural networks have already been established in [52, 27], we do not pursue in this direction.
+Instead, we study the convergence rates of over-parameterized neural networks by using the idea discussed in [24].
 The reason for studying such networks is that, in modern applications of deep learning, the number of parameters in the networks is often much larger than the number of samples.
-However, in the convergence analysis of [Schmidt-Hieber, 2020; Kohler and Langer, 2021], the network that achieves the optimal rate is under-parameterized (see also the choice of $N_{n}$ in Theorem 4.2).
-Hence, these analysis can not explain the empirical performance of deep learning models used in practice.
-Following Jiao et al.
-[2023], we consider deep neural networks with norm constraints on weight matrices.
+However, in the convergence analysis of [52, 27], the network that achieves the optimal rate is under-parameterized (see also the choice of $N_{n}$ in Theorem 4.2).
+Hence, the analysis can not explain the empirical performance of deep learning models used in practice.
+Following [24], we consider deep neural networks with norm constraints on weight matrices.
 For $W,L\in\mathbb{N}$, we denote by $\mathcal{NN}(W,L)$ the set of functions that can be parameterized by ReLU neural networks in the form
@@
 where we use $\|A\|:=\sup_{\|x\|_{\infty}\leq 1}\|Ax\|_{\infty}$ to denote the operator norm (induced by the $\ell^{\infty}$ norm) of a matrix $A=(a_{i,j})\in\mathbb{R}^{m\times n}$.
-It is well-known that $\|A\|$ is the maximum $1$-norm of the rows of $A$:
+It is well-known that $\|A\|$ is the maximum 11-norm of the rows of AA:
 $\|A\|=\max_{1\leq i\leq m}\sum_{j=1}^{n}|a_{i,j}|.$
-The motivation for such a definition of $\kappa(\theta)$ is discussed in [Jiao et al., 2023].
-For $M\geq 0$, we denote by $\mathcal{NN}(W,L,M)$ as the set of functions $f_{\theta}\in\mathcal{NN}$ that satisfy $\kappa(\theta)\leq M$.
-It is shown by [Jiao et al., 2023, Proposition 2.5] that, if $W_{1}\leq W_{2},L_{1}\leq L_{2},M_{1}\leq M_{2}$, then $\mathcal{NN}(W_{1},L_{1},M_{1})\subseteq\mathcal{NN}(W_{2},L_{2},M_{2})$.
-(Strictly speaking, Jiao et al.
-[2023] use the convention that the bias $b^{(L)}=0$ in the last layer.
-But the results can be easily generalized to the case $b^{(L)}\neq 0$, see [Yang, 2022, Section 2.1] for details.)
+The motivation for such a definition of $\kappa(\theta)$ is discussed in [24].
+For $M\geq 0$, we denote by $\mathcal{NN}(W,L,M)$ as the set of functions $f_{\theta}\in\mathcal{NN}(W,L)$ that satisfy $\kappa(\theta)\leq M$.
+It is shown by [24, Proposition 2.5] that, if $W_{1}\leq W_{2},L_{1}\leq L_{2},M_{1}\leq M_{2}$, then $\mathcal{NN}(W_{1},L_{1},M_{1})\subseteq\mathcal{NN}(W_{2},L_{2},M_{2})$.
+(Strictly speaking, [24] use the convention that the bias $b^{(L)}=0$ in the last layer.
+But the results can be easily generalized to the case $b^{(L)}\neq 0$, see [62, Section 2.1] for details.)
 To derive approximation bounds for deep neural networks, we consider the relationship of $\mathcal{F}_{\sigma}(N,M)$ and $\mathcal{NN}(N,1,M)$.
 The next proposition shows the function classes $\mathcal{NN}(N,1,M)$ and $\mathcal{F}_{\sigma}(N,M)$ have essentially the same approximation power.
-Proposition 4.4.
+Proposition $4.$
 For any $N\in\mathbb{N}$ and $M>0$, we have $\mathcal{F}_{\sigma}(N,M)\subseteq\mathcal{NN}(N,1,\sqrt{d+1}M)$ and $\mathcal{NN}(N,1,M)\subseteq\mathcal{F}_{\sigma}(N+1,M)$.
@@
 Since $v_{i}\in\mathbb{S}^{d}$, it is easy to see that $\kappa(\theta)\leq\sqrt{d+1}M$. Hence, $\mathcal{F}_{\sigma}(N,M)\subseteq\mathcal{NN}(N,1,\sqrt{d+1}M)$.
-On the contrary, let $f_{\theta}\in\mathcal{NN}(N,1,M)$ be a function parameterized in the form (4.3) with $(A^{(0)},b^{(0)})=(a^{(0)}_{1},\dots,a^{(0)}_{N})^{\intercal}$ and $(A^{(1)},b^{(1)})=(a^{(1)}_{1},\dots,a^{(1)}_{N},b^{(1)})$, where $a^{(0)}_{i}\in\mathbb{R}^{d+1}$ and $a^{(1)}_{i},b^{(1)}\in\mathbb{R}$.
+On the other side, let $f_{\theta}\in\mathcal{NN}(N,1,M)$ be a function parameterized in the form (4.4) with $(A^{(0)},b^{(0)})=(a^{(0)}_{1},\dots,a^{(0)}_{N})^{\intercal}$ and $(A^{(1)},b^{(1)})=(a^{(1)}_{1},\dots,a^{(1)}_{N},b^{(1)})$, where $a^{(0)}_{i}\in\mathbb{R}^{d+1}$ and $a^{(1)}_{i},b^{(1)}\in\mathbb{R}$.
 Then, $f_{\theta}$ can be represented as
@@
 Corollary 4.5.
-Let $d\geq 4$.
-  1. (1)
 For $\mathcal{H}^{\alpha}$ with $0<\alpha<(d+3)/2$, we have
 $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{NN}(W,L,M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim W^{-\frac{\alpha}{d}}\lor M^{-\frac{2\alpha}{d+3-2\alpha}}.$
-  2. (2)
 For $\mathcal{F}_{\sigma}(1)$, there exists a constant $M\geq 1$ such that
@@
 Proof.
-Part (1) is a direct consequence of Corollary 2.4 and $\mathcal{F}_{\sigma}(W,M)\subseteq\mathcal{NN}(W,L,\sqrt{d+1}M)$. Part (2) follows from Lemma 2.3. ∎
+The first part is a direct consequence of Corollary 2.4 and the inclusion $\mathcal{F}_{\sigma}(W,M)\subseteq\mathcal{NN}(W,L,\sqrt{d+1}M)$.
+The second part follows from Lemma 2.3 and we can choose $M=\sqrt{d+1}$.
+∎
 In the first part of Corollary 4.5, if we allow the width $W$ to be arbitrary large, say $W\gtrsim M^{2d/(d+3-2\alpha)}$, then we can bound the approximation error by the size of weights.
 Hence, this result can be applied to over-parameterized neural networks.
-(Note that, in Theorem 4.2, we use a different regime of the bound.) In part (2), the size of weights is bounded by a constant.
+(Note that, in Theorem 4.2, we use a different regime of the bound.) For the approximation of $\mathcal{F}_{\sigma}(1)$, the size of weights is bounded by a constant.
 We will show that this constant can be used to control the generalization error.
@@
 Hence, it can also be applied to over-parameterized neural networks.
-The approximation rate $M^{-\frac{2\alpha}{d+3-2\alpha}}$ for $\mathcal{H}^{\alpha}$ in Corollary 4.5 improves the rate $M^{-\frac{\alpha}{d+1}}$ proved by Jiao et al.
-[2023].
-Using the upper bound for Rademacher complexity of $\mathcal{NN}(W,L,M)$ (see Lemma 4.6), Jiao et al.
-[2023] also gave an approximation lower bound $(M\sqrt{L})^{-\frac{2\alpha}{d-2\alpha}}$.
+The approximation rate $M^{-\frac{2\alpha}{d+3-2\alpha}}$ for $\mathcal{H}^{\alpha}$ in Corollary 4.5 improves the rate $M^{-\frac{\alpha}{d+1}}$ proven by [24].
+Using the upper bound for Rademacher complexity of $\mathcal{NN}(W,L,M)$ (see Lemma 4.6), [24] also gave an approximation lower bound $(M\sqrt{L})^{-\frac{2\alpha}{d-2\alpha}}$.
 For fixed depth $L$, our upper bound is very close to this lower bound.
 We conjecture that the rate in Corollary 4.5 is optimal with respect to $M$ (for fixed depth $L$).
-The discussion of optimality in the last part of Section 2 implies that the conjecture is true for shallow neural networks.
+The discussion of optimality at the end of Section 2 implies that the conjecture is true for shallow neural networks (i.e.
+$L=1$).
 To control the generalization performance of over-parameterized neural networks, we need to have size-independent sample complexity bounds for such networks.
-Several methods have been applied to obtain such kind of bounds in recent works [Neyshabur et al., 2015, 2018; Bartlett et al., 2017; Golowich et al., 2020].
-Here, we will use the result of Golowich et al.
-[2020], which estimates the Rademacher complexity of deep neural networks [Bartlett and Mendelson, 2002].
+Several methods have been applied to obtain such kind of bounds in recent works [43, 42, 5, 21].
+Here, we will use the result of [21], which estimates the Rademacher complexity of deep neural networks [7].
 For a set $S\subseteq\mathbb{R}^{n}$, let us denote its Rademacher complexity by
 $\mathcal{R}_{n}(S):=\mathbb{E}_{\xi_{1:n}}\left[\sup_{(s_{1},\dots,s_{n})\in S}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}s_{i}\right],$
-where $\xi_{1:n}=(\xi_{1},\dots,\xi_{n})$ is a sequence of i.i.d. Rademacher random variables. The following lemma is from Golowich et al. [2020, Theorem 3.2] and Jiao et al. [2023, Lemma 2.3].
-Lemma 4.6 (Golowich et al. [2020]).
+where $\xi_{1:n}=(\xi_{1},\dots,\xi_{n})$ is a sequence of i.i.d. Rademacher random variables. The following lemma is from [21, Theorem 3.2] and [24, Lemma 2.3].
+Lemma 4.6.
 For any $x_{1},\dots,x_{n}\in[-1,1]^{d}$, let $S:=\{(f(x_{1}),\dots,f(x_{n})):f\in\mathcal{NN}(W,L,M)\}\subseteq\mathbb{R}^{n}$, then
@@
 As usual, we decompose the excess risk of the ERM into approximation error and generalization error, and bound them by Corollary 4.5 and Lemma 4.6, respectively.
-Note that the convergence rates in the following theorem are very close to the optimal rates in Theorem 4.2.
+Note that the convergence rates in the following theorem are worse than the optimal rates in Theorem 4.2.
 Theorem 4.7.
-Let $d\geq 4$ and $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\{\mathcal{T}_{B_{n}}f:f\in\mathcal{NN}(W_{n},L,M_{n})\}$, where $L\in\mathbb{N}$ is a fixed constant, $1\leq B_{n}\lesssim\log n$ in case (1) and $\sqrt{2}\leq B_{n}\lesssim\log n$ in case (2).
+Let $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\{\mathcal{T}_{B_{n}}f:f\in\mathcal{NN}(W_{n},L,M_{n})\}$, where $L\in\mathbb{N}$ is a fixed constant, $1\leq B_{n}\lesssim\log n$ in case (1) and $\sqrt{2}\leq B_{n}\lesssim\log n$ in case (2).
   1. (1)
     If $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, we choose
-    $M_{n}\asymp n^{\frac{1}{2}-\frac{2\alpha}{d+3+2\alpha}},\quad W_{n}\gtrsim n^{\frac{d}{d+3+2\alpha}},$
+    $W_{n}\gtrsim n^{\frac{d}{d+3+2\alpha}},\quad M_{n}\asymp n^{\frac{1}{2}-\frac{2\alpha}{d+3+2\alpha}},$
     then
@@
   2. (2)
-    If $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, then there exist a constant $M\geq 1$ such that for $M_{n}=M$ and any $W_{n}\gtrsim n^{d/(2d+6)}$, we have
-    $\mathbb{E}_{\mathcal{D}_{n}}\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-1/2}\log n.$
+    If $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, we choose a large enough constant $M$ and let
+    $W_{n}\gtrsim n^{\frac{d}{2d+6}},\quad M_{n}=M,$
+    then
+    $\mathbb{E}_{\mathcal{D}_{n}}\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{1}{2}}\log n.$
 Proof.
-The proof is essentially the same as Jiao et al. [2023, Theorem 4.1]. Observe that, for any $f\in\mathcal{F}_{n}$,
+The proof is essentially the same as [24, Theorem 4.1]. Observe that, for any $f\in\mathcal{F}_{n}$,
 $\|f_{n}^{*}-h\|^{2}_{L^{2}(\mu)}=\mathcal{L}(f_{n}^{*})-\mathcal{L}(h)$
@@
 $=$ $\mathbb{E}_{\mathcal{D}_{n}}\left[\|f_{n}^{*}-h\|^{2}_{L^{2}(\mu)}+V^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}(f_{n}^{*}(X_{i})-h(X_{i}))^{2}-2\eta_{i}(f_{n}^{*}(X_{i})-h(X_{i}))+\eta_{i}^{2}\right)\right]$
-$=$ $\mathbb{E}_{X_{1:n}}\left[\|f_{n}^{*}-h\|^{2}_{L^{2}(\mu)}-\frac{1}{n}\sum_{i=1}^{n}(f_{n}^{*}(X_{i})-h(X_{i}))^{2}\right]+2\mathbb{E}_{\mathcal{D}_{n}}\left[\frac{1}{n}\sum_{i=1}^{n}\eta_{i}(f_{n}^{*}(X_{i})-h(X_{i}))\right]$
+$=$ $\mathbb{E}_{\mathcal{D}_{n}}\left[\|f_{n}^{*}-h\|^{2}_{L^{2}(\mu)}-\frac{1}{n}\sum_{i=1}^{n}(f_{n}^{*}(X_{i})-h(X_{i}))^{2}\right]+2\mathbb{E}_{\mathcal{D}_{n}}\left[\frac{1}{n}\sum_{i=1}^{n}\eta_{i}(f_{n}^{*}(X_{i})-h(X_{i}))\right]$
 $\leq$ $\mathbb{E}_{X_{1:n}}\left[\sup_{\phi\in\Phi_{n}}\left(\mathbb{E}_{X}[\phi^{2}(X)]-\frac{1}{n}\sum_{i=1}^{n}\phi^{2}(X_{i})\right)\right]+2\mathbb{E}_{X_{1:n}}\mathbb{E}_{\eta_{1:n}}\left[\sup_{\phi\in\Phi_{n}}\frac{1}{n}\sum_{i=1}^{n}\eta_{i}\phi(X_{i})\right],$ (4.6)
-where we denote $\Phi_{n}:=\{f-h:f\in\mathcal{F}_{n}\}$. By a standard symmetrization argument (see Wainwright [2019, Theorem 4.10]), we can bound the first term in (4.5) by the Rademacher complexity:
+where we denote $\Phi_{n}:=\{f-h:f\in\mathcal{F}_{n}\}$. By a standard symmetrization argument (see [60, Theorem 4.10]), we can bound the first term in (4.6) by the Rademacher complexity:
 $\mathbb{E}_{X_{1:n}}\left[\sup_{\phi\in\Phi_{n}}\left(\mathbb{E}_{X}[\phi^{2}(X)]-\frac{1}{n}\sum_{i=1}^{n}\phi^{2}(X_{i})\right)\right]\leq 2\mathbb{E}_{X_{1:n}}\left[\mathcal{R}_{n}(\Phi_{n}^{2}(X_{1:n}))\right],$
@@
 Hence, we always have $\|h\|_{L^{\infty}(\mathbb{B}^{d})}\leq\sqrt{2}$ and $\|\phi\|_{L^{\infty}(\mathbb{B}^{d})}\leq 2B_{n}$ for any $\phi\in\Phi_{n}$.
-By the structural properties of Rademacher complexity [Bartlett and Mendelson, 2002, Theorem 12],
+By the structural properties of Rademacher complexity [7, Theorem 12],
 $\mathbb{E}_{X_{1:n}}\left[\mathcal{R}_{n}(\Phi_{n}^{2}(X_{1:n}))\right]$ $\leq 8B_{n}\mathbb{E}_{X_{1:n}}\left[\mathcal{R}_{n}(\Phi_{n}(X_{1:n}))\right]$
@@
 $\lesssim\frac{M_{n}\log n}{\sqrt{n}},$
-where we apply Lemma 4.6 in the last inequality. Note that the second term in (4.5) is a Gaussian complexity. We can also bound it by the Rademacher complexity [Bartlett and Mendelson, 2002, Lemma 4]:
+where we apply Lemma 4.6 in the last inequality. Note that the second term in (4.6) is a Gaussian complexity. We can also bound it by the Rademacher complexity [7, Lemma 4]:
 $\mathbb{E}_{X_{1:n}}\mathbb{E}_{\eta_{1:n}}\left[\sup_{\phi\in\Phi_{n}}\frac{1}{n}\sum_{i=1}^{n}\eta_{i}\phi(X_{i})\right]\lesssim\mathbb{E}_{X_{1:n}}\left[\mathcal{R}_{n}(\Phi_{n}(X_{1:n}))\right]\log n\lesssim\frac{M_{n}\log n}{\sqrt{n}}.$
@@
 $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{n}}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim W_{n}^{-\frac{\alpha}{d}}\lor M_{n}^{-\frac{2\alpha}{d+3-2\alpha}}.$
-Combining with (4.4) and (4.6), if we choose $M_{n}\asymp n^{\frac{1}{2}-\frac{2\alpha}{d+3+2\alpha}}$ and $W_{n}\gtrsim n^{\frac{d}{d+3+2\alpha}}$, then
+Combining with (4.5) and (4.7), we know that if we choose $M_{n}\asymp n^{\frac{1}{2}-\frac{2\alpha}{d+3+2\alpha}}$ and $W_{n}\gtrsim n^{\frac{d}{d+3+2\alpha}}$, then
 $\mathbb{E}_{\mathcal{D}_{n}}\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim W_{n}^{-\frac{2\alpha}{d}}\lor M_{n}^{-\frac{4\alpha}{d+3-2\alpha}}+\frac{M_{n}\log n}{\sqrt{n}}\lesssim n^{-\frac{2\alpha}{d+3+2\alpha}}\log n.$
@@
 4.3 Convolutional neural networks
-In contrast to the vast amount of theoretical studies on fully connected neural networks, there are only a few papers analyzing the performance of convolutional neural networks [Zhou, 2020a, b; Fang et al., 2020; Petersen and Voigtlaender, 2020; Lin et al., 2022].
-The recent work [Lin et al., 2022] showed the universal consistency of CNNs for nonparametric regression.
+In contrast to the vast amount of theoretical studies on fully connected neural networks, there are only a few papers analyzing the performance of convolutional neural networks [66, 67, 16, 47, 30, 20, 35, 68].
+The recent work [30] showed the universal consistency of CNNs for nonparametric regression.
 In this section, we show how to use our approximation results to analyze the convergence rates of CNNs.
-Following [Zhou, 2020b], we introduce a sparse convolutional structure on deep neural networks.
+Following [67], we introduce a sparse convolutional structure on deep neural networks.
 Let $s\geq 2$ be a fixed integer, which is used to control the filter length.
@@
 The assumption on the special form (4.8) of biases is used to reduce the free parameters in the network.
-As in [Zhou, 2020b], one can compute that the number of free parameters in $\mathcal{CNN}(s,L)$ is (5s+2)L+2d2s(5s+2)L+2d-2s, which grows linearly on $L$.
+As in [67], one can compute that the number of free parameters in $\mathcal{CNN}(s,L)$ is (5s+2)L+2d2s(5s+2)L+2d-2s, which grows linearly on $L$.
 The next proposition shows that all functions in $\mathcal{NN}(N,1)$ can be implemented by CNNs.
-Proposition 4.8 (Zhou [2020b]).
+Proposition 4.8 ([67]).
 If $N,L\in\mathbb{N}$ satisfy $L\geq\lfloor\frac{Nd}{s-1}+1\rfloor$, then $\mathcal{NN}(N,1)\subseteq\mathcal{CNN}(s,L)$.
 Proof.
-This result is proved in Zhou [2020b, Proof of Theorem 2].
+This result is proven in [67, Proof of Theorem 2].
 We only give a sketch of the construction for completeness.
@@
 $(v_{Nd-1},\dots,v_{0})=(a_{N}^{\intercal},\dots,a_{1}^{\intercal}).$
-Applying Zhou [2020b, Theorem 3] to the sequence $v$, we can construct filters $\{w^{(\ell)}\}_{\ell=0}^{L-1}$ supported on $\{0,1,\dots,s\}$ such that $v=w^{(L-1)}*w^{(L-2)}*\dots*w^{(0)}$, which implies $A^{w^{(L-1)}}\cdots A^{w^{(0)}}=A^{v}\in\mathbb{R}^{(d+Ls)\times d}$.
-Note that, by definition, for $i=1,\dots,N$, the $id$-th row of $A^{v}$ is exactly $a_{i}^{\intercal}$.
+Applying [67, Theorem 3] to the sequence $v$, we can construct filters $\{w^{(\ell)}\}_{\ell=0}^{L-1}$ supported on $\{0,1,\dots,s\}$ such that $v=w^{(L-1)}*w^{(L-2)}*\dots*w^{(0)}$, which implies $A^{w^{(L-1)}}\cdots A^{w^{(0)}}=A^{v}\in\mathbb{R}^{(d+Ls)\times d}$.
+Note that, by definition, for $i=1,\dots,N$, the idid-th row of $A^{v}$ is exactly $a_{i}^{\intercal}$.
 Then, for $\ell=0,\dots,L-2$, we can choose $b^{(\ell)}$ satisfying (4.8) such that $f^{(\ell+1)}(x)=A^{w^{(\ell)}}\cdots A^{w^{(0)}}x+B^{(\ell)}$, where $B^{(\ell)}>0$ is a sufficiently large constant that makes the components of $f^{(\ell+1)}(x)$ positive for all $x\in\mathbb{B}^{d}$.
@@
 Corollary 4.9.
-Let $s\geq 2$ and $d\geq 4$ be integers.
+Let $s\geq 2$ be an integer.
   1. (1)
@@
 ∎
-Since the number of parameters in $\mathcal{CNN}(s,L)$ is approximately $L$, the rate $\mathcal{O}(L^{-\alpha/d})$ in part (1) of Corollary 4.9 is the same as the rate in [Yarotsky, 2017] for fully connected neural networks.
-However, Yarotsky [2018]; Lu et al.
-[2021] showed that this rate can be improved to $\mathcal{O}(L^{-2\alpha/d})$ for fully connected neural networks by using the bit extraction technique [Bartlett et al., 2019].
+Since the number of parameters in $\mathcal{CNN}(s,L)$ is approximately $L$, the rate $\mathcal{O}(L^{-\alpha/d})$ in part (1) of Corollary 4.9 is the same as the rate in [64] for fully connected neural networks.
+However, [65, 31] showed that this rate can be improved to $\mathcal{O}(L^{-2\alpha/d})$ for fully connected neural networks by using the bit extraction technique [6].
 It would be interesting to see whether this rate also holds for $\mathcal{CNN}(s,L)$.
-As in Theorem 4.2, we use Lemma 4.1 to decompose the error and bound the approximation error by Corollary 4.9.
-But we need to use pseudo-dimension to bound the covering number, because we do not have estimate for the size of weights in CNNs.
+As in Theorem 4.2, we use Lemma 4.1 to decompose the error and bound the approximation error by Corollary 4.9. The covering number is bounded again by pseudo-dimension.
 Theorem 4.10.
-Let $d\geq 4$ and $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\mathcal{CNN}(s,L_{n})$, where $s\geq 2$ is a fixed integer, and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$.
+Let $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\mathcal{CNN}(s,L_{n})$, where $s\geq 2$ is a fixed integer, and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$.
   1. (1)
@@
 Proof.
-To apply Lemma 4.1, we need to estimate the covering number $\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})$.
-The classical result of Haussler [1992, Theorem 6] showed that the covering number can be bounded by pseudo-dimension:
-$\log\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})\lesssim\,{\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})\log(B_{n}/\epsilon),$
-where ${\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})$ is the pseudo-dimension of the function class $\mathcal{T}_{B_{n}}\mathcal{F}_{n}$, see (2.2).
-For ReLU neural networks, Bartlett et al.
-[2019] gave the following estimate of the pseudo-dimension:
+The proof is the same as Theorem 4.2 and [68].
+We can use (4.3) to bound the covering number by the pseudo-dimension.
+For convolutional neural networks, [6] gave the following estimate of the pseudo-dimension:
 ${\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})\lesssim L_{n}p(s,L_{n})\log(q(s,L_{n}))\lesssim L_{n}^{2}\log L_{n},$
@@
 We choose $L_{n}\asymp n^{d/(3d+3)}$, then $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-(d+3)/(3d+3)}(\log n)^{4}$. ∎
+Finally, we note that the recent paper [68] also studied the convergence of CNNs and proved the rate $\mathcal{O}(n^{-1/3}(\log n)^{2})$ for $\mathcal{H}^{\alpha}$ with $\alpha>(d+4)/2$.
+The convergence rate we obtained in Theorem 4.10 for $\mathcal{F}_{\sigma}(1)$, which includes $\mathcal{H}^{\alpha}$ with $\alpha>(d+3)/2$ by Theorem 2.1, is slightly better than their rate.
 5 Conclusion
@@
 There are a few interesting questions we would like to propose for future research.
-First, for shallow neural networks approximation, the optimal rate $\mathcal{O}(N^{-\alpha/d})$ in the supremum norm (in Corollary 2.4) is established only for ReLU and Heaviside activations.
-Can this result be extended to all ReLUk activations?
+First, for approximation by shallow neural networks, we establish the optimal rate in the supremum norm by using the results of [55] (Lemma 2.3).
+The paper [55] actually showed that approximation bounds similar to Lemma 2.3 also hold in Sobolev norms.
+We think it is a promising direction to extend our approximation results in the supremum norm (Theorem 2.1 and Corollary 2.4) to the Sobolev norms.
 Second, it is unclear whether over-parameterized neural networks can achieve the optimal rate for learning functions in $\mathcal{H}^{\alpha}$.
 It seems that refined generalization error analysis is needed.
-Finally, it would be interesting to extend the theory developed in this paper to general activation functions.
+Finally, it would be interesting to extend the theory developed in this paper to general activation functions and study how the results are affected by the activation functions.
 Acknowledgments
 The work described in this paper was partially supported by InnoHK initiative, The Government of the HKSAR, Laboratory for AI-Powered Financial Technologies, the Research Grants Council of Hong Kong [Projects No.
-CityU 11308121, 11306220, 11308020 and No.
-N_CityU102/20] and National Natural Science Foundation of China [Projects No.
-11461161006].
+CityU 11306220 and 11308020] and National Natural Science Foundation of China [Project No.
+12371103] when the second author worked at City University of Hong Kong.
+We thank the referees for their helpful comments and suggestions on the paper.